### v1

# Weighted shift operators on spaces of analytic functions over an annulus - I

Abstract.
We initiate a study of the forward weighted shift operator, denoted by $F_{w}$, defined on the Banach spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ of analytic functions on a suitable annulus $\mathtt{A}$ in the complex plane. Here, $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a.b}(\mathtt{A})$, respectively, are the spaces of analytic functions on some annulus $\mathtt{A}$, having normalized Schauder bases of the form $\{(a_{n}+b_{n}z)z^{n},~{}~{}n\in\mathbb{Z}\}$, which are equivalent to the standard basis in $\ell^{p}(\mathbb{Z})$ and $c_{0}(\mathbb{Z})$. We first obtain necessary and sufficient conditions for $F_{w}$ to be bounded, and then we primarily study when $F_{w}$ and its adjoint $F_{w}^{*}$ are hypercyclic, supercyclic, and chaotic. It is proved that, under certain conditions on $\{a_{n}\}$ and $\{b_{n}\}$, the operator $F_{w}$ similar to a compact perturbation of a classical bilateral weighted shift on $\ell^{p}(\mathbb{Z})$. Unlike the case of classical weighted shifts, the adjoint operator $F_{w}^{*}$ on the dual of $\ell^{p}_{a,b}(\mathtt{A})$ can have non-trivial periodic vectors, without being even hypercyclic. The zero-one law of orbital limit points fails for $F_{w}^{*}$.
Key words and phrases:
bilateral weighted shift operator, hypercyclicity, supercyclicity, chaos, periodic vector
2010 Mathematics Subject Classification:
Primary 47A16, 46E22, 32K05, 47B32; Secondary 47B37, 37A99.
1. Introduction
Among various classes of operators that are commonly investigated in operator theory and linear dynamics, the class of weighted shifts has a predominant role. Shift operators are fairly simple to define, yet they yield diverse and appealing results. In operator theory, weighted shifts are well studied in connections with model theory (representation of various operators using shifts), invariant subspaces, cyclic vectors, subnormality, and so on, cf. Halmos [24], and Shields [38]. In the context of linear dynamics it is a usual practice that dynamical properties are first tested for weighted shifts defined on the sequence spaces $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$. Important dynamical notions such as hypercyclicity, chaos, and ergodicity are well distinguishable using weighted shifts. Motivated these, in this paper we introduce and study weighted shift operators on spaces $\ell^{p}_{a,b}(\mathbb{D})$ and $\ell^{p}_{a,b}(\mathtt{A})$ having a close resemblance to the weighted $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$ spaces, highlighting the similarities and differences in the dynamics. These newly introduced spaces and weighted shifts are of independent interests as well.
Definition 1.1.
An operator $T$ on a separable Banach space $X$ is said to be
  • hypercyclic if there exists $x\in X$, known as a hypercyclic vector for $T$, such that the orbit $\{x,Tx,T^{2}x,\cdots\}$ is dense in $X$,
  • chaotic if $T$ is hypercyclic and the set $\{v\in X:T^{m}v=v,~{}for~{}some~{}m\in\mathbb{N}\}$ of periodic vectors is dense in $X$,
  • topologically transitive if, for two non-empty open sets $U_{1}$ and $U_{2}$ of $X$, there exists a natural number $k$ such that $T^{k}(U_{1})\cap U_{2}\neq\phi$.
  • topologically mixing on $X$ if, for any two non-empty open sets $U_{1}$ and $U_{2}$ of $X$, there exists $N$, a natural number, such that $T^{n}(U_{1})\cap U_{2}\neq\phi$ for all $n\geq N$.
We recall that the periodic vectors for an operator $T$ defined on a complex vector space are exactly the linear combinations of eigenvectors corresponding to “rational” unimodular eigenvalues.
Rolewicz [35] showed that $\lambda B$ is hypercyclic on $\ell^{p}(\mathbb{N})$, $1\leq p<\infty$, where $|\lambda|>1$ and $B$ is the unweighted backward shift operator. Also, for particular classes of hypercyclic weighted shifts, see Kitai [25], Gethner and Shapiro [19], and Godefroy and Shapiro [20]. A complete characterization of hypercyclic weighted shifts on $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$ was obtained by Salas [36]. In [20] the authors introduced and studied the notion of chaos for operators, and established that several familiar operators including weighted shifts are chaotic. A complete characterization of hypercyclic and chaotic weighted shifts in a general $F$-spaces was obtained by Grosse-Erdmann [22]. We also note that the weighted shifts which are (topologically) mixing were characterized by Costakis and Sambarino [13]. The dynamics of the unweighted backward shift on certain Bergman spaces was studied in [32]. For account on the theory on hypercyclicity and related topics, we refer to the monographs by Bayart and Matheron [3], and Grosse-Erdmann and Peris [23]. In practice, the existence of hypercyclic and chaotic properties of operators is established using well known criteria that are stated below, cf. [3] and [23]. We will make use of these standard criteria to obtain the hypercyclic and chaotic properties of weighted shifts defined on the weighted spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ which are introduced in Section $2$.
Theorem 1.2.
(Gethner-Shapiro Criterion [19]) Let $T$ be a bounded operator on a separable Banach space $X$, and let $X_{0}$ be a dense subset of $X$. If $\{n_{k}\}\subseteq\mathbb{N}$ is a strictly increasing sequence and $S:X_{0}\mapsto X_{0}$ is a map such that, for each $x\in X_{0}$, $lim_{k\rightarrow\infty}T^{n_{k}}x=0=\lim_{k\rightarrow\infty}S^{n_{k}}x,$ and $TSx=x$, then $T$ is hypercyclic. in addition, if $n_{k}=k$ for all $k\geq 1$, then $T$ is mixing on $X$.
A similar criterion, known as the chaoticity criterion, has been used to obtain chaotic operators in $F$-spaces, and we refer to [9] where it is originally stated. This criterion is very strong: indeed, if an operator $T$ satisfies the chaoticity criterion, then $T$ is topologically mixing, chaotic, frequently hypercyclic, and it has invariant measures of full support, cf. [3].
Theorem 1.3.
(Chaoticity Criterion [9]) Let $X$ be a separable Banach space, $X_{0}$ be a dense set in $X$, and let $T$ be a bounded operator on $X$. If there exists a map $D:X_{0}\rightarrow X_{0}$ such that $\sum_{n\geq 0}T^{n}x$ and $\sum_{n\geq 0}S^{n}x$ are unconditionally convergent, and $TSx=x$, for each $x\in X_{0}$, then the operator $T$ is chaotic and mixing on $X$.
An operator $T$ on a separable Banach space $X$ over $\mathbb{K}:=\mathbb{R}$ or $\mathbb{C}$ is supercyclic if the projective orbit $\{\alpha T^{n}(x):~{}\alpha\in\mathbb{K},n\geq 0\}$ is dense in $X$. Supercyclicity was initially studied by Hilden and Wallen [26] who provided classes of supercyclic weighted shifts. A complete characterization of supercyclic unilateral and bilateral weighted shifts was obtained by Salas [37], and further studied by Montes-Rodriguez and Salas [31]. Along the same lines of the hypercyclicity and chaoticity criteria, a general sufficient condition for supercyclicity is well-known, as stated below.
Theorem 1.4.
(Supercyclicity Criterion [31]) For a bounded operator $T$ on a separable Banach space $X$, if there exist a dense subset $X_{0}$, a sequence $(n_{k})$ of natural numbers, and a map $S:X_{0}\rightarrow X_{0}$ such that, for all $x,y\in X_{0}$, $\lim_{k\rightarrow\infty}\|T^{n_{k}}x\|\|S^{k}y\|=0$ and $TSx=x$, then $T$ is supercyclic.
Motivated by the classical bilateral shifts, in this paper we initiate a study of weighted bilateral forward shift operators $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. In Section $2$, we introduce the aforementioned spaces as analytic function spaces and prove that the co-ordinate functionals on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ are continuous. (Although sequential analogues of these spaces can be defined, we prefer to stick to the analytic function space set up since we believe that our study will be useful in the operator theoretic aspects such as descriptions of commutants, cyclic vectors, and so on.) In Section $3$ we obtain necessary and sufficient conditions for the boundedness of $F_{w}$, and as main results, we study when $F_{w}$ and its adjoint $F_{w}^{*}$ are hypercyclic, mixing, supercyclic, and chaotic. Dynamics of the adjoint of a unilateral forward shift is studied in Section $4$. In Section $5$, it is shown that the bilateral shifts on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ and also the unilateral shifts have distinguishing properties compared to classical weighted shifts . For example, among other results, it is shown that there are unilateral and bilateral forward shifts $F$ such that $F^{*}$ has a non-zero periodic vector, without being hypercyclic.
2. The spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$: evaluation and co-ordinate functionals
We first recall the notion of equivalent Schauder bases in Banach spaces, cf. Lindenstrauss and Tzafriri [28], page $5$. Let $X$ and $Y$ be Banach spaces. We say that the Schauder bases $\{u_{n}\}_{n=-\infty}^{\infty}$ of $X$ and $\{v_{n}\}_{n=-\infty}^{\infty}$ of $Y$, are equivalent if the convergence of $\sum_{n=-\infty}^{\infty}a_{n}u_{n}$ in $X$ is equivalent to that of $\sum_{n=-\infty}^{\infty}a_{n}v_{n}$ in $Y$.
Let $a:=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b:=\{b_{n}\}_{n=-\infty}^{\infty}$ be complex sequences, $a_{n}\neq 0$, and $b_{n}\neq 0$, for all $n\in\mathbb{Z}$. Define $\ell^{p}_{a,b}(\mathtt{A})$ to be the space of all analytic functions on an annulus $\mathtt{A}$, having a normalized Schauder basis $\{f_{n}\}_{n=-\infty}^{\infty}$, where
$f_{n}(z):=(a_{n}+b_{n}z)z^{n},\hskip 28.45274pt(n\in\mathbb{Z}),$
and it is assumed to be equivalent to the standard ordered basis in $\ell^{p}(\mathbb{Z})$, $1\leq p<\infty$, that is,
$f=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}\in\ell^{p}_{a,b}(\mathtt{A}),$ (2.1)
if and only if $\lVert f\rVert_{\ell^{p}_{a,b}}:=\left(\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}\rvert^{p}\right)^{\frac{1}{p}}<\infty.$ Note that for $b_{n}=0$ for all $n\in\mathbb{Z},$ we get a space that is equivalent to the standard weighted $\ell^{p}(\mathbb{Z})$. In a similar way, we define the space $c_{0,a,b}(\mathtt{A})$ which consists of analytic functions, for which $\{(a_{n}+b_{n}z)z^{n}:n\in\mathbb{Z}\}$ forms a normalized Schauder basis, equivalent to the standard basis in $c_{0}(\mathbb{Z})$. Moreover, for $f=\sum_{n}\lambda_{n}f_{n}$ in $c_{0,a,b}(\mathtt{A})$, we have $\lambda_{n}\rightarrow 0$, as $|n|\rightarrow\infty$. Define its norm as
$\|f\|_{c_{0,a,b}(\mathtt{A})}:=\sup_{n\in\mathbb{Z}}|\lambda_{n}|.$
We can similarly consider the spaces $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ over the unit disc $\mathbb{D}$ of the complex plane.
The following theorem shows the annulus $\mathtt{A}$ occurs naturally, which depends on $p$, $\{a_{n}\}$ and $\{b_{n}\}$. The series representation of $f$ as in (2.1) can be rearranged to a convergent Laurent series, and the evaluation functionals are continuous on $\mathtt{A}$. Moreover, each function in $\ell^{p}_{a,b}(\mathtt{A})$ as well as $c_{0,a,b}(\mathtt{A})$, is analytic on the natural annulus $\mathtt{A}$.
Theorem 2.1.
Consider the spaces $\ell^{p}_{a.b}(\mathtt{A}),$ $1\leq p<\infty,$ and $c_{0,a,b}(\mathtt{A})$, defined as above corresponding to $f_{n}(z)=(a_{n}+b_{n}z)z^{n}$, $n\in\mathbb{Z}$. Set
$\mathtt{A}_{1}:=\left\{\zeta\in\mathbb{C}:\sum_{n=-\infty}^{\infty}\big{(}(\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert\zeta\rvert^{n}\big{)}^{q}<\infty\right\},$
where $\frac{1}{p}+\frac{1}{q}=1,$
$\mathtt{A}_{2}:=\left\{\zeta\in\mathbb{C}:\sup_{n\in\mathbb{Z}}~{}(\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert\zeta\rvert^{n}<\infty\right\},$
and
$\mathtt{A}_{3}:=\left\{\zeta\in\mathbb{C}:\sum_{n=-\infty}^{\infty}(\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert\zeta\rvert^{n}<\infty\right\}.$
The following hold.
  1. (i)
    If $f=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}\in\ell^{p}_{a,b}(\mathtt{A}),$ $1<p<\infty,$ then the series $\sum_{n=-\infty}^{\infty}\lambda_{n}(a_{n}+b_{n}z)z^{n}$ converges uniformly and absolutely on compact subsets of $\mathtt{A}_{1}$.
    The same conclusions hold true with respect to the domains $\mathtt{A}_{2}$ and $\mathtt{A}_{3}$, respectively, for the spaces $\ell^{1}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$.
  2. (ii)
    The evaluation functionals $ev_{\zeta}$ given by $f\mapsto f(\zeta)$ are bounded on $\ell^{p}_{a.b}(\mathtt{A}),$ and $c_{0,a,b}(\mathtt{A})$, where $1\leq p<\infty$, and $\zeta\in\mathtt{A}$.
Proof.
We prove the result for the case $1<p<\infty$ only. Assume that $f=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}$ converges in $\ell^{p}_{a,b}(\mathtt{A}),$ and $\zeta\in\mathtt{A}_{1}.$ Fix a closed ball $K$ that lies in $\mathtt{A}_{1}$. Let $\zeta\in K$. We now get
$\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}(a_{n}+b_{n}\zeta)\zeta^{n}\rvert$$\leq$ $\text{max}\{1,\lvert\zeta\rvert\}\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}\rvert(\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert\zeta\rvert^{n}$
$\leq$ $\max_{\zeta\in K}\{1,\lvert\zeta\rvert\}\left(\sum_{n=-\infty}^{\infty}(\lvert a_{n}\rvert+\lvert b_{n}\rvert)^{q}\lvert\zeta\rvert^{qn}\right)^{\frac{1}{q}}\left(\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}\rvert^{p}\right)^{\frac{1}{p}}$
$=$ $M\lVert f\rVert,$
[…]
Further, we have that $|f(\zeta)|\leq M\|f\|$, and hence, evaluation functionals are continuous. (In fact, we have actually established that $\sup_{\zeta\in K}|f(\zeta)|\leq M_{K}\|f\|$ for every compact subset $K$ in $\mathtt{A}$, where the constant $M_{K}$ depends only on $K$.) ∎
A few comments are in order. The natural domains $\mathtt{A}_{j}$, $j=1,2,3$, will always be assumed to be non-degenerate. Secondly, in the case of the usual weighted spaces, the natural domains of the map $z\mapsto ev_{z}$ are open or closed annuli. However, this is not so for $\ell^{p}_{a,b}(\mathtt{A})$ which is an important property that distinguishes these spaces from the classical weighted $\ell^{p}$ spaces, which will be discussed and utilized in Section $5$. Also, the series $\sum_{n=-\infty}^{\infty}\lambda_{n}(a_{n}+b_{n}z)z^{n}$ can still converge at a point $z$ that lies outside $A_{j}$, and hence the domain of $z\mapsto ev_{z}$ can be extended to “extra points” outside $\mathtt{A_{j}}$, cf. Section $5$. We note that such a phenomenon in the unilateral case when $p=2$, was observed by Adams and McGuire [1] who introduced the space the analytic reproducing kernel Hilbert space $\ell^{2}_{a,b}$ having an orthonormal basis of the form $(a_{n}+b_{n}z)z^{n}$, $n\geq 0$.
Since the evaluation functional $ev_{z}$ is continuous for all $z$, we now get the continuity of the co-ordinate functionals which are defined as follows. Since each $f(z)$ in $\ell^{p}_{a,b}(\mathtt{A})$ is analytic on the annulus $\mathtt{A}_{1}$, we have the Laurent expansion $f(z)=\sum_{n=-\infty}^{\infty}\widehat{f}(n)z^{n}$, $z\in\mathtt{A}_{1}$, where $\widehat{f}(n)$ is the $n$-th Laurent coefficient of $f(z)$. The $n$-th co-ordinate functional (or rather, a coefficient functional) $k_{n}$ is given by
$k_{n}(f)=\widehat{f}(n),$
and it is defined on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. Also, let $f_{n}^{*}$ denote co-ordinate functional with respect to the basis $\{f_{n}\}$ in $\ell^{p}_{a,b}(\mathtt{A})$, $1\leq p<\infty$, or $c_{0,a,b}(\mathtt{A})$, that is,
$f_{n}^{*}(f)=\lambda_{n},\hskip 17.07182ptf=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}.$
(We add here that, both $k_{n}$ and $f_{n}$ will be utilized in understanding the dynamics of a weighted forward shift $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$. Indeed, the adjoint $F_{w}^{*}$ will “act like a weighted shift” with respect to $k_{n}$, see Proposition 3.1.)
Proposition 2.2.
Let $k_{n}$ denote the co-ordinate functional
$f\mapsto\widehat{f}(n),$
defined on $\ell^{p}_{a,b}(\mathtt{A})$ or $c_{0,a,b}(\mathtt{A})$, where $n\in\mathbb{Z}$. Then the following hold.
  • (i)
    For $1<p<\infty$, $11/p+1/q=$ 1,
    $\|k_{n}\|\leq(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q}.$
    For $p=1$, we have $\|k_{n}\|\leq\max\{|a_{n}|,|b_{n-1}|\}$. Also, for the case of $c_{0,a,b}(\mathtt{A})$, $\|k_{n}\|\leq|a_{n}|+|b_{n-1}|$.
  • (ii)
    $\{f_{n},f_{n}^{*}\}$ is a biorthogonal system, and for $1<p<\infty$, $\{f_{n}^{*}\}$ is an unconditional Schauder basis for the dual space.
  • (iii)
    For $n\in\mathbb{Z}$, we have $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$.
Proof.
$(i)$ If $f\in\ell^{p}_{a,b}(\mathtt{A}),$ then $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}(a_{k}+b_{k}z)z^{k}$, and $\|f\|=\left(\sum_{k=-\infty}^{\infty}\lvert\lambda_{k}\rvert^{p}\right)^{1/p}$. Rearrangement into a Laurent series yields $f(z)=\sum_{k=-\infty}^{\infty}(\lambda_{k}a_{k}+\lambda_{k-1}b_{k-1})z^{k},$ and so, $k_{n}(f)=\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}.$ It follows, by Hölder inequality, that $|k_{n}(f)|\leq\|f\|_{\ell^{p}_{a,b}(\mathtt{A})}(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q}$, and hence $\|k_{n}\|\leq(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q}$.
$(ii)$ This is immediate as $f_{n}^{*}(f_{n})=1$, and $f_{n}^{*}(f_{m})=0$ for $n\neq m$. Also, the equivalence of $\{f_{n}\}$ to the standard basis $\{e_{n}\}$ shows that $\{f_{n}^{*}\}$ and $\{e_{n}^{*}\}$ are equivalent bases which follow from reflexivity.
$(iii)$ For $f\in\ell^{p}_{a,b}(\mathtt{A}),$ we have
$(a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*})(f)$ $=$ $a_{n}f_{n}^{*}(f)+b_{n-1}f_{n-1}^{*}(f)$
$=$ $a_{n}f_{n}^{*}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}\big{)}+b_{n-1}f_{n-1}^{*}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}\big{)}$
$=$ $\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}=k_{n}(f).$
Hence, $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\in\mathbb{Z}$. ∎
We recall that the unilateral versions $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ were introduced studied in [15]. These spaces correspond to the normalized bases $(a_{n}+b_{n}z)z^{n}$, where $n\geq 0$. The results in Proposition 2.2 hold for $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ as well.
The following is a slight generalization on the continuity of the co-ordinate functionals $k_{n}$ defined on analytic function spaces over an annulus and disc centered at the origin.
Proposition 2.3.
Let $\mathcal{E}$ be a Banach space of analytic functions on an annulus $\mathtt{A}$ or a disc centered around 00, having evaluation functionals bounded at each $z\in\mathtt{A}$. Then, $k_{n}$ is bounded for all $n\in\mathbb{Z}$. Further, $\text{span}~{}\big{\{}k_{n}:~{}n\in\mathbb{Z}\big{\}}$ is dense in $\mathcal{E}^{*}$ with respect to the weak-topology, and also, in the norm topology of $\mathcal{E}^{*}$ if it is reflexive.
Proof.
Let $ev_{z}$ denote the evaluation functional acting on $\mathcal{E}$. Since the analyticity of $z\mapsto ev_{z}$ in the strong and norm operator topologies are equivalent, we see that $z\mapsto ev_{z}$ is an $\mathcal{E}^{*}$-valued norm-analytic function. It then admits a norm-convergent Laurent series form $ev_{z}=\sum_{n=-\infty}^{\infty}L_{n}z^{n},\hskip 14.22636ptz\in\mathtt{A}$, for some $L_{n}\in\mathcal{E}^{*}$. For $f\in\mathcal{E}$, we have $f(z)=ev_{z}(f)=\sum_{n=-\infty}^{\infty}L_{n}(f)z^{n}.$ Expanding $f(z)$ in its Laurent series, we must then have $L_{n}=k_{n}$ for all $n$.
To show that $\{k_{n}\}$ is a weak-total set in $\mathcal{E}^{*}$, let $\Lambda\in\mathcal{E}^{**}$ be such that $\Lambda(k_{n})=0$ for all $n$. Hence, there is some $f\in\mathcal{E}$ such that $\lambda(k_{n})=k_{n}(f)=0,~{}\forall n\in\mathbb{Z}$, which shows that all the coefficients of $f(z)$ vanish. Consequently, $f=0$.
We now obtain conditions so that the Laurent monomials $z^{\nu}$ belong to the space $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. The norm estimates, stated below, will be later used in the dynamics of weighted forward shifts.
Proposition 2.4.
Consider the spaces $\ell^{p}_{{a},{b}}(\mathtt{A}),$ $1\leq p<\infty$, and $c_{0,a,b}(\mathtt{A})$. Assume that
$\limsup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}\limsup_{n\rightarrow+\infty}\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<1,$
then there exist constants $M_{1},M_{2}>0$ such that
$\lVert z^{\nu}\rVert\leq M_{1}\left\lvert\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\right\rvert~{}~{}~{}\text{and}~{}~{}~{}\lVert z^{-\nu}\rVert\leq M_{2}~{}~{}\left\lvert\frac{b_{-1}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}\right\rvert,$ (2.2)
for all $\nu\in\mathbb{N}.$ The same holds for $c_{0,a,b}(\mathtt{A})$ as well. In particular, the Laurent polynomials form a dense subspace in $c_{0,a,b}(\mathtt{A})$ and $\ell^{p}_{{a},{b}}(\mathtt{A})$.
Proof.
By the basis expansion in $\ell^{p}_{{a},{b}}(\mathtt{A})$ and a rearrangement into Laurent series, we can find some $\{\lambda_{j}\}_{j=-\infty}^{\infty}\in\ell^{p}(\mathbb{Z})$ such that
$1=\sum_{j=-\infty}^{\infty}\lambda_{j}(a_{j}z^{j}+b_{j}z^{j+1})=\sum_{j=-\infty}^{\infty}\left(\lambda_{j}a_{j}+\lambda_{j-1}b_{j-1}\right)z^{j},$ (2.3)
for all $z$ in an annulus around 00. Now equating the coefficients of like-powers, we have
$\lambda_{j}=\begin{cases}(-1)^{j}\lambda_{0}\frac{b_{0}b_{1}b_{2}\cdots b_{j-1}}{a_{1}a_{2}a_{3}\cdots a_{j}},\hskip 88.2037ptj\geq 1\\ \lambda_{0},\hskip 165.02606ptj=0,\\ \frac{1}{b_{-1}}\left(1-\lambda_{0}a_{0}\right),\hskip 105.2751ptj=-1,\\ (-1)^{-j-1}\frac{1}{b_{-1}}\left(1-\lambda_{0}a_{0}\right)\frac{a_{-1}a_{-2}\cdots a_{j+1}}{b_{-2}b_{-3}\cdots b_{j}},~{}~{}j\leq-2.\\ \end{cases}$
Then, for a fixed $\nu\geq 2,$
$a_{0}$ $=$ $f_{0}-b_{0}z$
$=$ $f_{0}-\frac{b_{0}}{a_{1}}f_{1}+\frac{b_{0}b_{1}}{a_{1}}z^{2}$
$=$ $f_{0}-\frac{b_{0}}{a_{1}}f_{1}+\frac{b_{0}b_{1}}{a_{1}}f_{2}+\cdots+(-1)^{\nu-1}\frac{b_{0}b_{1}\cdots b_{\nu-2}}{a_{1}a_{2}\cdots a_{\nu-1}}f_{\nu-1}+(-1)^{\nu}\frac{b_{0}b_{1}\cdots b_{\nu-1}}{a_{1}a_{2}\cdots a_{\nu-1}}z^{\nu}.$
It follows that
$z^{\nu}=(-1)^{\nu}\left(\frac{a_{0}a_{1}\cdots a_{\nu-1}}{b_{0}b_{1}\cdots b_{\nu-1}}\right)-(-1)^{\nu}\left(\frac{a_{1}a_{2}\cdots a_{\nu-1}}{b_{0}a_{1}\cdots b_{\nu-1}}f_{0}-\cdots+(-1)^{\nu-1}\frac{1}{b_{\nu-1}}f_{\nu-1}\right).$
Again we can write
$z^{\nu}=(-1)^{\nu}\left(\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\right)\sum_{j=-\infty}^{\infty}\lambda_{j}f_{j}-(-1)^{\nu}\left(\frac{a_{1}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}f_{0}-\cdots+(-1)^{\nu-1}\frac{1}{b_{\nu-1}}f_{\nu-1}\right).$
Finally, we get
$z^{\nu}$ $=$ $(-1)^{\nu}\left(\frac{1-\lambda_{0}a_{0}}{b_{-1}}\right)\left(\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\right)\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{\nu}\left(\lambda_{0}a_{0}-1\right)\left(\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\right)\left(\frac{1}{a_{0}}f_{0}-\frac{b_{0}}{a_{0}a_{1}}f_{1}+\cdots+(-1)^{\nu-1}\frac{b_{0}\cdots b_{\nu-2}}{a_{0}\cdots a_{\nu-1}}f_{\nu-1}\right)$
$+\frac{\lambda_{0}a_{0}}{a_{\nu}}\left(f_{\nu}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{\nu}b_{\nu+1}\cdots b_{\nu+j-1}}{a_{\nu+1}a_{\nu+2}\cdots a_{\nu+j}}f_{\nu+j}\right).$
Now, since $\lim\sup_{n\rightarrow+\infty}|b_{n}/a_{n+1}|<1,$ we can find $r<1$ and an integer $N\geq 1,$ such that $\big{|}\frac{b_{n}}{a_{n+1}}\big{|}<r,$ for all $n\geq N.$ Thus, there exists a constant $M_{1}>0$ such that
$\left\lVert z^{j}\right\rVert_{\ell^{p}_{{a},{b}}(\mathtt{A})}\leq M_{1}\left\lvert\frac{a_{0}a_{1}\cdots a_{j-1}}{b_{0}b_{1}\cdots b_{j-1}}\right\rvert,$
for all $j\geq N$. This proves the first part.
For $\nu\leq-1,$ we write
$b_{-1}=f_{-1}-a_{-1}z^{-1},$
and proceed as above using $\lim\sup_{n\rightarrow+\infty}|a_{-n+1}/b_{-n}|<1.$ Then, we can obtain that
$\left\{\frac{a_{-1}a_{-2}\cdots a_{-\nu}}{b_{-1}b_{-2}\cdots b_{-\nu}}z^{-\nu}:\nu\geq 1\right\}$
is norm-bounded in $\ell^{p}_{a,b}(\mathtt{A}),$ and hence the second estimate in the proposition holds. ∎
3. Bilateral weighted shifts $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$
In this section we introduce the bilateral forward shift operator $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. Necessary and sufficient conditions for $F_{w}$ to become a bounded operator are obtained in terms of $\{a_{n}\}$, $\{b_{n}\}$ and the weight sequence $\{w_{n}\}$. It is shown that operator $F_{w}$, under suitable assumptions, is similar to a compact perturbation of a classical bilateral weighted shift on $\ell^{p}(\mathbb{Z})$. Hypercyclicity, supercyclicity, and chaos for both $F_{w}$ and $F_{w}^{*}$ are studied. Illustrations using examples are provided.
Let $w:=\{w_{n}\}$ be a complex sequence. The operator $F_{w}$ is given by the action $F_{w}(z^{n})=w_{n}z^{n+1}$ for all $n\in\mathbb{Z}$, that is, $F_{w}(\sum_{n=-\infty}^{\infty}\lambda_{n}z^{n})=\sum_{n=-\infty}^{\infty}\lambda_{n}w_{n}z^{n+1}$. We first isolate certain properties of $F_{w}$ which are of general interest. Among them is an important property of weighted forward shifts, which loosely says that the adjoint $F_{w}^{*}$ whenever $F_{w}$ is bounded, acts like a “weighted backward shift” with respect to a total set in the dual of a Banach space of analytic functions. The unweighted case of $F$ defined on reproducing kernel space over the unit disc was discussed in [33].
Proposition 3.1.
Let $\mathcal{E}$ be a Banach space of analytic functions defined on some annulus $r<|z|<R$, having continuous evaluation functionals. Let $F_{w}$ be bounded on $\mathcal{E}$. Then,
$F_{w}^{*}(k_{n})=w_{n-1}k_{n-1},$
for all $n\in\mathbb{Z},$ where $k_{n}$ denotes the $n$-th co-ordinate functional.
Proof.
Note that $k_{n}\in\mathcal{E}^{*}$ for all $n$. For a function $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}z^{k}$ in $\mathcal{E}$, we have
$F_{w}^{*}\big{(}k_{n}\big{)}(f)$ $=$ $\big{(}k_{n}\circ F_{w}\big{)}(f)$
$=$ $k_{n}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}w_{k}z^{k+1}\big{)}$
$=$ $\lambda_{n-1}w_{n-1}.$
On the other hand, $k_{n-1}(f)=k_{n-1}\left(\sum_{k=-\infty}^{\infty}\lambda_{k}z^{k}\right)=\lambda_{n-1}.$ Hence, $F_{w}^{*}(k_{n})=w_{n-1}k_{n-1},$ for all $n\in\mathbb{Z}.$
An analogue for the unilateral weighted forward shift $F_{w}$ is stated below, using the same notation $k_{n}$ for $n$-th co-ordinate functional, $n\geq 0$. We omit its proof.
Proposition 3.2.
Let $F_{w}$ be a bounded forward weighted shift on a Banach space $\mathcal{E}$ consisting of analytic functions on the unit disc $\mathbb{D}$. Then
$F_{w}^{*}(k_{n})=\begin{cases}w_{n-1}k_{n-1},\hskip 14.22636ptn\geq 1,\\ 0,\hskip 56.9055ptn=0.\end{cases}$
In terms of $k_{n}$, a general necessary condition for $F_{w}$ to be bounded can be obtained:
Proposition 3.3.
Suppose $\mathcal{E}$ is a Banach space of analytic functions defined on an annulus $\mathtt{A}$, having continuous evaluation functionals. Suppose that $k_{n}\neq 0$ for all $n\in\mathbb{Z}$. If $F_{w}$ is a bounded operator on $\mathcal{E}$, then $\sup~{}\Big{\{}w_{n-1}\frac{\|k_{n-1}\|}{\|k_{n}\|}:~{}{n\in\mathbb{Z}}\Big{\}}<\infty.$
Proof.
Consider the vectors of unit norm, $u_{n}:=\frac{k_{n}}{\|k_{n}\|},~{}n\in\mathbb{Z}$. Since $F_{w}$ is a bounded operator, we get that $\{F_{w}^{*}(u_{n}):n\in\mathbb{Z}\}$ is a norm-bounded set in $\mathcal{E}^{*}$. This, along with the previous proposition, implies the result. ∎
3.1. Matrix representation of $F_{w}$, boundedness and a perturbation result
To get the boundedness, we compute the matrix representations of the operator $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ with respect to the ordered Schauder basis $\{f_{n}\}_{n\in\mathbb{Z}}$. Recall that normalized Schauder basis $f_{n}(z)=(a_{n}+b_{n}z)z^{n},n\in\mathbb{Z}$ in $\ell^{p}_{a,b}(\mathtt{A})$ is assumed to be equivalent to the standard basis in $\ell^{p}(\mathbb{Z})$. Then, the boundedness of $F_{w}$ is equivalent to that of its matrix operator acting on $\ell^{p}(\mathbb{Z})$.
Note that
$F_{w}(f_{n})(z)=w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}=\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+\left(\frac{w_{n+1}b_{n}}{a_{n+2}}-\frac{w_{n}a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right)a_{n+2}z^{n+2},$
by putting the value of $z^{n+1}$. Setting
$c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}},$
we immediately get for $n\geq-1,$
$F_{w}(f_{n})(z)$ $=$ $\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+c_{n}a_{n+2}z^{n+2}$
$=$ $\left(\frac{w_{n}a_{n}}{a_{n+1}}+\alpha c_{n}\frac{a_{n+2}}{b_{n+1}}\right)f_{{n+1}}$
$+(-1)^{n+2}\alpha c_{n}\left(\frac{a_{0}\cdots a_{n+2}}{b_{-1}\cdots b_{n+1}}\right)\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{n+3}\alpha c_{n}\left(\frac{a_{1}\cdots a_{n+2}}{b_{0}\cdots b_{n+1}}f_{0}-\frac{a_{2}\cdots a_{n+2}}{b_{1}\cdots b_{n+1}}f_{1}+\cdots+(-1)^{n}\frac{a_{n+1}a_{n+2}}{b_{n}b_{n+1}}f_{n}\right)$
$+(1-\alpha)c_{n}\left(f_{n+2}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{n+2}b_{n+3}\cdots b_{n+1+j}}{a_{n+3}a_{n+4}\cdots a_{n+2+j}}f_{n+2+j}\right).$
Again
$F_{w}(f_{n})(z)=w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}=\frac{w_{n+1}b_{n}}{b_{n+1}}f_{{n+1}}+\left(\frac{w_{n}a_{n}}{b_{n}}-\frac{w_{n+1}a_{n+1}}{b_{n+1}}\right)b_{n}z^{n+1},$
by substituting the expression of $z^{n+2}$. Setting
$d_{n}:=w_{n}\frac{a_{n}}{b_{n}}-w_{n+1}\frac{a_{n+1}}{b_{n+1}},$
we immediately get for $n\geq 2,$
$F_{w}(f_{-n})(z)$ $=$ $\frac{w_{-n+1}b_{-n}}{b_{-n+1}}f_{{-n+1}}+d_{-n}b_{-n}z^{-n+1}$
$=$ $\left(\frac{w_{-n+1}b_{-n}}{b_{-n+1}}+(1-\alpha)d_{-n}\frac{b_{-n}}{a_{-n+1}}\right)f_{-n+1}$
$+\alpha d_{-n}\left(f_{-n}+\sum_{j=0}^{\infty}(-1)^{j+1}\frac{a_{-n}a_{-n-1}\cdots a_{-n-j}}{b_{-n-1}b_{-n-2}\cdots b_{-n-j-1}}f_{-n-j-1}\right)$
$+(-1)^{n-1}(1-\alpha)d_{-n}\left(\frac{-b_{-2}\cdots b_{-n}}{a_{-1}\cdots a_{-n+1}}f_{-1}+\cdots+(-1)^{n-2}\frac{b_{-n+1}b_{-n}}{a_{-n+2}a_{-n+1}}f_{-n+2}\right)$
$+(-1)^{n-1}d_{-n}\lambda_{0}\left(\frac{b_{-1}\cdots b_{-n}}{a_{-1}\cdots a_{-n+1}}\right)\left(f_{0}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{1}a_{2}\cdots a_{j}}f_{j}\right),$
where $\alpha=1-\lambda_{0}a_{0}.$ Hence, the (bilateral) matrix $[F_{w}]$ of the operator $F_{w}$ with respect to the ordered normalized Schauder basis $\{f_{n}\}_{n\in\mathbb{Z}}$ is given by
$[F_{w}]=\begin{bmatrix}\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots\\ \ddots&\frac{\alpha d_{-2}a_{-3}a_{-2}}{b_{-4}b_{-3}}&\frac{\alpha c_{-1}a_{-3}\cdots a_{1}}{b_{-4}\cdots b_{0}}&\frac{-\alpha c_{0}a_{-3}\cdots a_{2}}{b_{-4}\cdots b_{1}}&\ddots&\ddots&\ddots\\ \ddots&\frac{-\alpha d_{-2}a_{-2}}{b_{-3}}&\frac{-\alpha c_{-1}a_{-2}\cdots a_{1}}{b_{-3}\cdots b_{0}}&\frac{\alpha c_{0}a_{-2}\cdots a_{2}}{b_{-3}\cdots b_{1}}&\frac{-\alpha c_{1}a_{-2}\cdots a_{3}}{b_{-3}\cdots b_{2}}&\ddots&\ddots\\ \ddots&\alpha d_{-2}&\frac{\alpha c_{-1}a_{-1}a_{0}a_{1}}{b_{-2}b_{-1}b_{0}}&\frac{-\alpha c_{0}a_{-1}\cdots a_{2}}{b_{-2}\cdots b_{1}}&\frac{\alpha c_{1}a_{-1}\cdots a_{3}}{b_{-2}\cdots b_{2}}&\frac{-\alpha c_{2}a_{-1}\cdots a_{4}}{b_{-2}\cdots b_{3}}&\ddots\\ \ddots&\gamma_{-2}&\frac{-\alpha c_{-1}a_{0}a_{1}}{b_{-1}b_{0}}&\frac{\alpha c_{0}a_{0}a_{1}a_{2}}{b_{-1}b_{0}b_{1}}&\frac{-\alpha c_{1}a_{0}\cdots a_{3}}{b_{-1}\cdots b_{2}}&\frac{\alpha c_{2}a_{0}a_{1}\cdots a_{4}}{b_{-1}\cdots b_{3}}&\ddots\\ \ddots&\frac{-\lambda_{0}d_{-2}b_{-1}b_{-2}}{a_{-1}}&\gamma_{-1}&\bm{\frac{-\alpha c_{0}a_{1}a_{2}}{b_{0}b_{1}}}&\frac{\alpha c_{1}a_{1}a_{2}a_{3}}{b_{0}b_{1}b_{2}}&\frac{-\alpha c_{2}a_{1}\cdots a_{4}}{b_{0}\cdots b_{3}}&\ddots\\ \ddots&\frac{\lambda_{0}d_{-2}b_{-2}b_{-1}b_{0}}{a_{-1}a_{1}}&(1-\alpha)c_{-1}&\gamma_{0}&\frac{-\alpha c_{1}a_{2}a_{3}}{b_{1}b_{2}}&\frac{\alpha c_{2}a_{2}a_{3}a_{4}}{b_{1}b_{2}b_{3}}&\ddots\\ \ddots&\frac{-\lambda_{0}d_{-2}b_{-2}\cdots b_{1}}{a_{-1}a_{1}a_{2}}&\frac{-(1-\alpha)c_{-1}b_{1}}{a_{2}}&(1-\alpha)c_{0}&\gamma_{1}&\frac{-\alpha c_{2}a_{3}a_{4}}{b_{2}b_{3}}&\ddots\\ \ddots&\ddots&\frac{(1-\alpha)c_{-1}b_{1}b_{2}}{a_{2}a_{3}}&\frac{-(1-\alpha)c_{0}b_{2}}{a_{3}}&(1-\alpha)c_{1}&\gamma_{2}&\ddots\\ \ddots&\ddots&\ddots&\frac{(1-\alpha)c_{0}b_{2}b_{3}}{a_{3}a_{4}}&\frac{-(1-\alpha)c_{1}b_{3}}{a_{4}}&(1-\alpha)c_{2}&\ddots\\ \ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots\end{bmatrix}.$
Here, the expression in bold refers to the $(0,0)$-th entry of the above matrix. Moreover, $\alpha=1-\lambda_{0}a_{0},$ and
[…]
We can now state certain sufficient conditions for $F_{w}$ to become a bounded operator.
Theorem 3.4.
If
$\limsup_{n\rightarrow+\infty}\left\{\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert,\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert,\left\lvert c_{n}\frac{a_{0}a_{1}\cdots a_{n+2}}{b_{-1}b_{0}\cdots b_{n+1}}\right\rvert,\left\lvert d_{-n}\frac{b_{-1}b_{-2}\cdots b_{-n}}{a_{-1}a_{-2}\cdots a_{-n+1}}\right\rvert\right\}<1,$
and
$\sup_{n\geq 0}~{}\left\{\left\lvert w_{n}\frac{a_{n}}{a_{n+1}}\right\rvert,\left\lvert w_{-n}\frac{b_{-n-1}}{b_{-n}}\right\rvert\right\}<\infty,$
then $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A})$.
Proof.
It suffices to show that the above infinite matrix acts as a bounded operator on $\ell^{p}(\mathbb{Z})$. To this end, write the matrix of $F_{w}$ as a formal series of infinite matrices as follows:
$[F_{w}]=\cdots+[E_{-2}]+[E_{-1}]+[D]+[E_{1}]+[E_{2}]+\cdots,$
where $[D]$ is the matrix of the diagonal operator
$\text{diag}~{}(\cdots,\alpha d_{-3},\alpha d_{-2},\alpha\frac{c_{-1}a_{0}a_{1}}{b_{-1}b_{0}},\bm{\alpha\frac{c_{0}a_{1}a_{2}}{b_{0}b_{1}}},\alpha\frac{c_{1}a_{2}a_{3}}{b_{1}b_{2}},\alpha\frac{c_{2}a_{3}a_{4}}{b_{2}b_{3}},\alpha\frac{c_{3}a_{4}a_{5}}{b_{3}b_{4}},\cdots)$
on $\ell^{p}(\mathbb{Z}).$ Also the matrix $[E_{n}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $n$-th subdiagonal, where $n\geq 1$. Observe that $[E_{n}]$ is the matrix of suitable powers of a bilateral weighted forward shift $E_{n}$ for $n\geq 1$. Again, the matrix $[E_{-n}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $n$-th superdiagonal, where $n\geq 1$. Observe that $[E_{-n}]$ is the matrix of suitable powers of a bilateral weighted backward shift $E_{-n}$ for $n\geq 1$.
Note that $[E_{1}]$ is the matrix of the standard weighted bilateral forward shift $E_{1}(e_{n})\mapsto\gamma_{n}e_{n+1}$ on $\ell^{p}(\mathbb{Z})$, $n\in\mathbb{Z}$, having weights
[…]
For each $n\geq 0,$ we have clearly
$c_{n}=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}=\frac{b_{n}}{a_{n+1}}\frac{w_{n+1}a_{n+1}}{a_{n+2}}-\frac{w_{n}a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}},$
and for $n<0$
$d_{n}=w_{n}\frac{a_{n}}{b_{n}}-w_{n+1}\frac{a_{n+1}}{b_{n+1}}=\frac{a_{n}}{b_{n-1}}\frac{w_{n}b_{n-1}}{b_{n}}-\frac{a_{n+1}}{b_{n}}\frac{w_{n+1}b_{n}}{b_{n+1}}.$
By the assumptions, it follows that $\{c_{n}\}$ and $\{d_{n}\}$ is a bounded sequence. Now
$\lVert D\rVert$$\leq$ $\alpha\sup_{n\geq 0}\left\{\left\lvert c_{n}\frac{a_{n+1}a_{n+2}}{b_{n}b_{n+1}}\right\rvert,|d_{-n-2}|\right\},$
$\sum_{i=1}^{\infty}\lVert E_{i}\rVert$$\leq$ $\sup_{n\geq 0}\left\{\left\lvert w_{n}\frac{a_{n}}{a_{n+1}}+\alpha\frac{c_{n}a_{n+2}}{b_{n+1}}\right\rvert,\left\lvert w_{-n-1}\frac{b_{-n-1}}{b_{-n}}+(1-\alpha)\frac{d_{-n-2}b_{-n-2}}{a_{-n-1}}\right\rvert\right\}$
$+(1-\alpha)\sup_{n\geq 0}\left\{\left\lvert c_{n}\right\rvert,\left\lvert\frac{d_{-n-3}b_{-n-2}b_{-n-3}}{a_{-n-1}a_{-n-1}}\right\rvert\right\}$
$+(1-\alpha)\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n}b_{n+2}}{a_{n+3}}\right\rvert,\left\lvert\frac{d_{-n-4}b_{-n-2}b_{-n-3}b_{-n-4}}{a_{-n-1}a_{-n-2}a_{-n-3}}\right\rvert\right\}$
$+(1-\alpha)\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n}b_{n+2}b_{n+3}}{a_{n+3}a_{n+4}}\right\rvert,\left\lvert\frac{d_{-n-5}b_{-n-2}b_{-n-3}b_{-n-4}b_{-n-5}}{a_{-n-1}a_{-n-2}a_{-n-3}a_{-n-4}}\right\rvert\right\}+\cdots,$
and
$\sum_{i=1}^{\infty}\lVert E_{-i}\rVert$$\leq$ $\alpha\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n}a_{n}a_{n+1}a_{n+2}}{b_{n-1}b_{n}b_{n+1}}\right\rvert,\left\lvert d_{-n-2}\frac{a_{-n-2}}{b_{-n-3}}\right\rvert\right\}$
$+\alpha\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n+1}a_{n}a_{n+1}a_{n+2}a_{n+3}}{b_{n-1}b_{n}b_{n+1}b_{n+2}}\right\rvert,\left\lvert d_{-n-2}\frac{a_{-n-2}a_{-n-3}}{b_{-n-2}b_{-n-3}}\right\rvert\right\}$
$+\alpha\sup_{n\geq 0}\left\{\left\lvert\frac{c_{n+2}a_{n}a_{n+1}a_{n+2}a_{n+3}a_{n+4}}{b_{n-1}b_{n}b_{n+1}b_{n+2}b_{n+3}}\right\rvert,\left\lvert d_{-n-2}\frac{a_{-n-2}a_{-n-3}a_{-n-4}}{b_{-n-3}b_{-n-4}b_{-n-5}}\right\rvert\right\}+\cdots.$
Moreover, our assumptions
$\limsup_{n\rightarrow+\infty}\left\{\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert,\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert,\left\lvert c_{n}\frac{a_{0}a_{1}\cdots a_{n+2}}{b_{0}b_{1}\cdots b_{n+1}}\right\rvert,\left\lvert d_{-n}\frac{b_{-1}b_{-2}\cdots b_{-n}}{a_{-1}a_{-2}\cdots a_{-n+1}}\right\rvert\right\}<1$
implies that there exists $r,R,s,S<1$ and $n_{0}\in\mathbb{N}$ such that
[…]
for $n\geq n_{0},$ from which it follows that $\sum_{i=1}^{\infty}\lVert E_{-i}\rVert+\lVert D\rVert+\sum_{i=1}^{\infty}\lVert E_{i}\rVert<\infty$ and this completes the proof of the theorem. ∎
A necessary condition for $F_{w}$ to be bounded on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ is stated below.
Proposition 3.5.
If $F_{w}$ is a bounded operator on any of the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$, then $\{\gamma_{n}\}_{n=-\infty}^{\infty}$ is bounded, where $\gamma_{n}$ is the sequence as in (3.1).
Proof.
If $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A})$, then its matrix $[F_{w}]$ acts as a bounded operator on the space $\ell^{p}(\mathbb{Z})$. Hence,
$\sup_{n\in\mathbb{Z}}||[F_{w}](e_{n})||_{\ell^{p}(\mathbb{Z})}<\infty,$
where $e_{n}$ is the standard unit vector in $\ell^{p}(\mathbb{Z})$. On the other hand, $[F_{w}](e_{n})$ is the $n$-th column of the bilateral matrix $[F_{w}]$. With the observation that
$|\gamma_{n}|\leq\|[F_{w}](e_{n})\|_{\ell^{p}(\mathbb{Z})},\hskip 5.69046ptn\in\mathbb{Z},$
we immediately get $\sup_{n}|\gamma_{n}|<\infty$. ∎
The matrix representation of $F_{w}$ hints that the operator $F_{w}$ is far from being similar to a usual weighted shift on $\ell^{p}(\mathbb{Z})$ in general. Indeed, we can produce a weight $w$ and some $a=\{a_{n}\}$ and $b=\{b_{n}\}$ such that the $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ is not similar any forward weighted shift on $\ell^{p}(\mathbb{Z})$, see Section $5$. Below, we provide sufficient conditions so that the operator becomes a compact perturbation of a usual weighted shift.
Theorem 3.6.
Let $F_{w}$ be bounded on $\ell^{p}_{{a},{b}}(\mathtt{A})$. If
$\lim_{n\rightarrow\infty}\left\lvert w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right\rvert=0~{}\text{and}~{}\lim_{n\rightarrow-\infty}\left\lvert w_{n}\frac{a_{n}}{b_{n}}-w_{n+1}\frac{a_{n+1}}{b_{n+1}}\right\rvert=0,$
then the operator $F_{w}$ on $\ell^{p}_{{a},{b}}(\mathtt{A})$ is similar to $E_{\alpha}+K$ for some compact operator $K$ and a bilateral weighted forward shift $E_{\alpha}$ acting on the sequence space $\ell^{p}(\mathbb{Z})$, where the weight sequence $\alpha=(\alpha_{n})$ is given by
$\alpha_{n}=\begin{cases}w_{n}\frac{a_{n}}{a_{n+1}},\hskip 28.45274ptn\geq-1\\ w_{n+1}\frac{b_{n}}{b_{n+1}},\hskip 19.91684ptn\leq-2.\end{cases}$
Proof.
This result follows from the above matrix representation of $F_{w}$. Consider the similarity operator $V:\ell^{p}_{a,b}(\mathtt{A})\rightarrow\ell^{p}(\mathbb{Z})$ given by $V\big{(}\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}\big{)}=\sum_{n=-\infty}^{\infty}\lambda_{n}e_{n}$, where $\{e_{n}\}_{n\in\mathbb{Z}}$ is the standard basis in $\ell^{p}(\mathbb{Z})$. Now, from the proof of Theorem 3.4 we understand that $F_{w}$ acting on $\ell^{p}_{a,b}\mathtt{(}A)$ is similar via $V$ to the operator given by the series (which is convergent absolutely with respect to the operator norm)
$\sum_{i=1}^{\infty}E_{-i}+D+E_{1}+\sum_{i=2}^{\infty}E_{i}.$
Further, by our assumptions, the operators $E_{-i},$ $D$ and $E_{i}$ are compact on $\ell^{p}(\mathbb{Z})$ as the entries in each of the matrices $E_{-i},$ $D$ and $E_{i}$ converge to 00 except for the case $i=1$. The later case corresponds to $E_{1}$ which is the weighted forward shift on $\ell^{p}(\mathbb{Z})$ with weights $\alpha=\{\alpha_{n}\}_{n=-\infty}^{\infty}$, as in the theorem. We take $E_{\alpha}:=E_{1}$, and $K:=\sum_{i=1}^{\infty}E_{-i}+D+\sum_{i=2}^{\infty}E_{i}$ which is a compact operator on $\ell^{p}(\mathbb{Z})$. This proves the theorem. ∎
3.2. Linear dynamics of $F_{w}$ and its adjoint $F_{w}^{*}$
In this subsection, we obtain necessary and sufficient conditions for $F_{w}$ and its adjoint operator $F_{w}^{*}$ to be hypercyclic, mixing, supercyclic, or chaotic on $\ell^{p}_{a,b}(\mathtt{A})$, $c_{0,a,b}(\mathtt{A})$ and their dual spaces, respectively.
Theorem 3.7.
The following hold for the bilateral weighted forward shift $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$, assuming that the conditions of Theorem 3.4 are satisfied:
  1. (i)
    $F_{w}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})$ if
    $\limsup_{n\rightarrow\infty}\left\lvert\frac{1}{w_{\nu}\cdots w_{\nu+n-1}}\frac{b_{0}\cdots b_{\nu+n-1}}{a_{0}\cdots a_{\nu+n-1}}\right\rvert=\infty=\limsup_{n\rightarrow\infty}\left\lvert w_{\nu-1}\cdots w_{\nu-n}\frac{a_{-1}\cdots a_{\nu-n}}{b_{-1}\cdots b_{\nu-n}}\right\rvert,$ (3.2)
    ${}~{}\forall~{}~{}\nu\in\mathbb{N}.$
  2. (ii)
    $F_{w}$ is topologically mixing on $\ell^{p}_{a,b}(\mathtt{A})$ if
    $\limsup_{n\rightarrow\infty}\left\lvert\frac{1}{w_{\nu}\cdots w_{\nu+n-1}}\frac{b_{0}\cdots b_{\nu+n-1}}{a_{0}\cdots a_{\nu+n-1}}\right\rvert=\infty=\limsup_{n\rightarrow\infty}\left\lvert w_{\nu-1}\cdots w_{\nu-n}\frac{a_{-1}\cdots a_{\nu-n}}{b_{-1}\cdots b_{\nu-n}}\right\rvert,$ (3.3)
    ${}~{}\forall~{}~{}\nu\in\mathbb{N}.$
  3. (iii)
    $F_{w}$ is supercyclic on $\ell^{p}_{a,b}(\mathtt{A})$ if
    $\limsup_{n\rightarrow\infty}\left\lvert\frac{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\frac{b_{0}\cdots b_{\nu+n-1}}{a_{0}\cdots a_{\nu+n-1}}\frac{a_{-1}\cdots a_{\nu-n}}{b_{-1}\cdots b_{\nu-n}}\right\rvert=\infty,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.4)
Proof.
(i) Suppose that (3.2) holds, and we will use Gethner-Shapiro criterion to show that $F_{w}$ is hypercyclic. Let $D$ be the span of $\{z^{\nu}:\nu\in\mathbb{Z}\}$. Then, $D$ is dense in $\ell^{p}_{a,b}(\mathtt{A}),$ as it contains the normalized Schauder basis vectors $(a_{n}+b_{n}z)z^{n}$. Define $S:D\rightarrow D$ by
$s(z^{\nu})=\frac{1}{w_{n-1}}z^{\nu-1},~{}~{}\forall~{}~{}\nu\in\mathbb{Z}.$
Since $F_{w}S(f)=f,$ $\forall$ $f\in D,$ we just need to show that both the sequences $F_{w}^{n_{k}}(z^{\nu})$ and $S^{n_{k}}(z^{\nu})$ tend to 0 as $k\to\infty,$ for some increasing sequence $(n_{k})\subset\mathbb{N},$ and for all $\nu\in\mathbb{N}.$ Note that
$F^{n_{k}}_{w}(z^{\nu})=w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}z^{\nu+n_{k}},$
and
$S^{n_{k}}(z^{\nu})=\frac{z^{\nu-n_{k}}}{w_{\nu-1}\cdots w_{\nu-n_{k}}}.$
Combining these expressions with Proposition 2.4, and the assumption (3.2), we see that $F_{w}$ satisfies the Gethner-shapiro criterion. Hence, $F_{w}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A}).$
(ii) To conclude the mixing property of $F_{w}$, proceed as in the proof of (i) by taking $n_{k}=k$ for all $k\geq 1$, in the Genthner-Shapiro criterion.
(iii) We now prove the supercyclicity. Suppose that (3.4) holds. In the supercyclicity criterion, let $D$ and $S$ be defined as above. It suffices to verify that $\lVert F^{n_{k}}_{w}(z^{\nu})\rVert\lVert S^{n_{k}}(z^{\nu})\rVert$ tends to 0 as $k\to\infty,$ for some increasing sequence $(n_{k})\subset\mathbb{N},$ $\nu\in\mathbb{N}.$ Since we have
$\lVert F^{n_{k}}_{w}(z^{\nu})\rVert_{\ell^{p}_{a,b}(\mathtt{A})}\lVert S^{n_{k}}(z^{\nu})\rVert_{\ell^{p}_{a,b}(\mathtt{A})}=\lVert w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}z^{\nu+n_{k}}\rVert\left\lVert\frac{z^{\nu-n_{k}}}{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}}\right\rVert,$
Proposition 2.4, and the assumption (3.4), yield that $F_{w}$ satisfies the Supercyclicity criterion. ∎
A necessary condition for the supercyclicity of $F$ is stated below.
Proposition 3.8.
Let $X$ be a Banach space of analytic functions over an annulus $r<|z|<R$, having bounded evaluation functionals. Let the unweighted forward shift $F$ be bounded on $X$. Then $F$ cannot be supercyclic.
Proof.
Since $F$ is bounded on $X$, we have
$f\in X\Rightarrow zf\in X.$
Thus, $F$ is the multiplication operator $M_{\varphi}(f)=\varphi f$ defined on $X$, where $f\in X$, and $\varphi(z)=z$. However, multiplication operators cannot be supercyclic as its adjoint has more than two distinct eigenvalues with eigenvectors being the non-zero evaluation functionals, cf. [20]. ∎
The next result is on the chaos of $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ which involves lengthy computations.
Theorem 3.9.
Assume that $a=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b=\{b_{n}\}_{n=-\infty}^{\infty}$ satisfy the conditions of Theorem 3.4. Consider the following statements for the bilateral weighted forward shift $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ :
  1. (i)
    $F_{w}$ is chaotic on $\ell^{p}_{a,b}(\mathtt{A})$.
  2. (ii)
    $F_{w}$ has a non-trivial periodic vector.
  3. (iii)
    $\sum_{n=1}^{\infty}\left\lvert\frac{w_{0}w_{1}\cdots w_{n-1}}{a_{n}}\right\rvert^{p}<\infty~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n=1}^{\infty}\left\lvert\frac{1}{w_{-1}w_{-2}\cdots w_{-n}b_{-n-1}}\right\rvert^{p}<\infty.$
  1. (iv)
    For $\nu\in\mathbb{N}$,
    $\sum_{n=1}^{\infty}\left\lvert w_{\nu}\cdots w_{\nu+n-1}\frac{a_{0}\cdots a_{\nu+n-1}}{b_{0}\cdots b_{\nu+n-1}}\right\rvert^{p}<\infty~{}~{}\text{and}~{}~{}\displaystyle\sum_{n=1}^{\infty}\left\lvert\frac{1}{{w_{\nu-1}\cdots w_{\nu-n}}}\frac{b_{-1}\cdots b_{\nu-n}}{a_{-1}\cdots a_{\nu-n}}\right\rvert^{p}<\infty.$
Then $(i)\Rightarrow(ii)\Rightarrow(iii).$ Also, the conditions $(iii)$ and $(iv)$ together yield $(i)$.
Proof.
$(i)\Rightarrow(ii)$ is trivial.
To prove $(ii)\Rightarrow(iii)$, let $f(z)=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}(z)$ be a non-zero periodic vector for $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A}).$ Rearranging into Laurent series, we have $f(z)=\sum_{n=-\infty}^{\infty}A_{n}z^{n}$, where $A_{n}=\lambda_{n-1}b_{n-1}+\lambda_{n}a_{n}$. Let us choose $m\in\mathbb{N}$ such that $F_{w}^{m}f(z)=f(z),$ for all $z$ in some annulus $\mathtt{A}$. Since $F_{w}^{km}f(z)=f(z)$ for all $k\geq 1$, it follows that
$\cdots+w_{-km-1}w_{-km}\cdots w_{-2}A_{-km-1}\frac{1}{z}+w_{-km}w_{-km+1}\cdots w_{-1}A_{-km}+\cdots$ $=$
$\ldots+A_{2}z^{-2}+A_{-1}z^{-1}+A_{0}+\cdots+A_{n}z^{n}\cdots,$
for all $z\in\mathtt{A}$. Using the coefficients of like powers for comparison, we obtain $A_{j+km}=w_{j}w_{j+1}\cdots w_{j+km-1}A_{j}$, $\forall~{}~{}k\in\mathbb{N}$, where $10\leq j\leq m-$ 1. For $j=0,$ we have $\lambda_{km}a_{km}+\lambda_{km-1}b_{km-1}=w_{0}w_{1}\cdots w_{km-1}(\lambda_{0}a_{0}+\lambda_{-1}b_{-1})~{}\forall~{}~{}k\geq 1.$ Hence, we get
$\lvert\lambda_{0}a_{0}+\lambda_{-1}b_{-1}\rvert^{p}\sum_{k=1}^{\infty}\left\lvert\frac{w_{0}w_{1}\cdots w_{km-1}}{a_{km}}\right\rvert^{p}\leq C_{1}\left(r^{p}\sum_{k=1}^{\infty}\left\lvert\lambda_{km-1}\right\rvert^{p}+\sum_{k=1}^{\infty}\left\lvert\lambda_{km}\right\rvert^{p}\right),$
for some $C_{1}>0$. Here, $r:=\sup_{n\geq 0}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert$. As $\{\lambda_{n}\}\in\ell^{p}(\mathbb{Z})$, we have
$\sum_{k=1}^{\infty}\left\lvert\frac{w_{0}w_{1}\cdots w_{km-1}}{a_{km}}\right\rvert^{p}<\infty.$
Along the same lines, for $j=1,2,\ldots,m-1,$ we have $\lambda_{j+km}a_{j+km}+\lambda_{j+km-1}b_{j+km-1}=\\ w_{j+1}w_{j+2}\cdots w_{j+km}(\lambda_{j}a_{j}+\lambda_{j-1}b_{j-1})$, and
$\sum_{k=1}^{\infty}\left\lvert\frac{w_{j}w_{j+1}\cdots w_{j+km-1}}{a_{j+km}}\right\rvert^{p}<\infty.$
Consequently, the series $\sum_{k=1}^{\infty}\left\lvert\frac{w_{0}w_{1}\cdots w_{k-1}}{a_{k}}\right\rvert^{p}$ is convergent. In a similar way, the series
$\sum_{k=1}^{\infty}\left\lvert\frac{1}{w_{-1}w_{-2}\cdots w_{-k}b_{-k-1}}\right\rvert^{p}$ can be shown to be convergent, and thus the implication $(ii)\Rightarrow(iii)$ follows.
We now prove that $(iv)$ implies $(i)$ by applying the chaoticity criterion. Take $D$ to be the space of all Laurent polynomials, and define $S:D\rightarrow D$ by $S(z^{\nu})=\frac{1}{w_{\nu-1}}z^{\nu-1},~{}~{}\nu\in\mathbb{Z}.$ Clearly, $F_{w}S=I$ on $D.$ It remains to show that the series $\sum_{n=0}^{\infty}F_{w}^{n}(f)$ and $\sum_{n=0}^{\infty}S^{n}(f)$ converge unconditionally. It suffices to prove that
$\sum_{n=1}^{\infty}w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n=1}^{\infty}\frac{1}{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}}z^{\nu-n}$
are unconditionally convergent in $\ell^{p}_{a,b}(\mathtt{A}),$ for all $\nu\in\mathbb{N}$. Recalling the Schauder basis expansion from (2.3) and using the Proposition 2.4, for a fixed $\nu\in\mathbb{N}$, we obtain
$\left\lVert w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}\right\rVert_{\ell^{p}_{a,b}(\mathtt{A})}\leq M_{1}\left\lvert\frac{(w_{\nu}\cdots w_{\nu+n-1})(a_{0}\cdots a_{\nu+n-1})}{b_{0}\cdots b_{\nu+n-1}}\right\rvert$
and
$\left\lVert\frac{1}{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}}z^{\nu-n}\right\rVert_{\ell^{p}_{a,b}(\mathtt{A})}\leq M_{2}\left\lvert\frac{b_{-1}\cdots b_{\nu-n}}{(w_{\nu-1}\cdots w_{\nu-n})(a_{-1}\cdots a_{\nu-n})}\right\rvert,$
for some constants $M_{1}$ and $M_{2}$. It follows that the right hand side of the above inequality is $p$-summable. Consequently, the series $\sum_{n=1}^{\infty}w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}$ and $\sum_{n=1}^{\infty}\frac{1}{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}}z^{\nu-n}$ are convergent. The unconditional convergence follows because $\{f_{n}\}$ is a normalized Schauder basis for $\ell^{p}_{a,b}(\mathtt{A})$ which is invariant under $\ell^{1}$ convolutions. We can now conclude that $F_{w}$ satisfies the chaoticity criterion, and (i) follows. ∎
We now turn to study the hypercyclicity, supercyclicity, and chaos for the adjoint $F_{w}^{*}$ using Proposition 3.1. In the following theorem, we derive sufficient conditions for the same.
Theorem 3.10.
For $1<p<\infty$, assume that $a=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b=\{b_{n}\}_{n=-\infty}^{\infty}$ satisfy the conditions of Theorem 3.4. Then, the following hold for adjoint weighted forward shift operator $F_{w}^{*}$:
  1. (i)
    $F_{w}^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})^{*}$ if
    $\liminf_{n\rightarrow\infty}~{}\lvert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}b_{\nu-n-1}\rvert=0=\liminf_{n\rightarrow\infty}~{}\left\lvert\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.5)
  2. (ii)
    $F_{w}^{*}$ is topologically mixing on $\ell^{p}_{a,b}(\mathtt{A})^{*}$ if
    $\lim_{n\rightarrow\infty}\lvert~{}w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}b_{\nu-n-1}\rvert=0=\lim_{n\rightarrow\infty}\left\lvert~{}\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.6)
  3. (iii)
    $F_{w}^{*}$ is supercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$ if
    $\liminf_{n\rightarrow\infty}\lvert~{}w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n}b_{\nu-n-1}\rvert\left\lvert\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.7)
Proof.
Suppose that (3.5) holds, and we will use Gethner-Shapiro criterion to show that $F_{w}^{*}$ is hypercyclic by taking $D=\text{span}\{k_{n}:n\in\mathbb{Z}\}$, and $S$ as the map defined by
$S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\in\mathbb{Z}.$
By Proposition 2.3, the set $D$ is dense in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Since $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ we just need to show that $(F_{w}^{*})^{n_{k}}(k_{\nu})$ and $S^{n_{k}}(k_{\nu})$ both tends to 0 as $k\to\infty,$ for any increasing sequence $(n_{k})\subset\mathbb{N},$ $\nu\in\mathbb{N}.$ But this is clear since by Proposition (2.2), $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<1,$ and $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$
$\lVert(F_{w}^{*})^{n_{k}}(k_{\nu})\rVert=\lVert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}k_{\nu-n_{k}}\rVert\leq M_{1}\lvert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}b_{\nu-n_{k}-1}\rvert$
and
$\lVert S^{n_{k}}(k_{\nu})\rVert=\left\lVert\frac{k_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rVert\leq M_{2}\left\lvert\frac{a_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rvert,$
for some constants $M_{1},M_{2}>0$ and $n_{k}>\nu.$ Now by the assumption (3.5), we get that $F_{w}^{*}$ satisfies the Gethner-shapiro criterion. Hence, $F_{w}^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})^{*}.$
(ii) The proof here is verbatim. Taking $n_{k}=k$ for all $k\geq 1$, and pursue exactly as in the proof of (i).
(iii) Suppose that (3.7) holds. In the supercyclicity criterion, take the set $D$ and the map defined as above. It suffices to prove that $\lVert(F_{w}^{*})^{n_{k}}(k_{\nu})\rVert\lVert S^{n_{k}}(k_{\nu})\rVert$ tends to zero as $k\to\infty,$ for any increasing sequence $(n_{k})\subset\mathbb{N},$ $\nu\in\mathbb{N}.$ But this is clear since by Proposition (2.2), $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<1,$ and $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$
$\lVert(F_{w}^{*})^{n_{k}}(k_{\nu})\rVert\lVert S^{n_{k}}(k_{\nu})\rVert$ $=$ $\lVert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}k_{\nu-n_{k}}\rVert\left\lVert\frac{k_{\nu+n_{k}}}{w_{\nu}w_{t+1}\cdots w_{t+n_{k}-1}}\right\rVert$
$\leq$ $M\lvert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}b_{\nu-n_{k}-1}\rvert\left\lvert\frac{a_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rvert,$
for some constant $M>0$ and $n_{k}>\nu.$ Now with the assumption (3.7), we get that $F_{w}^{*}$ satisfies the Supercyclicity criterion. Hence, $F_{w}^{*}$ is supercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$
Theorem 3.11.
Let $\ell^{p}_{a,b}(\mathtt{A})$ $1<p<\infty,$ be an analytic Banach space over the unit annulus corresponding to $a=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b=\{b_{n}\}_{n=-\infty}^{\infty}$ satisfying the conditions of Theorem 3.4. Consider the following statements:
  1. (i)
    $F_{w}^{*}$ is chaotic on $\ell^{p}_{a,b}(\mathtt{A})^{*}$.
  2. (ii)
    $F_{w}^{*}$ has a non-trivial periodic vector.
  3. (iii)
    $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{p}<\infty$ and $\sum_{n=1}^{\infty}\lvert w_{-1}w_{-2}\cdots w_{-n}b_{-n-1}\rvert^{p}<\infty.$
Then $(iii)\Rightarrow(i)\Rightarrow(ii).$
Proof.
(i) implies (ii) trivial.
To see that (iii) implies (i), suppose that $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}\cdots w_{n-1}}\right\rvert^{p}<\infty$ and $\sum_{n=1}^{\infty}\lvert w_{-1}\cdots w_{-n}b_{-n-1}\rvert^{p}<\infty$ hold. We apply the chaoticity criterion to show that $F_{w}^{*}$ is chaotic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Consider $D=\text{span}\{k_{n}(z):n\in\mathbb{Z}\}$. Then by the Propostion 2.3, the set $D$ is dense in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Define $S:D\rightarrow D$ by $S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\in\mathbb{Z}.$ Since $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ we just need to show that $\sum_{n=0}^{\infty}(F_{w}^{*})^{n}(f)$ and $\sum_{n=0}^{\infty}S^{n}(f)$ converge unconditionally, for each $f\in D.$ We prove that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ and $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ are unconditionally convergent in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Since $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\in\mathbb{Z}$, we have
$\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ $=$ $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}(a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*})$
$=$ $\frac{a_{1}}{w_{0}}\frac{b_{0}}{a_{1}}f_{0}^{*}+\sum_{n=1}^{\infty}\left(\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}+\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\frac{b_{n}}{a_{n+1}}\right)f_{n}^{*}.$
As $\limsup_{n}|b_{n}/a_{n+1}|<1$, one gets $N\in\mathbb{N}$ and $r<1$ such that $|b_{n}/a_{n+1}|<r$ for all $n\geq N$. Hence,
$\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}+\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\frac{b_{n}}{a_{n+1}}\right\rvert\leq\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert+r\left\lvert\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\right\rvert,$
for all $n\geq N$. The right hand side of the above inequality is the $\ell^{p}$ element $\left\{\frac{a_{n+1}}{w_{0}\cdots w_{n}}\right\}_{n=1}^{\infty},$ and hence it is absolutely $p$-summable. Consequently, the series $\sum_{n\geq N}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is convergent. The unconditional convergence follows because $\{f_{n}^{*}\}$ is an unconditional basis. Similarly, to show that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ is unconditionally convergent, we write
$\sum_{n=1}^{\infty}w_{-1}\cdots w_{-n}k_{-n}$ $=$ $\sum_{n=1}^{\infty}w_{-1}\cdots w_{-n}(a_{-n}f_{-n}^{*}+b_{-n-1}f_{-n-1}^{*})$
$=$ $w_{-1}b_{-2}\frac{a_{-1}}{b_{-2}}f_{-1}^{*}+\sum_{n=2}^{\infty}\left(w_{-1}\cdots w_{-n+1}b_{-n}+w_{-1}\cdots w_{-n}\frac{b_{-n-1}a_{-n}}{b_{-n-1}}\right)f_{-n}^{*}.$
As $\limsup_{n}|a_{-n}/b_{-n-1}|<1$, one gets $N\in\mathbb{N}$ and $R<1$ such that $|a_{-n}/b_{-n-1}|<R$ for all $n\geq N$. Hence,
$\left\lvert w_{-1}\cdots w_{-n+1}b_{-n}+w_{-1}\cdots w_{-n}b_{-n-1}\frac{a_{-n}}{b_{-n-1}}\right\rvert\leq\left\lvert w_{-1}\cdots w_{-n+1}b_{-n}\right\rvert+R\left\lvert w_{-1}\cdots w_{-n}b_{-n-1}\frac{a_{-n}}{b_{-n-1}}\right\rvert$
for all $n\geq N$. The right hand side of the above inequality is the $\ell^{p}$ element $\left\{w_{-1}\cdots w_{-n}b_{-n-1}\right\}_{n\geq 1}$ and hence it is absolutely $p$-summable. Consequently, the series $\sum_{n\geq N}\frac{1}{w_{-1}w_{-2}\cdots w_{-n}}k_{-n}$ is unconditionally convergent. Hence, $F_{w}^{*}$ satisfies the chaoticity criterion, and (i) follows.
3.3. Examples and Illustrations
We illustrate the above results on dynamics with explicit examples. Recall that a Laurent series $\sum_{n=-\infty}^{\infty}a_{n}z^{n}$ converges for all $z$ such that $r<\lvert z\rvert<R,$ where
[…]
Example 3.12.
Let $a_{0}=b_{0}=w_{0}=1,$ and $a_{n}=1,$ $b_{n}=\frac{1}{n},$ $w_{n}=\frac{1}{(n+1)!},$ for all $n\geq 1,$, along with $a_{n}=\frac{n2^{n}}{2^{2}},$ $b_{n}=n2^{n},$ $w_{n}=8,$ for all $n\leq-1.$ Now by Theorem 2.1, for $1<p<\infty,$
$\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}=\sum_{n=1}^{\infty}\left(\left(\frac{n}{2^{n}}\frac{1}{2^{2}}+\frac{n}{2^{n}}\right)\lvert z\rvert^{-n}\right)^{q}+\sum_{n=0}^{\infty}\left(\left(1+\frac{1}{n}\right)\lvert z\rvert^{n}\right)^{q},$
where $11/p+1/q=$ 1. A simple computation shows that the natural annulus $\mathtt{A}$ of $\ell^{p}_{a,b}(\mathtt{A})$ is $r<|z|<R$, where $r=\frac{1}{2}$ and $R=1.$ Therefore $\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}<\infty$ if $\frac{1}{2}<\lvert z\rvert<1.$ In view of Theorem 3.4, $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}),$ where $\mathtt{A}$ is the given by annulus $\frac{1}{2}<\lvert z\rvert<1.$ Additionally, $F_{w}$ satisfies the conditions of Theorems 3.7 and 3.9, and so $F_{w}$ is mixing and chaotic on $\ell^{p}_{a,b}(\mathtt{A}).$
Example 3.13.
In the previous example, if we let $w_{n}=4,$ for all $n\leq-1,$ $w_{n}=\frac{1}{(n+1)!},$ for all $n\geq 0,$ and $\{a_{n}\},\{b_{n}\}$ being the same as mentioned above, then we see that by Theorem 3.7, $F_{w}$ is supercyclic but not hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})$ as
$\lim_{n\rightarrow\infty}\left\lvert\frac{(w_{\nu-1}\cdots w_{\nu-n})(a_{-1}\cdots a_{\nu-n})}{b_{-1}\cdots b_{\nu-n}}\right\rvert<\infty,$
for $\nu\in\mathbb{N}$ with $n>\nu.$
Next, we provide an example of a space of the form $\ell^{p}_{a,b}(\mathtt{A})$ such that $F_{w}^{*}$ is mixing and chaotic.
Example 3.14.
Let $a_{0}=b_{0}=w_{0}=1,$ and $a_{n}=\frac{n}{2^{n}},$ $b_{n}=\frac{n}{2^{n+2}},$ $w_{n}=8,$ for all $n\geq 1,$ and $a_{n}=\frac{1}{n},$ $b_{n}=1,$ $w_{n}=\frac{1}{(-n+2)!},$ for all $n\leq-1.$ By Theorem 2.1, we have for $1<p<\infty,$
$\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}=\sum_{n=1}^{\infty}\left(\left(\frac{1}{n}+1\right)\lvert z\rvert^{-n}\right)^{q}+\sum_{n=0}^{\infty}\left(\left(\frac{n}{2^{n}}+\frac{n}{2^{n+2}}\right)\lvert z\rvert^{n}\right)^{q},$
where $11/p+1/q=$ 1. In this case, $r=1$, $R=2$, and hence, the annulus $\mathtt{A}$ is $21<|z|<$ 2. Further, $\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}<\infty$ if $21<\lvert z\rvert<$ 2 . In addition, by the conditions in Theorem 3.4, the weighted shift $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}).$ In view of Theorems 3.10 and 3.11, the adjoint $F_{w}^{*}$ is mixing and chaotic on $\ell^{p}_{a,b}(\mathtt{A})^{*}.$ However, the $F_{w}$ itself is not supercyclic.
4. Adjoints of unilateral weighted shifts
In this section, we study the dynamical properties of the adjoint of the unilateral forward shift $F_{w}(\sum_{n=0}^{\infty}\lambda_{n}z^{n})=\sum_{n=0}^{\infty}\lambda_{n}w_{n}z^{n+1}$ defined on $\ell^{p}_{a,b}(\mathbb{D})$, where $w=\{w_{n}\}_{n=0}^{\infty}$ is a sequence of non-zero complex numbers. The differences between $F_{w}$ acting on $\ell^{p}_{a,b}(\mathbb{D})$ and the classical weighted forward shifts on $\ell^{p}$ will be pointed out in Section $5$.
First, we derive some sufficient conditions for the unilateral $F_{w}$ to bounded in terms of $\{a_{n}\}$, $\{b_{n}\}$ and the weight sequence $\{w_{n}\}$. For the unweighted right right $F$, some necessary and sufficient conditions for $F$ to be bounded on $\ell^{2}_{a,b}(\mathbb{D})$ were obtained by Adams and McGuire [1]. Weighted backward shifts $B_{w}$ on $\ell^{p}_{a,b}(\mathbb{D})$ were studied by Das and Mundayadan [15].
We compute the matrix representation of the operator $F_{w}$ acting on $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ with respect to the ordered Schauder basis $\{f_{n}\}_{n\geq 0}$. For this, we express a monomial $z^{n}$ as an expansion in the basis $\{f_{n}\}$; see (4.1) below. Indeed, fix $n\geq 0,$ and write $z^{n}=\sum_{j=0}^{\infty}\alpha_{j}f_{j}$ for some $\alpha_{j}\in\mathbb{C}.$ Then $z^{n}=\alpha_{0}a_{0}+\sum_{j=1}^{\infty}(\alpha_{j-1}b_{j-1}+\alpha_{j}a_{j})z^{j}.$ Comparing the coefficients of the same powers, a simple computation gives that $\alpha_{0}=\alpha_{1}=\cdot\cdot\cdot=\alpha_{n-1}=0,$ $\alpha_{n}=\frac{1}{a_{n}}$, and $\alpha_{n+k}=\frac{(-1)^{k}}{a_{n}}\frac{b_{n}b_{n+1}\cdot\cdot\cdot b_{n+k-1}}{a_{n+1}a_{n+2}\cdot\cdot\cdot a_{n+k}}$, where $k\geq 1$. From these expressions, we obtain
$z^{n}=\frac{1}{a_{n}}\sum_{j=0}^{\infty}(-1)^{j}\left(\frac{\prod_{k=0}^{j-1}b_{n+k}}{\prod_{k=0}^{j-1}a_{n+k+1}}\right)f_{n+j},\,\,\,\,(n\geq 0),$ (4.1)
where the coefficient for $j=0$ should be understood as $1$. (We refer to [1] for the previous calculations in the context of the study of tridiagonal kernels. Also, see [15].)
Acting $F_{w}$ on the normalized Schauder basis $f_{n}(z)=(a_{n}+b_{n}z)z^{n},n\geq 0$ in $\ell^{p}_{a,b}(\mathbb{D})$, we obtain
$F_{w}(f_{n})(z)$ $=$ $w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}$
$=$ $\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+\left(\frac{w_{n+1}b_{n}}{a_{n+2}}-\frac{w_{n}a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right)a_{n+2}z^{n+2}$
$=$ $\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+c_{n}\sum_{j=0}^{\infty}(-1)^{j}\left(\frac{\prod_{k=0}^{j-1}b_{n+2+k}}{\prod_{k=0}^{j-1}a_{n+3+k}}\right)f_{n+2+j},$
where
$c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}.$
Therefore, the matrix of $F_{w}$ with respect to the basis $\{f_{n}\}_{n\geq 0}$ is
$[F_{w}]:=\begin{bmatrix}0&0&0&0&\cdots\\ w_{0}\frac{a_{0}}{a_{1}}&0&0&0&\ddots\\ c_{0}&w_{1}\frac{a_{1}}{a_{2}}&0&0&\ddots\\ -c_{0}\frac{b_{2}}{a_{3}}&c_{1}&w_{2}\frac{a_{2}}{a_{3}}&0&\ddots\\ c_{0}\frac{b_{2}b_{3}}{a_{3}a_{4}}&-c_{1}\frac{b_{3}}{a_{4}}&c_{2}&w_{3}\frac{a_{3}}{a_{4}}&\ddots\\ -c_{0}\frac{b_{2}b_{3}b_{4}}{a_{3}a_{4}a_{5}}&c_{1}\frac{b_{3}b_{4}}{a_{4}a_{5}}&-c_{2}\frac{b_{4}}{a_{5}}&c_{3}&\ddots\\ \vdots&\ddots&\ddots&\ddots&\ddots\end{bmatrix}.$
In the following theorem we provide a sufficient condition for $F_{w}$ to be bounded. For the particular cases of $p=2$ and the unweighted forward shift, we refer to [1], [16], and [17].
Theorem 4.1.
If
$\sup_{n\geq 0}~{}\Big{\{}\left\lvert\frac{w_{n}a_{n}}{a_{n+1}}\right\rvert,~{}|c_{n}|\Big{\}}<\infty,~{}\text{and}~{}\sum_{n=3}^{\infty}\sup_{j\geq 0}~{}\left|\frac{c_{j}b_{j+2}b_{j+3}\cdots b_{j+n-1}}{a_{j+3}a_{j+4}\cdots a_{j+n}}\right|<\infty,$
then $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$, where $1\leq p<\infty$.
Proof.
We give only an outline of the proof. The matrix of $F_{w}$ can be formally written as $[F_{\alpha}]$+$\sum_{n=2}^{\infty}[F_{n}]$. Here, $[F_{\alpha}]$ is the matrix of the standard weighted forward shift $F_{\alpha}(e_{i})\mapsto\alpha_{i}e_{i+1}$ on $\ell^{p}$, $i\geq 1$, having weights
$\alpha_{i}=\frac{w_{i}a_{i}}{a_{i+1}},\hskip 28.45274pt(i\geq 0),$
on $\ell^{p}$. The matrix $[F_{n}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $n$-th subdiagonal, where $n\geq 2$. Observe that $[F_{n}]$ is the matrix of suitable powers of a weighted forward shift $F_{n}$ for $n\geq 2$.
It follows, respectively by the first two assumptions in the theorem, that the weighted shift $F_{\alpha}$ and $F_{2}$ are bounded on $\ell^{p}$. Since
$\|F_{n}\|=\sup_{j\geq 0}~{}\left|\frac{c_{j}b_{j+2}b_{j+3}\cdots b_{j+n-1}}{a_{j+3}a_{j+4}\cdots a_{j+n}}\right|,$
the third condition in the theorem gives that $F_{n}$ is bounded, and $\sum_{n\geq 3}\|F_{n}\|<\infty$, with respect to the operator norm. Hence, the shift $F_{w}$ is bounded. This completes the proof of the theorem. ∎
Remark 4.2.
The following conditions for the boundedness of $F_{w}$ are simpler and stronger, compared to those in the above theorem. If
$\sup_{n\geq 0}~{}\left\lvert\frac{w_{n}a_{n}}{a_{n+1}}\right\rvert<\infty\hskip 11.38092pt\text{and}\hskip 11.38092pt\lim\sup_{n}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$ (4.2)
then the assumptions in Theorem 4.1 are easily satisfied so that $F_{w}$ is bounded.
To infer the dynamics of $F_{w}^{*}$, we need the following. Its proof is omitted since it is similar to the bilateral version proved earlier. Also, we refer to [33] wherein the authors studied the chaos for $F^{*}$ in the tridiagonal spaces.
Proposition 4.3.
Let $X$ be a Banach space of analytic functions on the unit disc, having evaluation functionals continuous. Let $k_{n}$, $n\geq 0$, denote the $n$-th Taylor coefficient functional. Then, the span of $\{k_{n}:n\geq 0\}$ is weak-dense in $X^{*}$, and norm dense if $X^{*}$ is reflexive.
Using above result, we prove:
Theorem 4.4.
Let $\ell^{p}_{a,b}(\mathbb{D}),$ $1<p<\infty,$ be the analytic Banach space over the unit disc $\mathbb{D}$ corresponding to $a=\{a_{n}\}_{n=0}^{\infty}$ and $b=\{b_{n}\}_{n=0}^{\infty}$ satisfying the conditions in (4.2). Then, the following hold for the adjoint $F_{w}^{*}$:
  1. (i)
    $F_{w}^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathbb{D})^{*}$ if and only if
    $\liminf_{n\rightarrow\infty}\Big{|}\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\Big{|}=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (4.3)
  2. (ii)
    $F_{w}^{*}$ is topologically mixing if and only if
    $\lim_{n\rightarrow\infty}\Big{\lvert}\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\Big{\rvert}=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (4.4)
  3. (iii)
    The adjoint operator $F_{w}^{*}$ is always supercyclic.
Proof.
(i) First suppose that (4.3) holds, and we will use Gethner-shapiro criterion to show that $F_{w}^{*}$ is hypercyclic. Consider $D:=\text{span}\{k_{n}:n\geq 0\}$. Then $D$ is dense in $\ell^{p}_{a,b}(\mathbb{D})^{*},$ by the previous proposition. Define $S:D\rightarrow D$ by
$S(k_{n})=\frac{1}{w_{n}}k_{n+1},$
$\forall~{}~{}n\geq 0.$ Since $F_{w}^{*}S(f)=f$ and $(F_{w}^{*})^{n}f\to 0,$ as $n\to\infty,$ $\forall$ $f\in D,$ we just need to show that $S^{n_{k}}(k_{\nu})$ tends to 00, as $k\to\infty,$ for any increasing sequence $(n_{k})\subset\mathbb{N},$ $\nu\in\mathbb{N}.$ But this is clear since by Proposition (2.2) and $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$
$\lVert S^{n_{k}}(k_{\nu})\rVert=\left\lVert\frac{k_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rVert\leq M_{2}\left\lvert\frac{a_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rvert,$
for some constants $M_{1}>0.$ Now by the assumption (4.3), we get that $F_{w}^{*}$ satisfies the Gethner-Shapiro criterion. Hence, $F_{w}^{*}$ is hypercyclic.
Conversely, assume that $F_{w}^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathbb{D})^{*}$ and we have to show that (4.3) holds. Let $f=z^{\nu},\nu\in\mathbb{N}$ and $g$ be a non-zero vector in $\ell^{p}_{a,b}(\mathbb{D})^{*}.$ Then
$\langle(F_{w}^{*})^{n}g,f\rangle=\langle g,F_{w}^{n}f\rangle=\langle g,w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}\rangle.$
Now by Proposition 4.1 of [15] we know that
$\lVert w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}z^{\nu+n}\rVert\leq M\left\lvert\frac{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}{a_{\nu+n}}\right\rvert,$
for some constant $M>0.$ If possible, let $\lim\inf_{n\rightarrow\infty}\left\lvert\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert>0,$ then we get that $\{(F_{w}^{*})^{n}g\}_{n\geq 0}$ is a bounded set. This contradicts the fact that $F_{w}^{*}$ has a hyperyclic vector, and hence, $\lim\inf_{n\rightarrow\infty}\left\lvert\frac{a_{\nu+n}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n-1}}\right\rvert=0.$
(ii) Apply the Gethner-Shapiro criterion with $n_{k}=k$ for $k\geq 1$, and proceed as above.
(iii) The supercyclicity criterion with $D$ and $S$ as above, taking $n_{k}=k$, $k\geq 1$, is satisfied.
Our next result on the chaos of $F_{w}^{*}$. In [33], the authors provided a sufficient condition for the adjoint $F^{*}$ to be chaotic on a general analytic reproducing kernel space over $\mathbb{D}$ which includes the case $\ell^{2}_{a,b}(\mathbb{D})$. The following representations of $f_{n}^{*}$ in terms of $k_{n}$ will be utilized to get the chaos of the adjoint of forward weighted shifts.
Proposition 4.5.
Let $k_{n}$ denote the co-ordinate functional
$f\mapsto\frac{f^{(n)}(0)}{n!},$
defined on $\ell^{p}_{a,b}(\mathbb{D})$ or $c_{0,a,b}(\mathbb{D})$, where $n\geq 0$. We have
$f_{n}^{*}=\frac{1}{a_{n}}k_{n}-\frac{b_{n-1}}{a_{n}a_{n-1}}k_{n-1}+\frac{b_{n-1}b_{n-2}}{a_{n}a_{n-1}a_{n-2}}k_{n-2}-\cdots+(-1)^{n}\frac{b_{n-1}\cdots b_{0}}{a_{n}\cdots a_{0}}k_{0},$
for all $n\geq 1$ and $f_{0}^{*}=\frac{1}{a_{0}}k_{0}.$
Proof.
If $f\in\ell^{p}_{a,b},$ then $f(z)=\sum_{n=0}^{\infty}\lambda_{n}f_{n}(z)=\sum_{n=0}^{\infty}\lambda_{n}(a_{n}+b_{n}z)z^{n}.$ Rearranging into a power series, we obtain $f(z)=\lambda_{0}a_{0}+\sum_{n\geq 1}(\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1})z^{n},$ and so, $k_{n}(f)=\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}$ and $k_{0}(f)=\lambda_{0}a_{0}.$ Now for $f\in\ell^{p}_{a,b},$ we have
$\frac{1}{a_{n}}k_{n}(f)-\cdots+(-1)^{n}\frac{b_{n-1}\cdots b_{0}}{a_{n}\cdots a_{0}}k_{0}(f)$ $=$ $\frac{1}{a_{n}}(\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1})-\cdots+(-1)^{n}\frac{b_{n-1}\cdots b_{0}}{a_{n}\cdots a_{0}}(\lambda_{0}a_{0})$
$=$ $\lambda_{n}$
$=$ $f_{n}^{*}(\sum_{m=0}^{\infty}\lambda_{m}f_{m})$
$=$ $f_{n}^{*}(f),$
for all $n\geq 1$ and $\frac{1}{a_{0}}k_{0}(f)=\frac{1}{a_{0}}(\lambda_{0}a_{0})=\lambda_{0}=f_{0}^{*}(\sum_{m=0}^{\infty}\lambda_{m}f_{m})=f_{0}^{*}(f).$
We now have necessary and sufficient conditions for $F_{w}^{*}$ to be chaotic.
Theorem 4.6.
Assume that the conditions stated in (4.2) are satisfied for $\ell^{p}_{a,b}(\mathbb{D})$, where $1<p<\infty$. Consider the following statements:
  1. (i)
    $F_{w}^{*}$ is chaotic on $\ell^{p}_{a,b}(\mathbb{D})^{*}.$
  2. (ii)
    $F_{w}^{*}$ has a non-trivial periodic vector.
  3. (iii)
    $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{p}<\infty.$
Then $(iii)\Rightarrow(i)\Rightarrow(ii)$. In addition, if $\frac{1}{p}+\frac{1}{q}=1,$ and
$\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{b_{n}b_{n+1}\cdots b_{n+j-1}}{a_{n+1}a_{n+2}\cdots a_{n+j}}\right\rvert^{q}\right)^{\frac{p}{q}}<\infty.$
then $(ii)\Rightarrow(iii)$.
Proof.
(i) implies (ii) trivial.
To see that (ii) implies (iii), let $F=\sum_{m=0}^{\infty}\lambda_{m}f_{m}^{*}$ be a non-zero periodic vector for $F_{w}^{*}$ on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$, where $f_{m}^{*}$, $m\geq 0$, forms an unconditional Schauder basis for $\ell^{p}_{a,b}(\mathbb{D})^{*}$. Now by Proposition 4.5, each $f_{m}^{*},m\geq 1$ can be written as
$f_{m}^{*}=\frac{1}{a_{m}}k_{m}-\frac{b_{m-1}}{a_{m}a_{m-1}}k_{m-1}+\frac{b_{m-1}b_{m-2}}{a_{m}a_{m-1}a_{m-2}}k_{m-2}-\cdots+(-1)^{m}\frac{b_{m-1}\cdots b_{0}}{a_{m}\cdots a_{0}}k_{0},$
and $f_{0}^{*}=\frac{1}{a_{0}}k_{0}.$ Let $\nu\in\mathbb{N}$ such that $(F_{w}^{*})^{\nu}F=F$ for $z\in\mathbb{D}$. Then $(F_{w}^{*})^{n\nu}F=F$ for all $n\geq 1$. Equating the coefficient of $k_{0}$ both sides, we get
$\frac{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}{a_{n\nu}}\left(\lambda_{n\nu}-\lambda_{n\nu+1}\frac{b_{n\nu}}{a_{n\nu+1}}+\cdots\right)=\frac{1}{a_{0}}\left(\lambda_{0}-\lambda_{1}\frac{b_{0}}{a_{1}}+\lambda_{2}\frac{b_{0}b_{1}}{a_{1}a_{2}}-\cdots\right),$
implies that
$\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{p}\left\lvert\lambda_{0}-\lambda_{1}\frac{b_{0}}{a_{1}}+\lambda_{2}\frac{b_{0}b_{1}}{a_{1}a_{2}}-\cdots\right\rvert^{p}=\left\lvert a_{0}\right\rvert^{p}\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}-\lambda_{n\nu+1}\frac{b_{n\nu}}{a_{n\nu+1}}+\cdots\right\rvert^{p}.$
Using Hölder inequality, we can get
$\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}-\lambda_{n\nu+1}\frac{b_{n\nu}}{a_{n\nu+1}}+\lambda_{n\nu+2}\frac{b_{n\nu}b_{n\nu+1}}{a_{n\nu+1}a_{n\nu+2}}-\cdots\right\rvert^{p}<\infty,$
as $\{\lambda_{n}\}_{n}\in\ell^{p}(\mathbb{N})$ and by assumption $\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{b_{n}b_{n+1}\cdots b_{n+j-1}}{a_{n+1}a_{n+2}\cdots a_{n+j}}\right\rvert^{q}\right)^{\frac{p}{q}}<\infty.$ Therefore
$\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{p}<\infty.$ Similarly, equating the coefficient of $k_{j},j=1,2,\cdots,\nu-1$ both sides we get $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu+j}}{w_{n\nu+j-1}w_{n\nu+j-2}\cdots w_{j}}\right\rvert^{p}<\infty.$ Consequently, the series $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{p}<\infty.$
To see that (iii) implies (i), suppose that $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{p}<\infty$ holds. We apply the chaoticity criterion to show that $F_{w}^{*}$ is chaotic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Consider $D=\text{span}\{k_{n}(z):n\geq 0\}$. Then by the previous Proposition, the set $D$ is dense in $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Define $S:D\rightarrow D$ by $S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\geq 0.$ Clearly, $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ and $\sum_{n=0}^{\infty}(F_{w}^{*})^{n}(f)$ converges unconditionally for each $f\in D.$ We just need to show that $\sum_{n=0}^{\infty}S^{n}(f)$ converges unconditionally, for each $f\in D.$ We prove that $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is unconditionally convergent in $\ell^{p}_{a,b}(\mathbb{D})^{*}.$ Since $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\geq 1$ and $k_{0}=a_{0}f_{0}^{*}.$ Therefore, we have
$\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ $=$ $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}(a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*})$
$=$ $\frac{a_{1}}{w_{0}}\frac{b_{0}}{a_{1}}f_{0}^{*}+\sum_{n=1}^{\infty}\left(\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}+\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\frac{b_{n}}{a_{n+1}}\right)f_{n}^{*}.$
As $\limsup_{n}|b_{n}/a_{n+1}|<1$, one gets $N\in\mathbb{N}$ and $r<1$ such that $|b_{n}/a_{n+1}|<r$ for all $n\geq N$. Hence,
$\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}+\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\frac{b_{n}}{a_{n+1}}\right\rvert\leq\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert+r\left\lvert\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\right\rvert,$
for all $n\geq N$. The right hand side of the above inequality is the $\ell^{p}$ element $\left\{\frac{a_{n+1}}{w_{0}\cdots w_{n}}\right\}_{n=1}^{\infty},$ and hence it is absolutely $p$-summable. Consequently, the series $\sum_{n\geq N}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is convergent. The unconditional convergence occurs because $\{f_{n}^{*}\}$ is an unconditional Schauder basis. Hence, $F_{w}^{*}$ satisfies the chaoticity criterion, and (i) follows.
We illustrate the above results on the unilateral weighted shift $F_{w}$ with an example.
Example 4.7.
($F_{w}=$ nuclear $+$ a standard weighted shift) Choose $a_{0}=b_{0}=w_{0}=1,$ and
$a_{n}=\frac{1}{n2^{n-1}},~{}b_{n}=\frac{1}{(n+1)2^{2n}},~{}w_{n}=4,~{}n\geq 1.$
Then $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}),$ as it satisfies the conditions of Theorem 4.1. By Theorems 4.4 and 4.6, the adjoint $F_{w}^{*}$ is mixing and chaotic on $\ell^{p}_{a,b}(\mathbb{D})^{*}.$
We close this section with a remark. From the matrix representation of $F_{w}$, it follows that $F_{w}$ is similar to an operator of the form $A+F_{\alpha}$ for some bounded operator $A$ and a weighted forward shift $F_{\alpha}$ on $\ell^{p}(\mathbb{N})$. Hence, the adjoint $F_{w}^{*}$ is similar to a perturbation of a weighted backward shift on $\ell^{q}(\mathbb{N})$.
5. Notable differences: Eigenvectors, periodic vectors, and a zero-one law
We show that the forward weighted shifts $F_{w}$ on our spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $\ell^{p}_{a,b}(\mathbb{D})$ are in general very different from the classical bilateral/unilateral weighted shifts. Below, we show in both the cases, that the adjoint of an unweighted forward shift operator can have a non-zero periodic vector without being hypercyclic. It is well known that if a bilateral weighted shift on $\ell^{p}(\mathbb{Z})$ has an orbit having a non-zero limit point, then the operator has a dense orbit, cf. [12] and [11]. Such a property is called as a zero-one law of orbital limit points for an operator. Other distinguishing properties are stated in this section.
Proposition 5.1.
Let $|\lambda|=1$. Consider the space $\ell^{p}_{a,b}(\mathbb{D})$, $1<p<\infty$, where
$a_{n}=1~{}\text{and}~{}b_{n}=\frac{-1}{\lambda}\frac{n+1}{n+2},~{}n\geq 0.$
Then the unweighted forward shift $F$ is bounded, non-hypercyclic, and it is similar to $S+N$, where $S$ is the right shift on $\ell^{p}$ and $N$ is some nuclear (hence, compact) operator. Further,
(i) if $\lambda$ is a root of the unity, then $F^{*}$ has a non-trivial periodic vector in $\ell^{p}_{a,b}(\mathbb{D})^{*}$.
(ii) if $\lambda$ is not a root of the unity, then the zero-one law of orbital limit points fails for $F^{*}$.
Proof.
By Theorem $4.1$ we see that $F$ is bounded. However, to get an exact picture of $F$, we recall the matrix of $F$:
$[F]:=\begin{bmatrix}0&0&0\cdots\\ 1&0&0\ddots\\ \frac{1}{\lambda}\frac{1}{2.3}&1&0\ddots\\ \frac{1}{\lambda^{2}}\frac{1}{2.4}&\frac{1}{\lambda}\frac{1}{3.4}&1\ddots\\ \frac{1}{\lambda^{3}}\frac{1}{2.5}&\frac{1}{\lambda^{2}}\frac{1}{3.5}&\frac{1}{\lambda}\frac{1}{4.5}\ddots\\ \frac{1}{\lambda^{4}}\frac{1}{2.6}&\frac{1}{\lambda^{3}}\frac{1}{3.6}&\frac{1}{\lambda^{2}}\frac{1}{4.6}\ddots\\ \vdots&\ddots&\ddots~{}~{}~{}\ddots\end{bmatrix}.$
Now by the Theorem 4.1 we see that $F$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}).$ Going back to the proof of Theorem $4.1$, we see that each $F_{n}$ is nuclear and $\|F\|_{n}\leq\frac{1}{n^{2}}$, $n\geq 1$. Hence, the sum $\sum_{n\geq 1}F_{n}$ is convergent in with respect to the nuclear norm. Consequently, $F$ is similar to “nuclear + right shift” on $\ell^{p}(\mathbb{N})$.
Again by Theorem 4.4 we know that $F^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathbb{D})^{*}$ if and only if
$\lim\inf_{n\to\infty}(\lvert a_{n}\rvert+\lvert b_{n-1}\rvert)=0.$
However, in our case this condition is not satisfied. Hence $F^{*}$ is not hypercyclic.
The forward shift operator $F$ in $\ell^{p}_{a,b}(\mathbb{D})$ is given by $F(f(z))=zf(z),$ for all $f\in\ell^{p}_{a,b}(\mathbb{D})$ and $z\in\mathbb{D}.$ Now if $F$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}),$ then for $f\in\ell^{p}_{a,b}(\mathbb{D})$ and $w\in\mathbb{D},$ we have
$F^{*}(ev_{z})f=(ev_{z}\circ F)f=ev_{z}(zf)=zf(z)=z(ev_{z})f,$
that is, $F^{*}(ev_{z})=z(ev_{z}).$ It shows that $z$ is an eigenvalue of $F^{*}$ corresponding to the eigenvector $ev_{z}$ on $\ell^{p}_{a,b}(\mathbb{D})^{*}.$ This is in fact a well known fact, cf. Godefroy and Shapiro [20]. When $a_{n}=1$ and $b_{n}=\frac{-1}{\lambda}\frac{n+1}{n+2},$ $n\geq 0,$ where $\lambda$ is a root of the unity. Now
$\sum_{n=0}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=0}^{\infty}\left\lvert\left(1-\frac{1}{\lambda}\frac{n+1}{n+2}\lambda\right)\lambda^{n}\right\rvert^{q}<\infty,$
for $1<q<\infty.$ Therefore the natural domain of $\ell^{p}_{a,b}(\mathbb{D})$ is given by $\text{Dom}_{(p)}=\mathbb{D}\cup\{\lambda\}.$ Since $F^{*}(ev_{z})=\lambda(ev_{z}),$ for all $z\in\mathbb{D}.$ It follows that $ev_{\lambda}\in\ell^{p}_{a,b}(\mathbb{D})^{*}.$ Also, we have $F^{*}(ev_{\lambda})=\lambda(ev_{\lambda}).$
If $\lambda$ is a square root of unity, say $\lambda^{m}=1$, then $F^{*m}(ev_{\lambda})=ev_{\lambda}$. Hence $F^{*}$ has a nontrivial periodic vector, namely $ev_{\lambda}$, which gives $(i)$.
Now, assume that $\lambda$ is not a root of unity. It is well known, for such $\lambda$, that the sequence $\{\lambda^{n}:n\geq 1\}$ is dense in the unit circle of $\mathbb{C}$. As
$\{F^{*n}({ev}_{\lambda}):n\geq 0\}=\{\lambda^{n}ev_{\lambda}:n\geq 0\},$
it follows that the orbit of $ev_{\lambda}$ under $F^{*}$ has several non-zero limit points, and thus $(ii)$ holds. We remind that $F^{*}$ is not hypercyclic. ∎
Similarly, we have a bilateral analogue of the above result.
Proposition 5.2.
Consider the space $\ell^{p}_{a,b}(\mathtt{A})$, $1\leq p<\infty$, having a normalized Schauder basis consisting of $\{f_{n}(z)=(a_{n}+b_{n}z)z^{n}:n\in\mathbb{Z},\}$ where $a_{0}=1,$ $b_{0}=-\frac{1}{\lambda}\frac{1}{2}$and $a_{n}=1,$ $b_{n}=-\frac{1}{\lambda}\frac{n+1}{n+2},$ for all $n\geq 1,$ and $a_{n}=\frac{2^{n}}{2^{2}},$ $b_{n}=-\frac{1}{\lambda}2^{n},$ for all $n\leq-1.$ Then the following hold:
(i) If $\lambda$ is a root of the unity, then $F^{*}$ has a non-trivial periodic vector in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$, but it is not hypercyclic.
(ii) If $\lambda$ is not a root of the unity, then the zero-one law of orbital limit points fails for $F^{*}$.
Proof.
From the given values of $\{a_{n}\}_{n\in\mathbb{Z}}$ and $\{b_{n}\}_{n\in\mathbb{Z}},$ we get
$\sum_{n=-\infty}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=-1}^{-\infty}\left\lvert\left(\frac{2^{n}}{2^{2}}-\frac{1}{\lambda}2^{n}\lambda\right)\lambda^{-n}\right\rvert^{q}+\sum_{n=0}^{\infty}\left\lvert\left(1-\frac{1}{\lambda}\frac{n+1}{n+2}\lambda\right)\lambda^{n}\right\rvert^{q}<\infty.$
for $1<q<\infty.$ Here $r=\frac{1}{2}$ and $R=1.$ Therefore the natural domain of $\ell^{p}_{a,b}(\mathtt{A})$ is given by $\text{Dom}_{(p)}=\mathtt{A}\cup\{\lambda\},$ where $\mathtt{A}$ is a annulus $\frac{1}{2}<\lvert z\rvert<1.$ Since $F^{*}(ev_{z})=\lambda(ev_{z}),$ for all $z\in\mathtt{A}.$ In-particular for $z=\lambda,$ that is for the extra point in the $\text{\text{Dom}}_{(p)}$ we get
$F^{*}(ev_{\lambda})=\lambda(ev_{\lambda}),$
and hence $F^{*}$ has a non-trivial periodic vector. Now let us show that $ev_{\lambda}\in(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ But this is cleared by Closed-graph theorem. For the given values of $\{a_{n}\}_{n\in\mathbb{Z}}$ and $\{b_{n}\}_{n\in\mathbb{Z}},$ we have
$c_{-1}=-\frac{1}{2\lambda}\frac{7}{8},~{}~{}~{}~{}~{}~{}~{}c_{n}=\frac{b_{n}}{a_{n+2}}-\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}=-\frac{1}{\lambda}\frac{n+1}{n+2}+\frac{1}{\lambda}\frac{n+2}{n+3}=\frac{1}{\lambda}\frac{1}{(n+2)(n+3)},~{}~{}~{}(n\geq 0)$
and
$d_{-n}=\frac{a_{-n}}{b_{-n}}-\frac{a_{-n+1}}{b_{-n+1}}=-\frac{\lambda 2^{n}}{2^{n+2}}+\frac{\lambda 2^{n-1}}{2^{n+1}}=0,~{}~{}~{}(n\geq 2).$
Therefore, we get the matrix of $F$ corresponding to the given Schauder basis as
$[F]:=\begin{bmatrix}\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \ddots&0&\frac{\alpha\lambda^{3}}{2}\frac{7}{8}&-\frac{\alpha\lambda^{4}}{2.2}&-\frac{\alpha\lambda^{5}}{2.3}&-\frac{\alpha\lambda^{6}}{2.4}&\cdots\\ \ddots&0&\alpha\lambda^{2}\frac{7}{8}&-\frac{\alpha\lambda^{3}}{2}&-\frac{\alpha\lambda^{4}}{3}&-\frac{\alpha\lambda^{5}}{4}&\ddots\\ \ddots&\frac{1}{2}&2\alpha\lambda\frac{7}{8}&-\alpha\lambda^{2}&-\frac{2\alpha\lambda^{3}}{3}&-\frac{2\alpha\lambda^{4}}{4}&\ddots\\ \ddots&0&\frac{1}{8}(1+7\alpha)&\bm{-\frac{\alpha\lambda}{1.2}}&-\frac{\alpha\lambda^{2}}{1.3}&-\frac{\alpha\lambda^{3}}{1.4}&\ddots\\ \ddots&0&-\frac{(1-\alpha)}{2\lambda}\frac{7}{8}&1-\frac{\alpha}{2^{2}}&-\frac{\alpha\lambda}{2.3}&-\frac{\alpha\lambda^{2}}{2.4}&\ddots\\ \ddots&0&-\frac{(1-\alpha)}{3\lambda^{2}}\frac{7}{8}&\frac{(1-\alpha)}{\lambda}\frac{1}{2.3}&1-\frac{\alpha}{3^{2}}&-\frac{\alpha\lambda}{3.4}&\ddots\\ \ddots&0&-\frac{(1-\alpha)}{4\lambda^{3}}\frac{7}{8}&\frac{(1-\alpha)}{\lambda^{2}}\frac{1}{2.4}&\frac{(1-\alpha)}{\lambda}\frac{1}{3.4}&1-\frac{\alpha}{4^{2}}&\ddots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots\end{bmatrix}.$ (5.1)
Clearly, $\sum_{i=1}^{\infty}\lVert E_{-i}\rVert+\lVert D\rVert+\sum_{i=1}^{\infty}\lVert E_{i}\rVert<\infty.$ Therefore $F$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}).$ Again by the Theorem 3.10 we know that $F^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$ if and only if $\lim\inf_{n\to\infty}(\lvert a_{n}\rvert+\lvert b_{n-1}\rvert)=0$ and $\lim\inf_{n\to\infty}(\lvert a_{-n}\rvert+\lvert b_{-n-1}\rvert)=0.$ But, in our case, $\lim\inf_{n\to\infty}(\lvert a_{n}\rvert+\lvert b_{n-1}\rvert)\neq 0.$ Hence $F^{*}$ is not hypercyclic. ∎
In view of these propositions, the following corollaries are immediate.
Corollary 5.3.
There exists a weighted backward shift $B_{w}$ with respect to the standard Schauder basis in the sequence space $\ell^{p}$ and a compact operator $K$ such that $B_{w}+K$ has non-trivial periodic vectors, but $B_{w}+K$ is not hypercyclic.
There is yet another difference.
Corollary 5.4.
There exists a weighted forward shift $F_{w}$ on a space of the form $\ell^{p}_{a,b}(\mathbb{D})$, which is not similar to a forward weighted shift on $\ell^{p}$. A similar conclusion holds for bilateral forward shifts.
Acknowledgments. Bibhash Kumar Das is supported financially by CSIR research fellowship (File No.: 09/1059(0037)/2020-EMR-I), and Aneesh Mundayadan is partially funded by a Start-Up Research Grant of SERB-DST (File. No.: SRG/2021/002418).