### v2

# Weighted shift operators on spaces of analytic functions over an annulus

Abstract.
This article aims to initiate a study of the forward weighted shift operator, denoted by $F_{w}$, defined on the Banach spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. Here, $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a.b}(\mathtt{A})$, respectively, are the Banach spaces of analytic functions on a suitable annulus $\mathtt{A}$ in the complex plane, having a normalized Schauder basis of the form $\{(a_{n}+b_{n}z)z^{n},~{}~{}n\in\mathbb{Z}\}$, which is equivalent to the standard basis in $\ell^{p}(\mathbb{Z})$ and $c_{0}(\mathbb{Z})$. We first obtain necessary and sufficient conditions for $F_{w}$ to be bounded, and show that, under certain conditions on $\{a_{n}\}$ and $\{b_{n}\}$, the operator $F_{w}$ is similar to a compact perturbation of a bilateral weighted shift on $\ell^{p}(\mathbb{Z})$. This also allows us to obtain the essential spectrum of $F_{w}$. Further, we study when $F_{w}$ and its adjoint $F_{w}^{*}$ are hypercyclic, supercyclic, and chaotic, and provide a class of chaotic operators that are compact perturbations of weighted shifts on $\ell^{p}(\mathbb{Z})$. Finally, it is proved that the adjoint of a shift on the dual of $\ell^{p}_{a,b}(\mathtt{A})$ can have non-trivial periodic vectors, without being even hypercyclic. Also, the zero-one law of orbital limit points fails for $F_{w}^{*}$.
Key words and phrases:
bilateral weighted shift operator, hypercyclicity, supercyclicity, chaos, periodic vector
2010 Mathematics Subject Classification:
Primary 47A16, 46E22, 32K05, 47B32; Secondary 47B37, 37A99.
1. Introduction
Among various classes of operators that are commonly investigated in operator theory and linear dynamics, the class of weighted shifts has a predominant role. Shift operators are fairly simple to define, yet they yield diverse and appealing results. In operator theory, weighted shifts are well studied in connections with model theory (representation of various operators using shifts), invariant subspaces, cyclic vectors, subnormality, and so on, cf. Halmos [24], and Shields [36]. In the context of linear dynamics it is a usual practice that dynamical properties are first tested for weighted shifts defined on the sequence spaces $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$. Important dynamical notions such as hypercyclicity, chaos, and ergodicity are well distinguishable using weighted shifts. Motivated by these, in this paper we introduce and study weighted shift operators on the spaces $\ell^{p}_{a,b}(\mathbb{D})$ and $\ell^{p}_{a,b}(\mathtt{A})$ having a close resemblance to the weighted $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$ spaces, highlighting the similarities and differences in the dynamics. These newly introduced spaces and weighted shifts are of independent interests as well. Our main focus is on the linear dynamics of shifts.
Definition 1.1.
An operator $T$ on a separable Banach space $X$ is said to be
  • hypercyclic if there exists $x\in X$, known as a hypercyclic vector for $T$, such that the orbit $\{x,Tx,T^{2}x,\cdots\}$ is dense in $X$,
  • chaotic if $T$ is hypercyclic and the set $\{v\in X:T^{m}v=v,~{}for~{}some~{}m\in\mathbb{N}\}$ of periodic vectors is dense in $X$,
  • topologically transitive if, for two non-empty open sets $U_{1}$ and $U_{2}$ of $X$, there exists a natural number $k$ such that $T^{k}(U_{1})\cap U_{2}\neq\phi$,
  • topologically mixing on $X$ if, for any two non-empty open sets $U_{1}$ and $U_{2}$ of $X$, there exists $N$, a natural number, such that $T^{n}(U_{1})\cap U_{2}\neq\phi$ for all $n\geq N$.
We recall that the periodic vectors for an operator $T$ defined on a complex vector space are exactly the linear combinations of eigenvectors corresponding to “rational” unimodular eigenvalues.
Rolewicz [33] showed that $\lambda B$ is hypercyclic on $\ell^{p}(\mathbb{N})$, $1\leq p<\infty$, where $|\lambda|>1$ and $B$ is the unweighted backward shift operator. Also, for particular classes of hypercyclic weighted shifts, see Kitai [25], Gethner and Shapiro [18], and Godefroy and Shapiro [19]. A complete characterization of hypercyclic weighted shifts on $\ell^{p}(\mathbb{N})$ and $\ell^{p}(\mathbb{Z})$ was obtained by Salas [34]. In [19] the authors introduced and studied the notion of chaos for operators, and established that several familiar operators including weighted shifts are chaotic. A complete characterization of hypercyclic and chaotic weighted shifts in a general $F$-spaces was obtained by Grosse-Erdmann [22]. We also note that the weighted shifts which are (topologically) mixing were characterized by Costakis and Sambarino [12]. For an account on the theory on hypercyclicity and related topics, we refer to the monographs by F. Bayart and E. Matheron [4], and K.-G. Grosse-Erdmann and A. Peris [23]. Also, see S. Grivaux et al., [21]. In practice, the existence of hypercyclic and chaotic properties of operators is established using well known criteria that are stated below, cf. [4] and [23]. We will make use of these standard criteria to obtain the hypercyclic and chaotic properties of weighted shifts defined on the weighted spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ which are introduced in Section $2$.
Theorem 1.2.
(Gethner-Shapiro Criterion [18]) Let $T$ be a bounded operator on a separable Banach space $X$, and let $D$ be a dense subset of $X$. If $\{n_{k}\}\subseteq\mathbb{N}$ is a strictly increasing sequence and $S:D\rightarrow D$ is a map such that, for each $x\in D$, $lim_{k\rightarrow\infty}T^{n_{k}}x=0=\lim_{k\rightarrow\infty}S^{n_{k}}x,$ and $TSx=x$, then $T$ is hypercyclic. In addition, if $n_{k}=k$ for all $k\geq 1$, then $T$ is mixing on $X$.
A similar criterion, known as the chaoticity criterion, has been used to obtain chaotic operators in $F$-spaces, and we refer to [9] where it is originally stated. This criterion is very strong: indeed, if an operator $T$ satisfies the chaoticity criterion, then $T$ is topologically mixing, chaotic, frequently hypercyclic, and it has invariant measures of full support, cf. [4].
Theorem 1.3.
(Chaoticity Criterion [9]) Let $X$ be a separable Banach space, $D$ be a dense set in $X$, and let $T$ be a bounded operator on $X$. If there exists a map $S:D\rightarrow D$ such that $\sum_{n\geq 0}T^{n}x$ and $\sum_{n\geq 0}S^{n}x$ are unconditionally convergent, and $TSx=x$, for each $x\in D$, then the operator $T$ is chaotic and mixing on $X$.
An operator $T$ on a separable Banach space $X$ over $\mathbb{K}:=\mathbb{R}$ or $\mathbb{C}$ is supercyclic if the projective orbit $\{\alpha T^{n}(x):~{}\alpha\in\mathbb{K},n\geq 0\}$ is dense in $X$. Supercyclicity was initially studied by Hilden and Wallen [26] who provided classes of supercyclic weighted shifts. A complete characterization of supercyclic unilateral and bilateral weighted shifts was obtained by Salas [35], and further studied by Montes-Rodriguez and Salas [31]. Along the same lines of the hypercyclicity and chaoticity criteria, a general sufficient condition for supercyclicity is well-known, as stated below.
Theorem 1.4.
(Supercyclicity Criterion [31]) For a bounded operator $T$ on a separable Banach space $X$, if there exists a dense subset $D$, a sequence $(n_{k})$ of natural numbers, and a map $S:D\rightarrow D$ such that, for all $x,y\in D$, $\lim_{k\rightarrow\infty}\|T^{n_{k}}x\|\|S^{n_{k}}y\|=0$ and $TSx=x$, then $T$ is supercyclic.
Motivated by the classical bilateral shifts, in this paper we initiate a study of weighted bilateral forward shift operators $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. In Section $2$, we introduce the aforementioned spaces as analytic function spaces and prove that the co-ordinate functionals on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ are continuous. In Section $3,$ we obtain necessary and sufficient conditions for the boundedness of $F_{w}$, and as main results, we study when $F_{w}$ and its adjoint $F_{w}^{*}$ are hypercyclic, mixing, supercyclic, and chaotic. Dynamics of the adjoint of a unilateral forward shift is studied in Section $4$. In Section $5$, it is shown that the bilateral shifts on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ and also the unilateral shifts have distinguishing properties compared to classical weighted shifts . For example, among other results, it is shown that the unweighted unilateral and bilateral forward shifts $F$ such that $F^{*}$ can admit a non-zero periodic vector, without being hypercyclic. It is also shown that the zero-one law of orbital limit points for the adjoint $F^{*}$.
2. The spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$: evaluation and co-ordinate functionals
We first recall the notion of equivalent Schauder bases in Banach spaces, cf. Lindenstrauss and Tzafriri [28], page $5$. Let $X$ and $Y$ be Banach spaces. We say that the Schauder bases $\{u_{n}\}_{n=-\infty}^{\infty}$ of $X$ and $\{v_{n}\}_{n=-\infty}^{\infty}$ of $Y$, are equivalent if the convergence of $\sum_{n=-\infty}^{\infty}a_{n}u_{n}$ in $X$ is equivalent to that of $\sum_{n=-\infty}^{\infty}a_{n}v_{n}$ in $Y$.
Let $a:=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b:=\{b_{n}\}_{n=-\infty}^{\infty}$ be complex sequences, $a_{n}\neq 0$, and $b_{n}\neq 0$, for all $n\in\mathbb{Z}$. Define $\ell^{p}_{a,b}(\mathtt{A})$ to be the space of all analytic functions on an annulus $\mathtt{A}$, having a normalized Schauder basis $\{f_{n}\}_{n=-\infty}^{\infty}$, where
$f_{n}(z):=(a_{n}+b_{n}z)z^{n},\hskip 28.45274pt(n\in\mathbb{Z}),$
and it is assumed to be equivalent to the standard ordered basis in $\ell^{p}(\mathbb{Z})$, $1\leq p<\infty$, that is,
$f=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}\in\ell^{p}_{a,b}(\mathtt{A}),$ (2.1)
if and only if $\lVert f\rVert_{\ell^{p}_{a,b}(\mathtt{A})}:=\left(\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}\rvert^{p}\right)^{\frac{1}{p}}<\infty.$ Note that for $b_{n}=0$ for all $n\in\mathbb{Z},$ we get a space that is equivalent to a weighted $\ell^{p}(\mathbb{Z})$. In a similar way, we define the space $c_{0,a,b}(\mathtt{A})$ which consists of analytic functions, for which $\{(a_{n}+b_{n}z)z^{n}:n\in\mathbb{Z}\}$ forms a normalized Schauder basis, equivalent to the usual basis in $c_{0}(\mathbb{Z})$. Moreover, for $f=\sum_{n}\lambda_{n}f_{n}$ in $c_{0,a,b}(\mathtt{A})$, we have $\lambda_{n}\rightarrow 0$, as $|n|\rightarrow\infty$. Define its norm as
$\|f\|_{c_{0,a,b}(\mathtt{A})}:=\sup_{n\in\mathbb{Z}}|\lambda_{n}|.$
We can similarly consider the spaces $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ over the unit disc $\mathbb{D}$ of the complex plane.
The following theorem shows the annulus $\mathtt{A}$ occurs naturally, which depends on $p$, $\{a_{n}\}$ and $\{b_{n}\}$. The series representation of $f$ as in (2.1) can be rearranged to a convergent Laurent series, and the evaluation functionals are continuous on $\mathtt{A}$. Moreover, each function in $\ell^{p}_{a,b}(\mathtt{A})$ as well as $c_{0,a,b}(\mathtt{A})$, is analytic on the natural annulus $\mathtt{A}$.
Theorem 2.1.
Consider the spaces $\ell^{p}_{a.b}(\mathtt{A}),$ $1\leq p<\infty,$ and $c_{0,a,b}(\mathtt{A})$, defined as above corresponding to $f_{n}(z)=(a_{n}+b_{n}z)z^{n}$, $n\in\mathbb{Z}$. Set
$\mathtt{A}_{1}:=\left\{\zeta\in\mathbb{C}:\sum_{n=-\infty}^{\infty}\big{(}(\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert\zeta\rvert^{n}\big{)}^{q}<\infty\right\}$,
where $\frac{1}{p}+\frac{1}{q}=1,$
$\mathtt{A}_{2}:=\left\{\zeta\in\mathbb{C}:\sup_{n\in\mathbb{Z}}~{}(\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert\zeta\rvert^{n}<\infty\right\},$
and
$\mathtt{A}_{3}:=\left\{\zeta\in\mathbb{C}:\sum_{n=-\infty}^{\infty}(\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert\zeta\rvert^{n}<\infty\right\}.$
The following hold.
  1. (i)
    If $f=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}\in\ell^{p}_{a,b}(\mathtt{A}),$ $1<p<\infty,$ then the series $\sum_{n=-\infty}^{\infty}\lambda_{n}(a_{n}+b_{n}z)z^{n}$ converges uniformly and absolutely on compact subsets of $\mathtt{A}_{1}$.
    The same conclusions hold true with respect to the domains $\mathtt{A}_{2}$ and $\mathtt{A}_{3}$, respectively, for the spaces $\ell^{1}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$.
  2. (ii)
    The evaluation functionals $ev_{\zeta}$ given by $f\mapsto f(\zeta)$ are bounded on $\ell^{p}_{a.b}(\mathtt{A}),$ and $c_{0,a,b}(\mathtt{A})$, where $1\leq p<\infty$, and $\zeta\in\mathtt{A}$.
Proof.
We prove the result for the case $1<p<\infty$ only. Assume that $f=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}$ converges in $\ell^{p}_{a,b}(\mathtt{A}),$ and $\zeta\in\mathtt{A}_{1}.$ Fix a closed ball $K$ that lies in $\mathtt{A}_{1}$. Let $\zeta\in K$. We now get
$\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}(a_{n}+b_{n}\zeta)\zeta^{n}\rvert$$\leq$ $\text{max}\{1,\lvert\zeta\rvert\}\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}\rvert(\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert\zeta\rvert^{n}$
$\leq$ $\max_{\zeta\in K}\{1,\lvert\zeta\rvert\}\left(\sum_{n=-\infty}^{\infty}(\lvert a_{n}\rvert+\lvert b_{n}\rvert)^{q}\lvert\zeta\rvert^{qn}\right)^{\frac{1}{q}}\left(\sum_{n=-\infty}^{\infty}\lvert\lambda_{n}\rvert^{p}\right)^{\frac{1}{p}}$
$=$ $M\lVert f\rVert,$
[…]
Further, we have that $|f(\zeta)|\leq M\|f\|$, and hence, evaluation functionals are continuous. (In fact, we have actually established that $\sup_{\zeta\in K}|f(\zeta)|\leq M_{K}\|f\|$ for every compact subset $K$ in $\mathtt{A}$, where the constant $M_{K}$ depends only on $K$.) ∎
A few comments are in order. The natural domains $\mathtt{A}_{j}$, $j=1,2,3$, will always be assumed to be non-degenerate. Secondly, in the case of the usual weighted spaces, the natural domains of the map $z\mapsto ev_{z}$ are open or closed annuli. However, this is not so for $\ell^{p}_{a,b}(\mathtt{A})$ which is an important property that distinguishes these spaces from the classical weighted $\ell^{p}$ spaces, which will be discussed and utilized in Section $5$. Also, the series $\sum_{n=-\infty}^{\infty}\lambda_{n}(a_{n}+b_{n}z)z^{n}$ can still converge at a point $z$ that lies outside $\mathtt{A}_{j}$, and hence the domain of $z\mapsto ev_{z}$ can be extended to “extra points” outside $\mathtt{A_{j}}$, cf. Section $5$. We note that such a phenomenon in the unilateral case when $p=2$, was observed by Adams and McGuire [2] who introduced the space the analytic reproducing kernel Hilbert space $\ell^{2}_{a,b}$ having an orthonormal basis of the form $(a_{n}+b_{n}z)z^{n}$, $n\geq 0$.
Next, we get the continuity of the co-ordinate functionals which are defined as follows. Since each $f(z)$ in $\ell^{p}_{a,b}(\mathtt{A})$ is analytic on the annulus $\mathtt{A}$, we have the Laurent expansion $f(z)=\sum_{n=-\infty}^{\infty}\widehat{f}(n)z^{n}$, $z\in\mathtt{A}$, where $\widehat{f}(n)$ is the $n$-th Laurent coefficient of $f(z)$. The $n$-th co-ordinate functional (or rather, a coefficient functional) $k_{n}$ is given by
$k_{n}(f)=\widehat{f}(n),$
and it is defined on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. Also, let $f_{n}^{*}$ denote co-ordinate functional with respect to the basis $\{f_{n}\}$ in $\ell^{p}_{a,b}(\mathtt{A})$, $1\leq p<\infty$, or $c_{0,a,b}(\mathtt{A})$, that is,
$f_{n}^{*}(f)=\lambda_{n},\hskip 17.07182ptf=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}.$
(We add here that, both $k_{n}$ and $f_{n}$ will be utilized in understanding the dynamics of a weighted forward shift $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$. Indeed, the adjoint $F_{w}^{*}$ will “act like a weighted shift” with respect to $k_{n}$, see Proposition 3.1.)
Proposition 2.2.
Let $k_{n}$ denote the co-ordinate functional $f\mapsto\widehat{f}(n)$, defined on $\ell^{p}_{a,b}(\mathtt{A})$ or $c_{0,a,b}(\mathtt{A})$, where $n\in\mathbb{Z}$. Then the following hold.
  • (i)
    For $1<p<\infty$, $11/p+1/q=$ 1,
    $\|k_{n}\|\leq(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q}.$
    For $p=1$, we have $\|k_{n}\|\leq\max\{|a_{n}|,|b_{n-1}|\}$. Also, for the case of $c_{0,a,b}(\mathtt{A})$, $\|k_{n}\|\leq|a_{n}|+|b_{n-1}|$.
  • (ii)
    $\{f_{n},f_{n}^{*}\}$ is a biorthogonal system, and for $1<p<\infty$, $\{f_{n}^{*}\}$ is an unconditional Schauder basis for the dual space.
  • (iii)
    For $n\in\mathbb{Z}$, we have $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$.
Proof.
$(i)$ If $f\in\ell^{p}_{a,b}(\mathtt{A}),$ then $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}(a_{k}+b_{k}z)z^{k}$, and $\|f\|=\left(\sum_{k=-\infty}^{\infty}\lvert\lambda_{k}\rvert^{p}\right)^{1/p}$. A rearrangement into a Laurent series yields $f(z)=\sum_{k=-\infty}^{\infty}(\lambda_{k}a_{k}+\lambda_{k-1}b_{k-1})z^{k},$ and so, $k_{n}(f)=\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}.$ It follows, by Hölder inequality, that
$|k_{n}(f)|\leq\|f\|_{\ell^{p}_{a,b}(\mathtt{A})}(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q},$
and hence $\|k_{n}\|\leq(|a_{n}|^{q}+|b_{n-1}|^{q})^{1/q}$.
$(ii)$ This is immediate as $f_{n}^{*}(f_{n})=1$, and $f_{n}^{*}(f_{m})=0$ for $n\neq m$. Also, the equivalence of $\{f_{n}\}$ to the standard basis $\{e_{n}\}$ shows that $\{f_{n}^{*}\}$ and $\{e_{n}^{*}\}$ are equivalent bases which follow from reflexivity.
$(iii)$ For $f\in\ell^{p}_{a,b}(\mathtt{A})$ and $n\in\mathbb{Z}$, we have
$(a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*})(f)$ $=$ $a_{n}f_{n}^{*}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}\big{)}+b_{n-1}f_{n-1}^{*}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}\big{)}$
$=$ $\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}=k_{n}(f).$
We recall that the unilateral versions $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ were introduced and studied in [13]. These spaces correspond to the normalized bases $(a_{n}+b_{n}z)z^{n}$, where $n\geq 0$. The results in Proposition 2.2 hold for $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ as well.
The following is a slight generalization on the continuity of the co-ordinate functionals $k_{n}$ defined on analytic function spaces over an annulus and disc centered at the origin.
Proposition 2.3.
Let $\mathcal{E}$ be a Banach space of analytic functions on an annulus $\mathtt{A}$ centered around 00, having evaluation functionals bounded at each $z\in\mathtt{A}$. Then, $k_{n}$ is bounded for all $n\in\mathbb{Z}$. Further,
$\text{span}~{}\big{\{}k_{n}:~{}n\in\mathbb{Z}\big{\}}$
is dense in $\mathcal{E}^{*}$ with respect to the weak-topology, and also, in the norm topology of $\mathcal{E}^{*}$ if it is reflexive.
Proof.
Let $ev_{z}$ denote the evaluation functional acting on $\mathcal{E}$. Since the analyticity of $z\mapsto ev_{z}$ in the strong and norm operator topologies are equivalent, we see that $z\mapsto ev_{z}$ is an $\mathcal{E}^{*}$-valued norm-analytic function. It then admits a norm-convergent Laurent series: $ev_{z}=\sum_{n=-\infty}^{\infty}L_{n}z^{n},\hskip 14.22636pt(z\in\mathtt{A}),$ for some $L_{n}\in\mathcal{E}^{*}$, $n\in\mathbb{Z}$. On the other hand, if $f\in\mathcal{E}$, we have $f(z)=ev_{z}(f)=\sum_{n=-\infty}^{\infty}L_{n}(f)z^{n}.$ Expanding $f(z)$ in its Laurent series, we must then have $L_{n}=k_{n}$ for all $n$.
To show that $\{k_{n}\}$ is a weak-total set in $\mathcal{E}^{*}$, let $\Lambda\in\mathcal{E}^{**}$ be such that $\Lambda(k_{n})=0$ for all $n$. Hence, there is some $f\in\mathcal{E}$ such that $\Lambda(k_{n})=k_{n}(f)=0,$ $\forall n\in\mathbb{Z}$, which shows that all the coefficients of $f(z)$ vanish. Consequently, $f=0$.
We now obtain conditions so that the Laurent monomials $z^{\nu}$ belong to the space $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. The norm estimates, stated below, will be later used in deducing the dynamical properties of weighted shifts.
Proposition 2.4.
Consider the spaces $\ell^{p}_{{a},{b}}(\mathtt{A}),$ $1\leq p<\infty$, and $c_{0,a,b}(\mathtt{A})$. Assume that
$\limsup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}\limsup_{n\rightarrow+\infty}\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<1,$
then there exist constants $M_{1},M_{2}>0$ such that
$\lVert z^{\nu}\rVert\leq M_{1}\prod_{k=0}^{\nu-1}\left|\frac{a_{k}}{b_{k}}\right|~{}~{}~{}\text{and}~{}~{}~{}\lVert z^{-\nu}\rVert\leq M_{2}~{}~{}\prod_{k=1}^{\nu}\left|\frac{b_{-k}}{a_{-k}}\right|,$ (2.2)
for all $\nu\in\mathbb{N}.$ The same holds for $c_{0,a,b}(\mathtt{A})$ as well. In particular, the Laurent polynomials form a dense subspace in $c_{0,a,b}(\mathtt{A})$ and $\ell^{p}_{{a},{b}}(\mathtt{A})$.
Proof.
Since $\lim\sup_{n\rightarrow+\infty}\big{|}\frac{b_{n}}{a_{n+1}}\big{|}<1,$ we can find $r<1$ and an integer $N_{1}\geq 1,$ such that $\big{|}\frac{b_{n}}{a_{n+1}}\big{|}<r,$ for all $n\geq N_{1}.$ Also, as $\lim\sup_{n\rightarrow+\infty}\big{|}\frac{a_{-n+1}}{b_{-n}}\big{|}<1$, there exists $R<1$ such that $\left|\frac{a_{-n+1}}{b_{-n}}\right|<R$, for all $n\geq N_{2}$.
The space $\ell^{p}_{a,b}(\mathtt{A})$ contains the constant functions. Indeed, by the basis expansion in $\ell^{p}_{{a},{b}}(\mathtt{A})$ and a rearrangement into Laurent series, we can find some $\{\lambda_{j}\}_{j=-\infty}^{\infty}\in\ell^{p}(\mathbb{Z})$ such that
$1=\sum_{j=-\infty}^{\infty}\lambda_{j}(a_{j}z^{j}+b_{j}z^{j+1})=\sum_{j=-\infty}^{\infty}\left(\lambda_{j}a_{j}+\lambda_{j-1}b_{j-1}\right)z^{j},$ (2.3)
for all $z$ in an annulus around 00. Now equating the coefficients of like-powers, we have
$\lambda_{j}=\begin{cases}(-1)^{j}\lambda_{0}\frac{b_{0}b_{1}b_{2}\cdots b_{j-1}}{a_{1}a_{2}a_{3}\cdots a_{j}},\hskip 88.2037ptj\geq 1\\ \lambda_{0},\hskip 165.02606ptj=0,\\ \frac{1}{b_{-1}}\left(1-\lambda_{0}a_{0}\right),\hskip 105.2751ptj=-1,\\ (-1)^{-j-1}\frac{1}{b_{-1}}\left(1-\lambda_{0}a_{0}\right)\frac{a_{-1}a_{-2}\cdots a_{j+1}}{b_{-2}b_{-3}\cdots b_{j}},~{}~{}j\leq-2.\\ \end{cases}$
By the choices of $N_{1}$ and $N_{2}$, we separately obtain the convergence of $\sum_{j>0}|\lambda_{j}|^{p}$ and $\sum_{j<0}|\lambda_{j}|^{p}$, and hence $1\in\ell^{p}_{a,b}(\mathtt{A})$.
For a fixed $\nu\geq 1,$
$a_{0}$ $=$ $f_{0}-b_{0}z=f_{0}-\frac{b_{0}}{a_{1}}f_{1}+\frac{b_{0}b_{1}}{a_{1}}z^{2}$
$=$ $f_{0}+\sum_{j=1}^{\nu-1}(-1)^{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{1}a_{2}\cdots a_{j}}f_{j}+(-1)^{\nu}\frac{b_{0}b_{1}\cdots b_{\nu-1}}{a_{1}a_{2}\cdots a_{\nu-1}}z^{\nu}.$
It follows that by using the expansion of the constant function $1$, we obtain
$z^{\nu}=(-1)^{\nu}\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\left(\sum_{j=-\infty}^{\infty}\lambda_{j}f_{j}\right)-(-1)^{\nu}\left(\frac{a_{1}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}f_{0}-\cdots+(-1)^{\nu-1}\frac{1}{b_{\nu-1}}f_{\nu-1}\right).$
Thus, we get
$z^{\nu}$ $=$ $(-1)^{\nu}\frac{1-\lambda_{0}a_{0}}{b_{-1}}~{}\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{\nu}(\lambda_{0}a_{0}-1)\frac{a_{0}\cdots a_{\nu-1}}{b_{0}\cdots b_{\nu-1}}\left(\frac{1}{a_{0}}f_{0}+\sum_{j=1}^{\nu-1}(-1)^{j}\frac{b_{0}\cdots b_{j-1}}{a_{0}\cdots a_{j}}f_{j}\right)$
$+\frac{\lambda_{0}a_{0}}{a_{\nu}}\left(f_{\nu}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{\nu}b_{\nu+1}\cdots b_{\nu+j-1}}{a_{\nu+1}a_{\nu+2}\cdots a_{\nu+j}}f_{\nu+j}\right),$
which is the expansion of $z^{\nu}$ in terms of the basis $\{f_{j}\}$. Therefore, for $\nu\geq 1$,
$\left\lVert z^{\nu}\right\rVert^{p}$ $=$ $\left\lvert(1-\lambda_{0}a_{0})\prod_{k=0}^{\nu-1}\frac{a_{k}}{b_{k}}\right\rvert^{p}\left(\left\lvert\frac{1}{b_{-1}}\right\rvert^{p}+\sum_{j=2}^{\infty}\left\lvert\frac{1}{b_{-1}}\prod_{k=1}^{j-1}\frac{a_{-k}}{b_{-k-1}}\right\rvert^{p}+\left\lvert\frac{1}{a_{0}}\right\rvert^{p}+\sum_{j=1}^{\nu-1}\left\lvert\frac{1}{a_{0}}\prod_{k=1}^{j}\frac{b_{k-1}}{a_{k}}\right\rvert^{p}\right)$
$+\left\lvert\frac{\lambda_{0}a_{0}}{a_{\nu}}\right\rvert^{p}\left(1+\sum_{j=1}^{\infty}\left\lvert\prod_{k=0}^{j-1}\frac{b_{\nu+k}}{a_{\nu+k+1}}\right\rvert^{p}\right).$
We immediately calculate the norm estimates of $z^{\nu}$ using the choices of $N_{1}$ and $N_{2}$. Thus, there exists a constant $M_{1}>0$ such that
$\left\lVert z^{\nu}\right\rVert_{\ell^{p}_{{a},{b}}(\mathtt{A})}\leq M_{1}\left\lvert\frac{a_{0}a_{1}\cdots a_{\nu-1}}{b_{0}b_{1}\cdots b_{\nu-1}}\right\rvert,$
for all $\nu\geq N_{1}$. This is the first estimate in the proposition.
To get the second estimate, we write
$b_{-1}$ $=$ $f_{-1}-a_{-1}z^{-1}$
$=$ $f_{-1}-\frac{a_{-1}}{b_{-2}}f_{-2}+\frac{a_{-1}a_{-2}}{b_{-2}}z^{-2}$
$=$ $f_{-1}+\sum_{j=2}^{\nu}(-1)^{j-1}\frac{a_{-1}a_{-2}\cdots a_{-j+1}}{b_{-2}b_{-3}\cdots b_{-j}}f_{-j}+(-1)^{\nu}\frac{a_{-1}\cdots a_{-\nu}}{b_{-2}\cdots b_{-\nu}}z^{-\nu},$
for a fixed $\nu>1.$ It follows that by using the expansion of the constant $1,$
$z^{-\nu}=(-1)^{\nu}\frac{b_{-1}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}\left(\sum_{j=-\infty}^{\infty}\lambda_{j}f_{j}\right)-(-1)^{\nu}\left(\frac{b_{-2}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}f_{-1}-\cdots+(-1)^{\nu-1}\frac{1}{a_{-\nu}}f_{-\nu}\right).$
Thus, we get
$z^{-\nu}$ $=$ $\frac{1-\lambda_{0}a_{0}}{b_{-\nu-1}}\left(f_{-\nu-1}+\sum_{j=0}^{\infty}(-1)^{j+1}\frac{a_{-\nu-1}\cdots a_{-\nu-1-j}}{b_{-\nu-2}\cdots b_{-\nu-2-j}}f_{-\nu-2-j}\right)$
$+(-1)^{\nu}\lambda_{0}a_{0}\frac{b_{-1}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}\left(\frac{-1}{b_{-1}}f_{-1}+\sum_{j=2}^{\nu}(-1)^{j}\frac{a_{-1}\cdots a_{-j+1}}{b_{-1}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{\nu}\lambda_{0}\frac{b_{-1}\cdots b_{-\nu}}{a_{-1}\cdots a_{-\nu}}\left(f_{0}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{1}a_{2}\cdots a_{j}}f_{j}\right),$
which is the expansion of $z^{-\nu}$ in terms of the basis $\{f_{j}\}.$ Computing the norm $\|z^{-\nu}\|$ using the choices of $N_{1}$ and $N_{2},$ the second estimate in the proposition holds. ∎
3. Bilateral weighted shifts $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$
In this section we introduce the bilateral weighted forward shift operator $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$. We find boundedness conditions for $F_{w}$. The essential spectrum of $F_{w}$ is obtained via a compact perturbation result. Further, hypercyclicity, supercyclicity, and chaos for both $F_{w}$ and $F_{w}^{*}$ are studied along with examples.
Let $w:=\{w_{n}\}$ be a complex sequence. The operator $F_{w}$ is defined by
$F_{w}(\sum_{n=-\infty}^{\infty}\lambda_{n}z^{n}):=\sum_{n=-\infty}^{\infty}\lambda_{n}w_{n}z^{n+1}.$
So, we have $F_{w}(z^{n})=w_{n}z^{n+1}$ for all $n\in\mathbb{Z}$. We isolate certain general properties of $F_{w}$ which are of independent interest. Among them is an important property of weighted forward shifts, which loosely says that the adjoint $F_{w}^{*}$ whenever $F_{w}$ is bounded, “acts like” a weighted backward shift with respect to a total set, as proved below. The unweighted case of $F$ defined on reproducing kernel space over the unit disc was discussed in [32].
Proposition 3.1.
Let $\mathcal{E}$ be a Banach space of analytic functions defined on some annulus $r<|z|<R$, having continuous evaluation functionals. Let $F_{w}$ be bounded on $\mathcal{E}$. Then,
$F_{w}^{*}(k_{n})=w_{n-1}k_{n-1},$
for all $n\in\mathbb{Z},$ where $k_{n}$ denotes the $n$-th co-ordinate functional.
Proof.
We have already showed that $k_{n}\in\mathcal{E}^{*}$ for all $n$. For a function $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}z^{k}$ in $\mathcal{E}$, we have $k_{n-1}(f)=k_{n-1}\left(\sum_{k=-\infty}^{\infty}\lambda_{k}z^{k}\right)=\lambda_{n-1}.$ On the other hand,
$F_{w}^{*}\big{(}k_{n}\big{)}(f)$ $=$ $\big{(}k_{n}\circ F_{w}\big{)}(f)$
$=$ $k_{n}\big{(}\sum_{k=-\infty}^{\infty}\lambda_{k}w_{k}z^{k+1}\big{)}=\lambda_{n-1}w_{n-1}.$
An analogue for the unilateral weighted forward shift $F_{w}$ is stated below, using the same notation $k_{n}$ for $n$-th co-ordinate functional, $n\geq 0$. We omit its proof.
Proposition 3.2.
Let $F_{w}$ be a bounded forward weighted shift on a Banach space $\mathcal{E}$ consisting of analytic functions on the unit disc $\mathbb{D}$, having continuous evaluation functionals. Then,
$F_{w}^{*}(k_{n})=\begin{cases}w_{n-1}k_{n-1},\hskip 14.22636ptn\geq 1,\\ 0,\hskip 56.9055ptn=0.\end{cases}$
3.1. The matrix of $F_{w}$, boundedness, compact perturbations, and the essential spectrum
To derive some necessary and sufficient conditions for the weighted forward shift $F_{w}$ to be bounded, we compute the matrix representation of $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ with respect to the ordered normalized Schauder basis $f_{n}(z)=(a_{n}+b_{n}z)z^{n},n\in\mathbb{Z}$ in $\ell^{p}_{a,b}(\mathtt{A})$. Then, the boundedness of $F_{w}$ is equivalent to that of its matrix operator acting on $\ell^{p}(\mathbb{Z})$.
We will use the basis expansions of $z^{n}$ given in the proof of Proposition 2.4. To find the matrix of $F_{w}$, we have
$F_{w}(f_{n})(z)=w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}=\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+\left(\frac{w_{n+1}b_{n}}{a_{n+2}}-\frac{w_{n}a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right)a_{n+2}z^{n+2},$
by substituting $z^{n+1}=\frac{1}{a_{n+1}}(f_{n+1}-b_{n+1}z^{n+2})$. Set
$c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}.$
In the following, we now get the $n$-th column of the matrix of $F_{w}$, for $n\geq-1$, by using the expansion of $z^{n+2}$ in terms of the basis elements $f_{n}$’s:
$F_{w}(f_{n})(z)$ $=$ $\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+c_{n}a_{n+2}z^{n+2}$
$=$ $\left(\frac{w_{n}a_{n}}{a_{n+1}}+\alpha c_{n}\frac{a_{n+2}}{b_{n+1}}\right)f_{{n+1}}$
$+(-1)^{n+2}\alpha c_{n}\frac{a_{0}\cdots a_{n+2}}{b_{-1}\cdots b_{n+1}}\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+(-1)^{n+3}\alpha c_{n}\left(\sum_{j=1}^{n+1}(-1)^{j-1}\frac{a_{n+2}a_{n+1}\cdots a_{j}}{b_{n+1}b_{n}\cdots b_{j-1}}f_{j-1}\right)$
$+(1-\alpha)c_{n}\left(f_{n+2}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{n+2}b_{n+3}\cdots b_{n+1+j}}{a_{n+3}a_{n+4}\cdots a_{n+2+j}}f_{n+2+j}\right).$
Similarly, by substituting the expression $z^{n+2}=\frac{1}{b_{n+1}}(f_{n+1}-a_{n+1}z^{n+1})$, we find that
$F_{w}(f_{n})(z)=w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}=\frac{w_{n+1}b_{n}}{b_{n+1}}f_{{n+1}}+\left(\frac{w_{n}a_{n}}{b_{n}}-\frac{w_{n+1}a_{n+1}}{b_{n+1}}\right)b_{n}z^{n+1}.$
Setting
$d_{n}:=w_{n}\frac{a_{n}}{b_{n}}-w_{n+1}\frac{a_{n+1}}{b_{n+1}},$
we can now get the remaining columns of $[F_{w}]$, as follows.
$F_{w}(f_{-n})(z)$ $=$ $\frac{w_{-n+1}b_{-n}}{b_{-n+1}}f_{{-n+1}}+d_{-n}b_{-n}z^{-n+1}$
$=$ $\left(\frac{w_{-n+1}b_{-n}}{b_{-n+1}}+(1-\alpha)d_{-n}\frac{b_{-n}}{a_{-n+1}}\right)f_{-n+1}$
$+\alpha d_{-n}\left(f_{-n}+\sum_{j=0}^{\infty}(-1)^{j+1}\frac{a_{-n}a_{-n-1}\cdots a_{-n-j}}{b_{-n-1}b_{-n-2}\cdots b_{-n-j-1}}f_{-n-j-1}\right)$
$+(-1)^{n-1}(1-\alpha)d_{-n}\left(\sum_{j=1}^{n-2}(-1)^{j}\frac{b_{-n}b_{-n+1}\cdots b_{-j-1}}{a_{-n+1}a_{-n+2}\cdots a_{-j}}f_{-j}\right)$
$+(-1)^{n-1}d_{-n}\lambda_{0}\frac{b_{-1}\cdots b_{-n}}{a_{-1}\cdots a_{-n+1}}\left(f_{0}+\sum_{j=1}^{\infty}(-1)^{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{1}a_{2}\cdots a_{j}}f_{j}\right),$
where $n\geq 2$, and $\alpha=1-\lambda_{0}a_{0}.$ Hence, the (bilateral) matrix $[F_{w}]$ of the operator $F_{w}$ with respect to the ordered normalized Schauder basis $\{f_{n}\}_{n\in\mathbb{Z}}$ is given by
$[F_{w}]=\begin{bmatrix}\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots\\ \ddots&\frac{\alpha d_{-2}a_{-3}a_{-2}}{b_{-4}b_{-3}}&\frac{\alpha c_{-1}a_{-3}\cdots a_{1}}{b_{-4}\cdots b_{0}}&\frac{-\alpha c_{0}a_{-3}\cdots a_{2}}{b_{-4}\cdots b_{1}}&\ddots&\ddots&\ddots\\ \ddots&\frac{-\alpha d_{-2}a_{-2}}{b_{-3}}&\frac{-\alpha c_{-1}a_{-2}\cdots a_{1}}{b_{-3}\cdots b_{0}}&\frac{\alpha c_{0}a_{-2}\cdots a_{2}}{b_{-3}\cdots b_{1}}&\frac{-\alpha c_{1}a_{-2}\cdots a_{3}}{b_{-3}\cdots b_{2}}&\ddots&\ddots\\ \ddots&\alpha d_{-2}&\frac{\alpha c_{-1}a_{-1}a_{0}a_{1}}{b_{-2}b_{-1}b_{0}}&\frac{-\alpha c_{0}a_{-1}\cdots a_{2}}{b_{-2}\cdots b_{1}}&\frac{\alpha c_{1}a_{-1}\cdots a_{3}}{b_{-2}\cdots b_{2}}&\frac{-\alpha c_{2}a_{-1}\cdots a_{4}}{b_{-2}\cdots b_{3}}&\ddots\\ \ddots&\gamma_{-2}&\frac{-\alpha c_{-1}a_{0}a_{1}}{b_{-1}b_{0}}&\frac{\alpha c_{0}a_{0}a_{1}a_{2}}{b_{-1}b_{0}b_{1}}&\frac{-\alpha c_{1}a_{0}\cdots a_{3}}{b_{-1}\cdots b_{2}}&\frac{\alpha c_{2}a_{0}a_{1}\cdots a_{4}}{b_{-1}\cdots b_{3}}&\ddots\\ \ddots&\frac{-\lambda_{0}d_{-2}b_{-1}b_{-2}}{a_{-1}}&\gamma_{-1}&\bm{\frac{-\alpha c_{0}a_{1}a_{2}}{b_{0}b_{1}}}&\frac{\alpha c_{1}a_{1}a_{2}a_{3}}{b_{0}b_{1}b_{2}}&\frac{-\alpha c_{2}a_{1}\cdots a_{4}}{b_{0}\cdots b_{3}}&\ddots\\ \ddots&\frac{\lambda_{0}d_{-2}b_{-2}b_{-1}b_{0}}{a_{-1}a_{1}}&(1-\alpha)c_{-1}&\gamma_{0}&\frac{-\alpha c_{1}a_{2}a_{3}}{b_{1}b_{2}}&\frac{\alpha c_{2}a_{2}a_{3}a_{4}}{b_{1}b_{2}b_{3}}&\ddots\\ \ddots&\frac{-\lambda_{0}d_{-2}b_{-2}\cdots b_{1}}{a_{-1}a_{1}a_{2}}&\frac{-(1-\alpha)c_{-1}b_{1}}{a_{2}}&(1-\alpha)c_{0}&\gamma_{1}&\frac{-\alpha c_{2}a_{3}a_{4}}{b_{2}b_{3}}&\ddots\\ \ddots&\ddots&\frac{(1-\alpha)c_{-1}b_{1}b_{2}}{a_{2}a_{3}}&\frac{-(1-\alpha)c_{0}b_{2}}{a_{3}}&(1-\alpha)c_{1}&\gamma_{2}&\ddots\\ \ddots&\ddots&\ddots&\frac{(1-\alpha)c_{0}b_{2}b_{3}}{a_{3}a_{4}}&\frac{-(1-\alpha)c_{1}b_{3}}{a_{4}}&(1-\alpha)c_{2}&\ddots\\ \ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots\end{bmatrix}.$
Here, the expression in bold refers to the $(0,0)$-th entry of the above matrix. Moreover, $\alpha=1-\lambda_{0}a_{0},$ and
[…]
We can now state certain sufficient conditions for $F_{w}$ to become a bounded operator.
Theorem 3.3.
If
$\limsup_{n\rightarrow+\infty}\left\{\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert,\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert,\left\lvert c_{n}\frac{a_{0}a_{1}\cdots a_{n+2}}{b_{-1}b_{0}\cdots b_{n+1}}\right\rvert,\left\lvert d_{-n}\frac{b_{-1}b_{-2}\cdots b_{-n}}{a_{-1}a_{-2}\cdots a_{-n+1}}\right\rvert\right\}<1,$
and
$\sup_{n\geq 0}~{}\left\{\left\lvert w_{n}\frac{a_{n}}{a_{n+1}}\right\rvert,\left\lvert w_{-n}\frac{b_{-n-1}}{b_{-n}}\right\rvert\right\}<\infty,$
then $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A})$.
Proof.
It suffices to show that the above infinite matrix acts as a bounded operator on $\ell^{p}(\mathbb{Z})$. To this end, write the matrix of $F_{w}$ as a formal series of infinite matrices as follows:
$[F_{w}]=\cdots+[E_{-2}]+[E_{-1}]+[D]+[E_{1}]+[E_{2}]+\cdots,$
where $[D]$ is the matrix of the diagonal operator
$\text{diag}~{}(\cdots,\alpha d_{-3},\alpha d_{-2},-\alpha\frac{c_{-1}a_{0}a_{1}}{b_{-1}b_{0}},\bm{-\alpha\frac{c_{0}a_{1}a_{2}}{b_{0}b_{1}}},-\alpha\frac{c_{1}a_{2}a_{3}}{b_{1}b_{2}},-\alpha\frac{c_{2}a_{3}a_{4}}{b_{2}b_{3}},-\alpha\frac{c_{3}a_{4}a_{5}}{b_{3}b_{4}},\cdots)$
on $\ell^{p}(\mathbb{Z}).$ Also, the matrix $[E_{i}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $i$-th subdiagonal, where $i\geq 1$. Observe that $[E_{i}]$ is the matrix of suitable powers of a bilateral weighted forward shift. On the other hand, the matrix $[E_{-i}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $i$-th superdiagonal, where $i\geq 1$. Again observe that, $[E_{-i}]$ is the matrix of a suitable power of a bilateral weighted backward shift. Note that $[E_{1}]$ is the matrix of the standard weighted bilateral forward shift $E_{1}(e_{n})\mapsto\gamma_{n}e_{n+1}$ on $\ell^{p}(\mathbb{Z})$, $n\in\mathbb{Z}$, having weights
[…]
For each $n\geq 0,$ we have
$c_{n}=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}=\frac{b_{n}}{a_{n+1}}\frac{w_{n+1}a_{n+1}}{a_{n+2}}-\frac{w_{n}a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}},$
and for $n<0$
$d_{n}=w_{n}\frac{a_{n}}{b_{n}}-w_{n+1}\frac{a_{n+1}}{b_{n+1}}=\frac{a_{n}}{b_{n-1}}\frac{w_{n}b_{n-1}}{b_{n}}-\frac{a_{n+1}}{b_{n}}\frac{w_{n+1}b_{n}}{b_{n+1}}.$
By the assumptions, it follows that $\{c_{n}\}$ and $\{d_{n}\}$ are bounded sequences. Moreover, we formally have
$\lVert D\rVert\leq\lvert\alpha\rvert\sup_{n\geq 0}\left\{\left\lvert c_{n}\frac{a_{n+1}a_{n+2}}{b_{n}b_{n+1}}\right\rvert,|d_{-n-2}|\right\},$
$\lVert E_{1}\rVert\leq\sup_{n\in\mathbb{Z}}~{}~{}\lvert\gamma_{n}\rvert,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\lVert E_{2}\rVert\leq|1-\alpha|\sup_{n\geq 0}\left\{\left\lvert c_{n}\right\rvert,\left\lvert\frac{d_{-n-3}b_{-n-2}b_{-n-3}}{a_{-n-1}a_{-n-2}}\right\rvert\right\},$
$\lVert E_{i}\rVert\leq|1-\alpha|\sup_{n\geq 0}\left\{\left\lvert c_{n}\prod_{k=2}^{i-1}\frac{b_{n+k}}{a_{n+k+1}}\right\rvert,\left\lvert d_{-n-i-1}\prod_{k=2}^{i+1}\frac{b_{-n-k}}{a_{-n-k+1}}\right\rvert\right\},~{}~{}~{}~{}\forall i\geq 3,$
and
$\lVert E_{-i}\rVert\leq|\alpha|\sup_{n\geq 0}\left\{\left\lvert c_{n+i-1}\prod_{k=0}^{i+1}\frac{a_{n+k}}{b_{n+k-1}}\right\rvert,\left\lvert d_{-n-2}\prod_{k=2}^{i+1}\frac{a_{-n-k}}{b_{-n-k-1}}\right\rvert\right\},~{}~{}~{}~{}\forall i\geq 1.$
On the other hand, by the conditions in the statement of the theorem, namely
$\limsup_{n\rightarrow+\infty}\left\{\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert,\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert,\left\lvert c_{n}a_{0}\prod_{k=1}^{n+2}\frac{a_{k}}{b_{k-1}}\right\rvert,\left\lvert d_{-n}b_{-1}\prod_{k=1}^{n-1}\frac{b_{-k-1}}{a_{-k}}\right\rvert\right\}<1,$ there are $r,R,s,S<1$ and $n_{0}\in\mathbb{N}$ such that
$\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<r,\hskip 5.69046pt\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<R,\hskip 5.69046pt\left\lvert c_{n}a_{0}\prod_{k=1}^{n+2}\frac{a_{k}}{b_{k-1}}\right\rvert<s,\hskip 5.69046pt\text{and}\hskip 5.69046pt\left\lvert d_{-n}b_{-1}\prod_{k=1}^{n-1}\frac{b_{-k-1}}{a_{-k}}\right\rvert<S,$
for $n\geq n_{0},$ from which it follows that
$\lVert E_{i}\rVert\leq|1-\alpha|\max\left\{r^{i-1},SR^{n-n_{0}+2}\right\},~{}~{}\forall i\geq 3~{}~{}\text{and}~{}~{}\lVert E_{-i}\rVert\leq|\alpha|\max\left\{sr^{n-n_{0}},R^{i+1}\right\},~{}~{}\forall i\geq 1.$
Hence, the series $\sum_{i=1}^{\infty}\lVert E_{-i}\rVert$ and $\sum_{i=1}^{\infty}\lVert E_{i}\rVert$ are convergent. Consequently, the matrix $[F_{w}]$ acts as a bounded operator on $\ell^{p}(\mathbb{Z})$. This completes the proof of the theorem. ∎
A necessary condition for $F_{w}$ to be bounded on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$ is stated below.
Proposition 3.4.
If $F_{w}$ is a bounded operator on any of the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$, then $\{\gamma_{n}\}_{n=-\infty}^{\infty}$ is bounded, where $\gamma_{n}$ is the sequence as in (3.1).
Proof.
If $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A})$, then its matrix $[F_{w}]$ acts as a bounded operator on the space $\ell^{p}(\mathbb{Z})$. Hence,
$\sup_{n\in\mathbb{Z}}||[F_{w}](e_{n})||_{\ell^{p}(\mathbb{Z})}<\infty,$
where $e_{n}$ is the standard unit vector in $\ell^{p}(\mathbb{Z})$. On the other hand, $[F_{w}](e_{n})$ is the $n$-th column of the bilateral matrix $[F_{w}]$. With the observation that
$|\gamma_{n}|\leq\|[F_{w}](e_{n})\|_{\ell^{p}(\mathbb{Z})},\hskip 5.69046ptn\in\mathbb{Z},$
we immediately get $\sup_{n}|\gamma_{n}|<\infty$. ∎
In terms of $k_{n}$, a general necessary condition for $F_{w}$ to be bounded can be obtained, as follows:
Proposition 3.5.
Suppose $\mathcal{E}$ is a Banach space of analytic functions defined on an annulus $\mathtt{A}$, having continuous evaluation functionals. Suppose that $k_{n}\neq 0$ for all $n\in\mathbb{Z}$. If $F_{w}$ is a bounded operator on $\mathcal{E}$, then $\sup~{}\Big{\{}w_{n-1}\frac{\|k_{n-1}\|}{\|k_{n}\|}:~{}{n\in\mathbb{Z}}\Big{\}}<\infty.$
Proof.
Consider the vectors of unit norm, $u_{n}:=\frac{k_{n}}{\|k_{n}\|},~{}n\in\mathbb{Z}$. Since $F_{w}$ is a bounded operator, we get that $\{F_{w}^{*}(u_{n}):n\in\mathbb{Z}\}$ is a norm-bounded set in $\mathcal{E}^{*}$. This, along with the previous proposition, implies the result. ∎
The matrix representation of $F_{w}$ hints that the operator $F_{w}$ is far from being similar to a usual weighted shift on $\ell^{p}(\mathbb{Z})$ in general. Indeed, we can produce a weight $w$, some $a=\{a_{n}\}$ and $b=\{b_{n}\}$ such that the $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ is not similar to any forward weighted shift on $\ell^{p}(\mathbb{Z})$, see Section $5$. Below, we provide sufficient conditions so that the operator becomes a compact perturbation of a usual weighted shift. A consequence is the description of the essential spectrum $\sigma_{e}(F_{w})$ of the weighted shift $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$, for $1<p<\infty$. For a bounded operator $T$ on a complex Banach space $X$, the essential spectrum is denoted and defined by
$\sigma_{e}(T):=\{\lambda\in\mathbb{C}:$   $T-\lambda I$ is not Fredholm}}\}},
cf. Shields [36]. See, also, Bayart and Matheron [4], and Douglas [17].
Theorem 3.6.
Let $F_{w}$ be bounded on $\ell^{p}_{{a},{b}}(\mathtt{A})$. If
$\lim_{n\rightarrow\infty}\left\lvert w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right\rvert=0~{}\text{and}~{}\lim_{n\rightarrow-\infty}\left\lvert w_{n}\frac{a_{n}}{b_{n}}-w_{n+1}\frac{a_{n+1}}{b_{n+1}}\right\rvert=0,$
then the operator $F_{w}$ on $\ell^{p}_{{a},{b}}(\mathtt{A})$ is similar to $E_{\alpha}+K$ for some compact operator $K$ and a bilateral weighted forward shift $E_{\alpha}$ acting on the sequence space $\ell^{p}(\mathbb{Z})$, where the weight sequence $\alpha=(\alpha_{n})$ is given by
$\alpha_{n}=\begin{cases}w_{n}\frac{a_{n}}{a_{n+1}},\hskip 28.45274ptn\geq-1\\ w_{n+1}\frac{b_{n}}{b_{n+1}},\hskip 19.91684ptn\leq-2.\end{cases}$
Consequently, $\sigma_{e}(F_{w})=\sigma_{e}(E_{\alpha})$ which is an annulus.
Proof.
This result follows from the above matrix representation of $F_{w}$. Consider the similarity operator $V:\ell^{p}_{a,b}(\mathtt{A})\rightarrow\ell^{p}(\mathbb{Z})$ given by $V\big{(}\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}\big{)}=\sum_{n=-\infty}^{\infty}\lambda_{n}e_{n}$, where $\{e_{n}\}_{n\in\mathbb{Z}}$ is the standard basis in $\ell^{p}(\mathbb{Z})$. Now, from the proof of Theorem 3.3 we understand that $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ is similar via $V$ to the operator given by the series (which is convergent absolutely with respect to the operator norm)
$\sum_{i=1}^{\infty}E_{-i}+D+E_{1}+\sum_{i=2}^{\infty}E_{i}.$
Further, by our assumptions, the operators $E_{-i},$ $D$ and $E_{i}$ are compact on $\ell^{p}(\mathbb{Z})$ as the entries in each of $E_{-i},$ $D$ and $E_{i}$ converge to 00 except for the case $i=1$. Also, the operator $E_{1}$ is the weighted forward shift on $\ell^{p}(\mathbb{Z})$ with weights $\alpha=\{\alpha_{n}\}_{n=-\infty}^{\infty}$, as in the statement of this theorem. We now take
$E_{\alpha}:=E_{1}$, and $K:=\sum_{i=1}^{\infty}E_{-i}+D+\sum_{i=2}^{\infty}E_{i}$,
required as in the theorem since $K$ is a compact operator on $\ell^{p}(\mathbb{Z})$.
Since the essential spectrum is invariant under compact perturbations, we have $\sigma_{e}(F_{w})=\sigma_{e}(E_{\alpha}+K)=\sigma_{e}(E_{\alpha})$. ∎
3.2. Linear dynamics of $F_{w}$ and its adjoint $F_{w}^{*}$
In this subsection, we obtain necessary and sufficient conditions for $F_{w}$ and its adjoint operator $F_{w}^{*}$ to be hypercyclic, mixing, supercyclic, or chaotic on $\ell^{p}_{a,b}(\mathtt{A})$, $c_{0,a,b}(\mathtt{A})$ and their dual spaces, respectively.
Theorem 3.7.
The following hold for the bilateral weighted forward shift $F_{w}$ acting on $\ell^{p}_{a,b}(\mathtt{A})$ and $c_{0,a,b}(\mathtt{A})$, assuming that the conditions of Theorem 3.3 are satisfied:
  1. (i)
    $F_{w}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})$ if
    $\limsup_{n\rightarrow\infty}\left\lvert\prod_{k=0}^{n-1}w_{\nu+k}^{-1}\prod_{k=0}^{\nu+n-1}\frac{b_{k}}{a_{k}}\right\rvert=\infty=\limsup_{n\rightarrow\infty}\left\lvert\prod_{k=1}^{n}w_{\nu-k}\prod_{k=1}^{-\nu+n}\frac{a_{-k}}{b_{-k}}\right\rvert,$ (3.2)
    ${}~{}\forall~{}~{}\nu\in\mathbb{N}.$
  2. (ii)
    $F_{w}$ is topologically mixing on $\ell^{p}_{a,b}(\mathtt{A})$ if
    $\lim_{n\rightarrow\infty}\left\lvert\prod_{k=0}^{n-1}w_{\nu+k}^{-1}\prod_{k=0}^{\nu+n-1}\frac{b_{k}}{a_{k}}\right\rvert=\infty=\lim_{n\rightarrow\infty}\left\lvert\prod_{k=1}^{n}w_{\nu-k}\prod_{k=1}^{-\nu+n}\frac{a_{-k}}{b_{-k}}\right\rvert$ (3.3)
    ${}~{}\forall~{}~{}\nu\in\mathbb{N}.$
  3. (iii)
    $F_{w}$ is supercyclic on $\ell^{p}_{a,b}(\mathtt{A})$ if
    $\limsup_{n\rightarrow\infty}\left\lvert\frac{\prod_{k=1}^{n}w_{\nu-k}}{\prod_{k=0}^{n-1}w_{\nu+k}}\prod_{k=0}^{\nu+n-1}\frac{b_{k}}{a_{k}}~{}\prod_{k=1}^{-\nu+n}\frac{a_{-k}}{b_{-k}}\right\rvert=\infty,\hskip 14.22636pt\forall~{}~{}\nu\in\mathbb{N}.$ (3.4)
Proof.
(i) Suppose that (3.2) holds, and we will use Gethner-Shapiro criterion to show that $F_{w}$ is hypercyclic. Let $D$ be the span of $\{z^{\nu}:\nu\in\mathbb{Z}\}$. Then, $D$ is dense in $\ell^{p}_{a,b}(\mathtt{A}),$ as it contains the normalized Schauder basis vectors $(a_{n}+b_{n}z)z^{n}$. Define $S:D\rightarrow D$ by
$S(z^{\nu})=\frac{1}{w_{\nu-1}}z^{\nu-1},~{}~{}\forall~{}~{}\nu\in\mathbb{Z}.$
Since $F_{w}S(f)=f,$ $\forall$ $f\in D,$ we just need to show that both the sequences $F_{w}^{n_{k}}(z^{\nu})$ and $S^{n_{k}}(z^{\nu})$ tend to 0 as $k\to\infty,$ for some increasing sequence $(n_{k})\subset\mathbb{N},$ and for all $\nu\in\mathbb{N}.$ Note that
$F^{n_{k}}_{w}(z^{\nu})=w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}z^{\nu+n_{k}},$
and
$S^{n_{k}}(z^{\nu})=\frac{z^{\nu-n_{k}}}{w_{\nu-1}\cdots w_{\nu-n_{k}}}.$
Combining these expressions with Proposition 2.4, and the assumption (3.2), we see that $F_{w}$ satisfies the Gethner-shapiro criterion. Hence, $F_{w}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A}).$
(ii) To conclude the mixing property of $F_{w}$, proceed as in the proof of (i) by taking $n_{k}=k$ for all $k\geq 1$, in the Genthner-Shapiro criterion.
(iii) We now prove the supercyclicity. Suppose that (3.4) holds. In the supercyclicity criterion, let $D$ and $S$ be defined as above. It suffices to verify that $\lVert F^{n_{k}}_{w}(z^{\nu})\rVert\lVert S^{n_{k}}(z^{\nu})\rVert$ tends to 0 as $k\to\infty,$ for some increasing sequence $(n_{k})\subset\mathbb{N},$ $\nu\in\mathbb{N}.$ Since we have
$\lVert F^{n_{k}}_{w}(z^{\nu})\rVert_{\ell^{p}_{a,b}(\mathtt{A})}\lVert S^{n_{k}}(z^{\nu})\rVert_{\ell^{p}_{a,b}(\mathtt{A})}=\lVert w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}z^{\nu+n_{k}}\rVert\left\lVert\frac{z^{\nu-n_{k}}}{w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}}\right\rVert,$
Proposition 2.4, and the assumption (3.4), yield that $F_{w}$ satisfies the Supercyclicity criterion. ∎
Remark 3.8.
It should be noted that the unweighted forward shift $F$ cannot be supercyclic on a Banach space $\mathcal{E}$ of analytic functions over an annulus $r<|z|<R$, having bounded evaluation functionals. Indeed, the boundedness of $F$ on $\mathcal{E}$ is equivalent to the following: $f\in\mathcal{E}\Rightarrow zf\in\mathcal{E}.$ Thus, $F$ is the multiplication operator $M_{\varphi}(f)=\varphi f$ defined on $\mathcal{E}$, where $f\in\mathcal{E}$, and $\varphi(z)=z$. However, a multiplication operator cannot be supercyclic as its adjoint has more than two distinct eigenvalues with eigenvectors being the non-zero evaluation functionals, cf. [19].
The next result is on the chaos of $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ which involves lengthy computations.
Theorem 3.9.
Assume that $a=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b=\{b_{n}\}_{n=-\infty}^{\infty}$ satisfy the conditions of Theorem 3.3. Consider the following statements for the bilateral weighted forward shift $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ :
  1. (i)
    $F_{w}$ is chaotic on $\ell^{p}_{a,b}(\mathtt{A})$.
  2. (ii)
    $F_{w}$ has a non-trivial periodic vector.
  3. (iii)
    $\sum_{n=1}^{\infty}\left\lvert\frac{1}{a_{n}}\prod_{k=0}^{n-1}w_{k}\right\rvert^{p}<\infty~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n=1}^{\infty}\left\lvert\frac{1}{b_{-n-1}}\frac{1}{\prod_{k=1}^{n}w_{-k}}\right\rvert^{p}<\infty.$
  1. (iv)
    $\sum_{n=1}^{\infty}\left\lvert\prod_{k=0}^{n-1}\frac{a_{k}w_{k}}{b_{k}}\right\rvert^{p}<\infty~{}~{}\text{and}~{}~{}\displaystyle\sum_{n=1}^{\infty}\left\lvert\prod_{k=1}^{n}\frac{b_{-k}}{w_{-k}a_{-k}}\right\rvert^{p}<\infty.$
Then $(i)\Rightarrow(ii)\Rightarrow(iii)$ and $(iv)\Rightarrow(iii)$ Also, the condition $(iv)\Rightarrow(i).$
Proof.
It is trivial that $(i)\Rightarrow(ii)$.
To prove $(ii)\Rightarrow(iii)$, let $f(z)=\sum_{n=-\infty}^{\infty}\lambda_{n}f_{n}(z)$ be a non-zero periodic vector for $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A}).$ Rearranging into Laurent series, we have $f(z)=\sum_{n=-\infty}^{\infty}A_{n}z^{n}$, where $A_{n}=\lambda_{n-1}b_{n-1}+\lambda_{n}a_{n}$. Let us choose $m\in\mathbb{N}$ such that $F_{w}^{m}f(z)=f(z),$ for all $z$ in some annulus $\mathtt{A}$. Since $F_{w}^{km}f(z)=f(z)$ for all $k\geq 1$, it follows that
$\sum_{n=-\infty}^{\infty}w_{n}w_{n+1}\cdots w_{n+km-1}A_{n}z^{n+km}=\sum_{n=-\infty}^{\infty}A_{n}z^{n},$
for all $z\in\mathtt{A}$. Using the coefficients of like powers for comparison, we obtain $A_{j+km}=w_{j}w_{j+1}\cdots w_{j+km-1}A_{j}$, $\forall~{}~{}k\geq 1$, where $10\leq j\leq m-$ 1. For $j=0,$ we have $\lambda_{km}a_{km}+\lambda_{km-1}b_{km-1}=w_{0}w_{1}\cdots w_{km-1}(\lambda_{0}a_{0}+\lambda_{-1}b_{-1})$, for $k\geq 1.$ Writing $s_{n,j}:=w_{j}w_{j+1}\cdots w_{j+n-1}$ for convenience, we see that
$\lvert\lambda_{0}a_{0}+\lambda_{-1}b_{-1}\rvert^{p}\sum_{k=1}^{\infty}\left\lvert\frac{s_{km,0}}{a_{km}}\right\rvert^{p}\leq C_{1}r^{p}\big{(}\sum_{k=1}^{\infty}\left\lvert\lambda_{km-1}\right\rvert^{p}+\sum_{k=1}^{\infty}\left\lvert\lambda_{km}\right\rvert^{p}\big{)},$
for some constant $C_{1}>0$. Here, $r:=\sup_{n\geq 0}\left\lvert b_{n}/a_{n+1}\right\rvert$. As $\{\lambda_{n}\}\in\ell^{p}(\mathbb{Z})$, we have $\sum_{k=1}^{\infty}\left\lvert\frac{s_{km,0}}{a_{km}}\right\rvert^{p}<\infty.$ Along the same lines, for $j=1,2,\ldots,m-1,$ we have $\lambda_{j+km}a_{j+km}+\lambda_{j+km-1}b_{j+km-1}=w_{j}w_{j+1}\cdots w_{j+km-1}(\lambda_{j}a_{j}+\lambda_{j-1}b_{j-1})$, and $\sum_{k=1}^{\infty}\left\lvert\frac{s_{km,j}}{a_{j+km}}\right\rvert^{p}<\infty.$ Consequently, we get the convergence of $\sum_{k=1}^{\infty}\left\lvert\frac{s_{k,0}}{a_{k}}\right\rvert^{p}$. In a similar way, it can be proved that $\sum_{k=1}^{\infty}\left\lvert w_{-1}w_{-2}\cdots w_{-k}b_{-k-1}\right\rvert^{-p}$ is convergent, and thus the implication $(ii)\Rightarrow(iii)$ follows.
The implications in $(iv)\Rightarrow(iii),$ are easily seen to hold, just by noting the expressions
$\frac{1}{a_{n}}=\frac{a_{0}\cdots a_{n-1}}{b_{0}\cdots b_{n-1}}\frac{b_{0}\cdots b_{n-1}}{a_{0}\cdots a_{n}}~{}\text{and}~{}\frac{1}{b_{-n-1}}=\frac{b_{-1}\cdots b_{-n}}{a_{-1}\cdots a_{-n}}\frac{a_{-1}\cdots a_{-n}}{b_{-1}\cdots b_{-n-1}},$
along with $\limsup_{n}|b_{n}/a_{n+1}|<1$, and $\limsup_{n}|a_{-n+1}/b_{-n}|<1$.
We now prove that $(iv)$ implies $(i)$ by applying the chaoticity criterion. Take $D$ and $S$ to be the same set and map as in the previous theorem. It suffices if we show that
$\sum_{n=1}^{\infty}w_{0}w_{1}\cdots w_{n-1}z^{n}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n=1}^{\infty}\frac{1}{w_{-1}w_{-2}\cdots w_{-n}}z^{-n}$
are unconditionally convergent in $\ell^{p}_{a,b}(\mathtt{A}).$ Recalling the expansion of $z^{n}$ in terms of the basis $\{f_{j}\}$, and to keep calculations simple, let us assume that $s_{n}=w_{0}w_{1}\cdots w_{n}$ and $u_{n}=\frac{a_{0}\cdots a_{n}}{b_{0}\cdots b_{n}}$ for all $n\geq 0$ and we obtain
$\sum_{n=1}^{\infty}s_{n-1}z^{n}$ $=$ $\frac{1-\lambda_{0}a_{0}}{b_{-1}}\left(\sum_{j=0}^{\infty}(-1)^{j+1}s_{j}u_{j}\right)\times\left(f_{-1}+\sum_{j=2}^{\infty}(-1)^{j-1}\frac{a_{-1}\cdots a_{-j+1}}{b_{-2}\cdots b_{-j}}f_{-j}\right)$
$+\frac{\lambda_{0}a_{0}-1}{a_{0}}\left(\sum_{j=0}^{\infty}(-1)^{j+1}s_{j}u_{j}\right)f_{0}+\frac{\lambda_{0}b_{0}}{a_{1}}\left(s_{0}u_{0}+\frac{\lambda_{0}a_{0}-1}{a_{0}}\sum_{j=1}^{\infty}(-1)^{j}s_{j}u_{j}\right)f_{1}$
$+\sum_{j=2}^{\infty}\left(\frac{\lambda_{0}a_{0}}{b_{j}u_{j}}\left(\sum_{k=1}^{j}(-1)^{k-1}s_{j-k}u_{j-k}+\frac{\lambda_{0}a_{0}-1}{a_{0}}\sum_{k=0}^{\infty}(-1)^{k+1}s_{j+k}u_{j+k}\right)\right)f_{j}.$
Now, using the conditions in (iv) we can show that the coefficients in the previous expansion in terms of $\{f_{n}\}$ form a member in $\ell^{p}(\mathbb{Z})$, and hence, it converges unconditionally as $\{f_{n}\}$ is an unconditional basis. Thus, series $\sum_{n=1}^{\infty}w_{0}\cdots w_{n-1}z^{n}$ is convergent. It remains to show that $\sum_{n=1}^{\infty}\frac{1}{w_{-1}w_{-2}\cdots w_{-n}}z^{-n}$ is unconditionally convergent, for that we will write the expansion of $z^{-n}$ in terms of the basis $\{f_{j}\}$ and then proceed as above. We can now conclude that $F_{w}$ satisfies the chaoticity criterion, and (i) follows. ∎
We now turn to study the hypercyclicity, supercyclicity, and chaos for the adjoint $F_{w}^{*}$ using Proposition 3.1. In the following theorem, we derive sufficient conditions for the same.
Theorem 3.10.
For $1<p<\infty$, assume that $a=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b=\{b_{n}\}_{n=-\infty}^{\infty}$ satisfy the conditions of Theorem 3.3. Then, the following hold for adjoint weighted forward shift operator $F_{w}^{*}$:
  1. (i)
    $F_{w}^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$ if
    $\liminf_{n\rightarrow\infty}~{}\left\lvert b_{\nu-n-1}\prod_{k=1}^{n}w_{\nu-k}\right\rvert=0=\liminf_{n\rightarrow\infty}~{}\left\lvert\frac{a_{\nu+n}}{\prod_{k=0}^{n-1}w_{\nu+k}}\right\rvert,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.5)
  2. (ii)
    $F_{w}^{*}$ is topologically mixing on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$ if
    $\lim_{n\rightarrow\infty}\left\lvert b_{\nu-n-1}\prod_{k=1}^{n}w_{\nu-k}\right\rvert=0=\lim_{n\rightarrow\infty}\left\lvert~{}\frac{a_{\nu+n}}{\prod_{k=0}^{n-1}w_{\nu+k}}\right\rvert,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.6)
  3. (iii)
    $F_{w}^{*}$ is supercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$ if
    $\liminf_{n\rightarrow\infty}\left|b_{\nu-n-1}a_{\nu+n}\frac{\prod_{k=1}^{n}w_{\nu-k}}{\prod_{k=0}^{n-1}w_{\nu+k}}\right|=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.7)
Proof.
Suppose that (3.5) holds, and we will use Gethner-Shapiro criterion to show that $F_{w}^{*}$ is hypercyclic by taking $D=\text{span}\{k_{\nu}:\nu\in\mathbb{Z}\}$, and $S$ as the map defined by
$S(k_{\nu})=\frac{1}{w_{\nu}}k_{\nu+1},~{}~{}\forall~{}~{}\nu\in\mathbb{Z}.$
The map $S$ is well-defined as $\{k_{\nu}:\nu\in\mathbb{Z}\}$ is linearly independent. Indeed, if $\sum_{\nu=m_{1}}^{m_{2}}\alpha_{\nu}k_{\nu}=0$ for some $m_{1}<m_{2}$, then $\alpha_{\nu}=0$ because of the following: $k_{\nu}(z^{m})=0$ if and only if $\nu\neq m$ and $k_{\nu}(z^{\nu})\neq 0$.
By Proposition 2.3, the set $D$ is dense in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Since $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ we just need to show that, for some increasing $(n_{k})\subset\mathbb{N},$ the sequences $(F_{w}^{*})^{n_{k}}(k_{\nu})$ and $S^{n_{k}}(k_{\nu})$ both tend to 0 as $k\to\infty,$ $\nu\in\mathbb{N}.$ But this is clear by Proposition (2.2), $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<1,$ $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$
$\lVert(F_{w}^{*})^{n_{k}}(k_{\nu})\rVert=\lVert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}k_{\nu-n_{k}}\rVert\leq M_{1}\lvert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}b_{\nu-n_{k}-1}\rvert$
and
$\lVert S^{n_{k}}(k_{\nu})\rVert=\left\lVert\frac{k_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rVert\leq M_{2}\left\lvert\frac{a_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rvert,$
for some constants $M_{1},M_{2}>0$. Now by the assumption (3.5), we get that $F_{w}^{*}$ satisfies the Gethner-shapiro criterion. Hence, $F_{w}^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$
(ii) The proof here is verbatim. Taking $n_{k}=k$ for all $k\geq 1$, and pursue exactly as in the proof of (i).
(iii) Suppose that (3.7) holds. In the supercyclicity criterion, take the set $D$ and the map defined as above. It suffices to prove that $\lVert(F_{w}^{*})^{n_{k}}(k_{\nu})\rVert\lVert S^{n_{k}}(k_{\nu})\rVert$ tends to zero as $k\to\infty,$ for any increasing sequence $(n_{k})\subset\mathbb{N},$ $\nu\in\mathbb{N}.$ But this is clear since by Proposition (2.2), $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{a_{-n+1}}{b_{-n}}\right\rvert<1,$ and $\lim\sup_{n\rightarrow+\infty}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$
$\lVert(F_{w}^{*})^{n_{k}}(k_{\nu})\rVert\lVert S^{n_{k}}(k_{\nu})\rVert$ $=$ $\lVert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}k_{\nu-n_{k}}\rVert\left\lVert\frac{k_{\nu+n_{k}}}{w_{\nu}w_{t+1}\cdots w_{t+n_{k}-1}}\right\rVert$
$\leq$ $M\lvert w_{\nu-1}w_{\nu-2}\cdots w_{\nu-n_{k}}b_{\nu-n_{k}-1}\rvert\left\lvert\frac{a_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rvert,$
for some constant $M>0$ and $n_{k}>\nu.$ Now with the assumption (3.7), we get that $F_{w}^{*}$ satisfies the Supercyclicity criterion. Hence, $F_{w}^{*}$ is supercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$
Our next aim is to obtain necessary and sufficient conditions for $F_{w}^{*}$ to be chaotic. We need a representation of the functional $f_{n}^{*}$ in terms of $k_{n}$.
Proposition 3.11.
Let $k_{n}$ denote the co-ordinate functional $f\mapsto\widehat{f}(n),$ defined on $\ell^{p}_{a,b}(\mathtt{A})$ or $c_{0,a,b}(\mathtt{A})$, where $n\in\mathbb{Z}$. The following hold.
  • (i)
    For $n\geq 1,$
    $f_{n}^{*}=\frac{1}{a_{n}}k_{n}+\sum_{j=1}^{n}(-1)^{j}\frac{b_{n-1}b_{n-2}\cdots b_{n-j}}{a_{n}a_{n-1}\cdots a_{n-j}}k_{n-j}+(-1)^{n+1}\frac{b_{n-1}\cdots b_{-1}}{a_{n}\cdots a_{0}}f_{-1}^{*}.$
  • (ii)
    For $n\geq 2,$
    $f_{-n}^{*}=\frac{1}{b_{-n}}k_{-n+1}+\sum_{j=2}^{n}(-1)^{j-1}\frac{a_{-n+1}a_{-n+2}\cdots a_{-n+j-1}}{b_{-n}b_{-n+1}\cdots b_{-n+j-1}}k_{-n+j}+(-1)^{n}\frac{a_{-n+1}\cdots a_{0}}{b_{-n}\cdots b_{-1}}f_{0}^{*}.$
Proof.
$(i)$ If $f\in\ell^{p}_{a,b}(\mathtt{A}),$ then $f(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}f_{k}(z)=\sum_{k=-\infty}^{\infty}\lambda_{k}(a_{k}+b_{k}z)z^{k}$, and $\|f\|=\left(\sum_{k=-\infty}^{\infty}\lvert\lambda_{k}\rvert^{p}\right)^{1/p}$. Rearranging into a Laurent series yields $f(z)=\sum_{k=-\infty}^{\infty}(\lambda_{k}a_{k}+\lambda_{k-1}b_{k-1})z^{k},$ and so, $k_{n}(f)=\lambda_{n}a_{n}+\lambda_{n-1}b_{n-1}.$ Now for $f\in\ell^{p}_{a,b}(\mathtt{A}),$ we have
$\frac{1}{a_{n}}k_{n}(f)-\cdots+(-1)^{n}\frac{b_{n-1}\cdots b_{0}}{a_{n}\cdots a_{0}}k_{0}(f)$ $=$ $\lambda_{n}+(-1)^{n}\frac{b_{n-1}\cdots b_{-1}}{a_{n}\cdots a_{0}}\lambda_{-1}$
$=$ $f_{n}^{*}(f)+(-1)^{n}\frac{b_{n-1}\cdots b_{-1}}{a_{n}\cdots a_{0}}f_{-1}^{*}(f),$
for all $n\geq 1$, which gives (i).
Proof of (ii) is similar. ∎
Theorem 3.12.
Let $\ell^{p}_{a,b}(\mathtt{A}),$ $1<p<\infty,$ be an analytic Banach space over the unit annulus corresponding to $a=\{a_{n}\}_{n=-\infty}^{\infty}$ and $b=\{b_{n}\}_{n=-\infty}^{\infty}$ satisfying the conditions of Theorem 3.3. Consider the following statements:
  1. (i)
    $F_{w}^{*}$ is chaotic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$.
  2. (ii)
    $F_{w}^{*}$ has a non-trivial periodic vector.
  3. (iii)
    $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{q}<\infty$ and $\sum_{n=1}^{\infty}\lvert w_{-1}w_{-2}\cdots w_{-n}b_{-n-1}\rvert^{q}<\infty.$
Then $(iii)\Rightarrow(i)\Rightarrow(ii).$ In addition, if for $\frac{1}{p}+\frac{1}{q}=1,$
$\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}b_{n+k-1}}{\prod_{k=1}^{j}a_{n+k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty~{}~{}~{}~{}~{}~{}~{}~{}\text{and}~{}~{}~{}~{}~{}~{}~{}~{}\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}a_{-n-k+1}}{\prod_{k=1}^{j}b_{-n-k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty,$
then $(ii)\Rightarrow(iii)$.
Proof.
It is trivial that (i) implies (ii).
To see that (ii) implies (iii), let $\psi=\sum_{m=-\infty}^{\infty}\lambda_{m}f_{m}^{*}$ be a non-zero periodic vector for $F_{w}^{*}$ on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$, where $f_{m}^{*}$, $m\in\mathbb{Z}$, forms an unconditional Schauder basis for $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$. Now by using Proposition 3.11, we get
$\psi=\sum_{m=-\infty}^{\infty}\lambda_{m}f_{m}^{*}:=\sum_{m=-\infty}^{\infty}A_{m}k_{m},$
where
$A_{0}=\displaystyle\sum_{j=2}^{\infty}(-1)^{j-1}\lambda_{-j}\frac{a_{-1}a_{-2}\cdots a_{-j+1}}{b_{-1}b_{-2}\cdots b_{-j}}+\displaystyle\sum_{j=1}^{\infty}(-1)^{j}\lambda_{j}\frac{b_{0}b_{1}\cdots b_{j-1}}{a_{0}a_{1}\cdots a_{j}},$
$A_{m}=\lambda_{m}\frac{1}{a_{m}}+\displaystyle\sum_{j=1}^{\infty}(-1)^{j}\lambda_{m+j}\frac{b_{m}b_{m+1}\cdots b_{m+j-1}}{a_{m}a_{m+1}\cdots a_{m+j}},$
and
$A_{-m}=\lambda_{-m-1}\frac{1}{b_{-m-1}}+\displaystyle\sum_{j=1}^{\infty}(-1)^{j}\lambda_{-m-1-j}\frac{a_{-m-1}a_{-m-2}\cdots a_{-m-j}}{b_{-m-1}b_{-m-2}\cdots b_{-m-1-j}},$
for all $m\in\mathbb{N}.$ Now, let $\nu\in\mathbb{N}$ such that $(F_{w}^{*})^{\nu}\psi=\psi$ for $z\in\mathtt{A}$. Then,
$(F_{w}^{*})^{n\nu}\psi=\psi$ for all $n\geq 1$.
Equating the coefficient of $k_{0}$ both sides, we get
$\frac{w_{n\nu-1}\cdots w_{0}}{a_{n\nu}}\left(\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right)=\cdots-\lambda_{-2}\frac{a_{-1}}{b_{-2}b_{-1}}-\lambda_{1}\frac{b_{0}}{a_{0}a_{1}}+\lambda_{2}\frac{b_{0}b_{1}}{a_{0}a_{1}a_{2}}-\cdots,$
Since the series in the right hand side converges, we get that
$C\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{q}=\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right\rvert^{q},$
for some $C>0$. Note that using Hölder inequality, we can get that
$\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right\rvert^{q}<\infty,$
as $\{\lambda_{n}\}_{n}\in\ell^{q}(\mathbb{Z})$ and $\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}b_{n+k-1}}{\prod_{k=1}^{j}a_{n+k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty.$ Therefore,
$\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{q}<\infty.$ Similarly, equating the coefficient of $k_{j},j=1,2,\cdots,\nu-1$ both sides we get $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu+j}}{w_{n\nu+j-1}w_{n\nu+j-2}\cdots w_{j}}\right\rvert^{q}<\infty.$ Consequently, the series $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{q}$ is convergent. It now remains to show that $\sum_{n=0}^{\infty}\lvert w_{-1}w_{-2}\cdots w_{-n}b_{-n-1}\rvert^{q}$ is convergent. To see that, we just need to equate the coefficient of $k_{-j-n\nu},j=0,1,\cdots,\nu-1$ both sides, and use the assumption $\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}a_{-n-k+1}}{\prod_{k=1}^{j}b_{-n-k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty$.
To get the implication in (iii) $\Rightarrow$ (i), we use the chaoticity criterion. As in the previous theorem, consider $D:=\text{span}\{k_{n}:n\in\mathbb{Z}\}$, and $S:D\rightarrow D$ which is given by $S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\in\mathbb{Z}.$ Observing that $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ we need only to prove that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ and $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ are unconditionally convergent in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ To this end, since $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\in\mathbb{Z}$, we have
$\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ $=$ $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}(a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*})$
$=$ $\frac{a_{1}}{w_{0}}\frac{b_{0}}{a_{1}}f_{0}^{*}+\sum_{n=1}^{\infty}\left(\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}+\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\frac{b_{n}}{a_{n+1}}\right)f_{n}^{*}.$
As $\limsup_{n}|b_{n}/a_{n+1}|<1$, one gets $N_{1}\in\mathbb{N}$ and $r<1$ such that $|b_{n}/a_{n+1}|<r$ for all $n\geq N_{1}$. Hence,
$\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}+\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\frac{b_{n}}{a_{n+1}}\right\rvert\leq\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert+r\left\lvert\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\right\rvert,$
for all $n\geq N_{1}$. By our hypothesis, the right hand side of the above inequality is $q$-absolutely summable, and so, the series $\sum_{n\geq 1}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is convergent. The unconditional convergence follows because $\{f_{n}^{*}\}$ is an unconditional basis. Similarly, to show that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ is unconditionally convergent, we write
$\sum_{n=1}^{\infty}w_{-1}\cdots w_{-n}k_{-n}$ $=$ $\sum_{n=1}^{\infty}w_{-1}\cdots w_{-n}(a_{-n}f_{-n}^{*}+b_{-n-1}f_{-n-1}^{*})$
$=$ $w_{-1}b_{-2}\frac{a_{-1}}{b_{-2}}f_{-1}^{*}+\sum_{n=2}^{\infty}\left(w_{-1}\cdots w_{-n+1}b_{-n}+w_{-1}\cdots w_{-n}\frac{b_{-n-1}a_{-n}}{b_{-n-1}}\right)f_{-n}^{*}.$
As $\limsup_{n}|a_{-n}/b_{-n-1}|<1$, we find $N_{2}\in\mathbb{N}$ and $R<1$ such that $|a_{-n}/b_{-n-1}|<R$ for all $n\geq N_{2}$. Hence,
$\left\lvert w_{-1}\cdots w_{-n+1}b_{-n}+w_{-1}\cdots w_{-n}b_{-n-1}\frac{a_{-n}}{b_{-n-1}}\right\rvert\leq\left\lvert w_{-1}\cdots w_{-n+1}b_{-n}\right\rvert+R\left\lvert w_{-1}\cdots w_{-n}b_{-n-1}\frac{a_{-n}}{b_{-n-1}}\right\rvert$
for all $n\geq N_{2}$. We can conclude that series $\sum_{n\geq 1}\frac{1}{w_{-1}w_{-2}\cdots w_{-n}}k_{-n}$ is unconditionally convergent. Hence, $F_{w}^{*}$ satisfies the chaoticity criterion, and (i) follows. ∎
The proofs of the main theorems on the hypercyclicity, supercyclicity, and chaos of $F_{w}^{*}$ on the dual of $\ell^{p}_{a,b}(\mathtt{A})$ show that we can have general results, stated below.
Theorem 3.13.
Let $\mathcal{E}$ be a Banach space of analytic functions on an annulus $\mathtt{A}$ concentric at 00, having bounded evaluation functionals. Let $\mathcal{E}^{*}$ be norm-separable, and let $\{k_{n}\}$ be linearly independent. If $F_{w}$ is bounded on $\mathcal{E}$, then the following hold.
  1. (i)
    $F_{w}^{*}$ is hypercyclic on $\mathcal{E}^{*}$ if
    $\liminf_{n\rightarrow\infty}~{}\Big{(}\big{\lvert}\prod_{j=1}^{n}w_{\nu-j}\big{\rvert}\|k_{\nu-n}\|\Big{)}=0=\liminf_{n\rightarrow\infty}~{}\Big{(}\frac{\|k_{\nu+n}\|}{\Big{\lvert}\prod_{j=0}^{n-1}w_{\nu+j}\Big{\rvert}}\Big{)},~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.8)
  2. (ii)
    $F_{w}^{*}$ is mixing if
    $\lim_{n\rightarrow\infty}~{}\Big{(}\big{\lvert}\prod_{j=1}^{n}w_{\nu-j}\big{\rvert}\|k_{\nu-n}\|\Big{)}=0=\lim_{n\rightarrow\infty}~{}\Big{(}\frac{\|k_{\nu+n}\|}{\left\lvert\prod_{j=0}^{n-1}w_{\nu+j}\right\rvert}\Big{)},~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (3.9)
  3. (iii)
    $F_{w}^{*}$ is supercyclic if
    $\liminf_{n\rightarrow\infty}\Big{(}\left|\frac{\prod_{j=1}^{n}w_{\nu-j}}{\prod_{j=0}^{n-1}w_{\nu+j}}\right|\|k_{\nu-n}\|\|k_{\nu+n}\|\Big{)}=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$
  4. (iv)
    Denote by $\widehat{f}(n)$, the $n$-th Laurent coefficient of $f\in\mathcal{E}$. If $\mathcal{E}$ is reflexive, then $F_{w}^{*}$ is chaotic provided
    $\sum_{n\geq 1}\Big{|}\Big{(}\prod_{j=1}^{n}w_{-j}\Big{)}\widehat{f}(-n)\Big{|}<\infty\hskip 11.38092pt\text{and}\hskip 11.38092pt\sum_{n\geq 1}\Big{|}\Big{(}\frac{1}{\prod_{j=1}^{n}w_{j}}\Big{)}\widehat{f}(n+1)\Big{|}<\infty,~{}~{}\forall~{}f\in\mathcal{E}.$
Proof.
Only an outline of the proofs are given. For all the implications, let $D$ and $S$ be as in the previous theorem. By a standard procedure, the results in (i), (ii), and (iii) follow from the hypercyclicity and supercyclicity criteria.
To see that (iv) also holds, via the chaoticity criterion we show that $\sum_{n=1}^{\infty}w_{-1}w_{-2}\cdots w_{-n}k_{-n}$ and $\sum_{n=1}^{\infty}\frac{1}{w_{1}w_{1}\cdots w_{n}}k_{n+1}$ are norm unconditionally convergent in $\mathcal{E}^{*}$. By the well known Orlicz-Pettis theorem (cf. [16]) it suffices to prove that these series are weakly unconditionally convergent. This is exactly the summability assumption in (iv). ∎
Next, we provide some necessary conditions for hypercyclicity, supercyclicity, and mixing of $F_{w}$ and its adjoint. We require the following basic result about the orbits under adjoints of hypercyclic and mixing operators.
Proposition 3.14.
(Bonet [6]) Let $T$ be a bounded operator on a Banach space $X$, and $x^{*}\neq 0$ be a bounded functional. If $T$ is hypercyclic, then $\sup_{n}\|T^{*n}x^{*}\|=\infty$.
If $T$ is mixing, then $\lim_{n\rightarrow\infty}\|T^{*n}x^{*}\|=\infty$.
We have:
Proposition 3.15.
Suppose that $F_{w}$ is bounded on a Banach space $\mathcal{E}$ of analytic functions over an annulus, having bounded evaluation functionals. For each fixed $\nu\in\mathbb{Z}$, the following hold.
  1. (i)
    If $F_{w}$ is hypercyclic, then $\sup_{n\geq 1}~{}(|w_{\nu-1}\cdots w_{\nu-n}|\|k_{\nu-n}\|)=\infty$. Also, if $F_{w}$ is mixing, then $\lim_{n\rightarrow\infty}~{}(|w_{\nu-1}\cdots w_{\nu-n}|\|k_{\nu-n}\|)=\infty.$
  2. (ii)
    Assume that $\mathcal{E}$ contains all Laurent polynomials. If the adjoint $F_{w}^{*}$ is hypercyclic,
    then $\sup_{n\geq 1}(|w_{\nu}\cdots w_{\nu+n}|\|z^{n+\nu}\|)=\infty.$ If $F_{w}^{*}$ is a mixing operator,
    then $\lim_{n\rightarrow\infty}(|w_{\nu}\cdots w_{\nu+n}|\|z^{n+\nu}\|)=\infty$.
Proof.
If $F_{w}$ is hypercyclic, then by the aforementioned result, we have $\sup_{n\geq 0}\|F_{w}^{*n}(k_{\nu})\|=\infty$. In view of Proposition 3.1, we have $F_{w}^{*n}(k_{\nu})=w_{\nu-1}\cdots w_{\nu-n}k_{\nu-n}$. So, the first part in (i) follows. Similarly, the other part also holds.
If $F_{w}^{*}$ is hypercyclic, then the orbit $\{F_{w}^{n}(z^{\nu}):n\geq 0\}$ is unbounded. Noting that $F_{w}^{n}(z^{\nu})=w_{\nu}\cdots w_{\nu+n}z^{\nu+n}$, we get (ii). ∎
Next, we state a result on the existence of hypercyclic subspaces for $F_{w}$ and $F_{w}^{*}$. It is well known that an operator $T$ on a complex Banach space, satisfying the hypercyclicity criterion has a hypercyclic subspace if and only if
$\sigma_{e}(T)\cap S^{1}\neq\phi,$
$S^{1}$ being the unit circle in $\mathbb{C}$. Every hypercyclic weighted shift on $\ell^{p}(\mathbb{Z})$ has a hypercyclic subspace. This is not true for hypercyclic unilateral shifts, and a characterization along with other results can be found in [4], [27], [29], and [30].
Theorem 3.16.
Suppose $F_{w}$ satisfies the conditions of Theorem 3.6. If $F_{w}$ is hypercyclic, then it has hypercyclic subspaces. A similar conclusion holds for $F_{w}^{*}$.
3.3. Examples and Illustrations
We provide explicit examples to illustrate our results on the linear dynamics of $F_{w}$ and $F_{w}^{*}$. Recall that a Laurent series $\sum_{n=-\infty}^{\infty}c_{n}z^{n}$ converges if $r<\lvert z\rvert<R,$ where
[…]
Indeed, $|z|<R$ and $\frac{1}{|z|}<\frac{1}{r}$ give, respectively, the convergence of $\sum_{n\geq 0}c_{n}z^{n}$ and $\sum_{n\leq 0}c_{n}z^{n}$.
Example 3.17.
Let $a_{0}=b_{0}=w_{0}=1,$ and $a_{n}=1,$ $b_{n}=\frac{1}{n},$ $w_{n}=\frac{1}{(n+1)!},$ for all $n\geq 1$, along with $a_{n}=n2^{n-2},$ $b_{n}=n2^{n},$ $w_{n}=8,$ for all $n\leq-1.$ By Theorem 2.1, for $1<p<\infty,$
$\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}=2^{q}+\sum_{n=1}^{\infty}\left(\left(\frac{n}{2^{n}}\frac{1}{2^{2}}+\frac{n}{2^{n}}\right)\lvert z\rvert^{-n}\right)^{q}+\sum_{n=1}^{\infty}\left(\left(1+\frac{1}{n}\right)\lvert z\rvert^{n}\right)^{q},$
where $11/p+1/q=$ 1. A simple computation shows that the natural annulus $\mathtt{A}$ of $\ell^{p}_{a,b}(\mathtt{A})$ is $r<|z|<R$, where $r=\frac{1}{2}$ and $R=1.$ Therefore $\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}<\infty$ if $\frac{1}{2}<\lvert z\rvert<1.$ In view of Theorem 3.3, $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}),$ where $\mathtt{A}$ is the given by annulus $\frac{1}{2}<\lvert z\rvert<1.$ Additionally, $F_{w}$ satisfies the conditions of Theorems 3.7 and 3.9, and so $F_{w}$ is mixing and chaotic on $\ell^{p}_{a,b}(\mathtt{A}).$ Indeed,
$\lim_{n\rightarrow\infty}\left\lvert\prod_{k=0}^{n-1}w_{k}^{-1}\prod_{k=0}^{n-1}\frac{b_{k}}{a_{k}}\right\rvert=\lim_{n\rightarrow\infty}\frac{\prod_{k=1}^{n}k!}{(n-1)!}=\infty~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\lim_{n\rightarrow\infty}\left\lvert\prod_{k=1}^{n}w_{-k}\prod_{k=1}^{n}\frac{a_{-k}}{b_{-k}}\right\rvert=\lim_{n\to\infty}2^{n}=\infty.$
Example 3.18.
In the previous example, if we let $w_{n}=1,$ for all $n\leq-1,$ $w_{n}=\frac{1}{(n+1)!},$ for all $n\geq 0,$ and $\{a_{n}\},\{b_{n}\}$ being the same as mentioned above, then we see that by Theorem 3.7, $F_{w}$ is supercyclic. Now, by Theorem 3.15, we know that if $F_{w}$ is hypercyclic on $\ell^{p}_{a,b}(\mathtt{A}),$ then $\sup_{n\geq 1}~{}(|w_{-1}\cdots w_{-n}|\|k_{-n}\|)=\infty$, which suggests that $\sup_{n\geq 1}~{}(|w_{-1}\cdots w_{-n}b_{-n-1}|)=\infty$, by using Proposition 2.2 and $\lim\sup_{n\rightarrow+\infty}\big{|}\frac{a_{-n+1}}{b_{-n}}\big{|}<1.$ However, in this case $F_{w}$ is not hypercyclic on $\ell^{p}_{a,b}(\mathtt{A})$ as
$\sup_{n\geq 1}\left\lvert b_{-n-1}\prod_{k=1}^{n}w_{-k}\right\rvert=\sup_{n\geq 0}\frac{n+1}{2^{n+1}}<\infty.$
Next, we provide an example of a space of the form $\ell^{p}_{a,b}(\mathtt{A})$ such that $F_{w}^{*}$ is mixing and chaotic.
Example 3.19.
Let $a_{0}=b_{0}=w_{0}=1,$ and $a_{n}=\frac{n}{2^{n}},$ $b_{n}=\frac{n}{2^{n+2}},$ $w_{n}=8,$ for all $n\geq 1,$ and $a_{n}=\frac{1}{n},$ $b_{n}=1,$ $w_{n}=\frac{1}{(-n+2)!},$ for all $n\leq-1.$ By Theorem 2.1, we have for $1<p<\infty,$
$\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}=2^{q}+\sum_{n=1}^{\infty}\left(\left(\frac{1}{n}+1\right)\lvert z\rvert^{-n}\right)^{q}+\sum_{n=1}^{\infty}\left(\left(\frac{n}{2^{n}}+\frac{n}{2^{n+2}}\right)\lvert z\rvert^{n}\right)^{q},$
where $11/p+1/q=$ 1. In this case, $r=1$, $R=2$, and hence, the annulus $\mathtt{A}$ is $21<|z|<$ 2. Further, $\sum_{n=-\infty}^{\infty}\left((\lvert a_{n}\rvert+\lvert b_{n}\rvert)\lvert z\rvert^{n}\right)^{q}<\infty$ if $21<\lvert z\rvert<$ 2 . In addition, by the conditions in Theorem 3.3, the weighted shift $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}).$ In view of Theorems 3.10 and 3.12, the adjoint $F_{w}^{*}$ is mixing and chaotic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ Indeed,
$\lim_{n\rightarrow\infty}\left\lvert b_{-n-1}\prod_{k=1}^{n}w_{-k}\right\rvert=\lim_{n\rightarrow\infty}\frac{1}{\prod_{k=3}^{n+2}k!}=0~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\lim_{n\rightarrow\infty}\left\lvert~{}\frac{a_{n}}{\prod_{k=0}^{n-1}w_{k}}\right\rvert=\lim_{n\rightarrow\infty}\frac{n}{2^{4n-3}}=0.$
Remark 3.20.
It should be noted that the operators $F_{w}$ considered in this section satisfy
$\lim_{n\rightarrow\infty}\left\lvert w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right\rvert=0~{}\text{and}~{}\lim_{n\rightarrow-\infty}\left\lvert w_{n}\frac{a_{n}}{b_{n}}-w_{n+1}\frac{a_{n+1}}{b_{n+1}}\right\rvert=0.$
Therefore, by the Theorem 3.6, we see that $F_{w}$ on $\ell^{p}_{a,b}(\mathtt{A})$ is similar to a compact perturbation of a bilateral weighted forward shift on $\ell^{p}(\mathbb{Z}).$ Consequently, we have a class of chaotic operators that are compact perturbation of some classical weighted shifts.
4. Adjoints of unilateral weighted shifts
In this section, we study the dynamical properties of the adjoint of the unilateral forward shift $F_{w}(\sum_{n=0}^{\infty}\lambda_{n}z^{n})=\sum_{n=0}^{\infty}\lambda_{n}w_{n}z^{n+1}$ defined on $\ell^{p}_{a,b}(\mathbb{D})$, where $w=\{w_{n}\}_{n=0}^{\infty}$ is a sequence of non-zero complex numbers. The differences between $F_{w}$ acting on $\ell^{p}_{a,b}(\mathbb{D})$ and the classical weighted forward shifts on $\ell^{p}$ will be pointed out in Section $5$. First, we derive some sufficient conditions for the unilateral $F_{w}$ to bounded in terms of $\{a_{n}\}$, $\{b_{n}\}$ and the weight sequence $\{w_{n}\}$. For the unweighted forward shift $F$, some necessary and sufficient conditions for $F$ to be bounded on $\ell^{2}_{a,b}(\mathbb{D})$ were obtained by Adams and McGuire [2]. Weighted backward shifts $B_{w}$ on $\ell^{p}_{a,b}(\mathbb{D})$ were studied by the authors in [13]. For operator theoretic results in the $p=2$ case, see [14], and [15].
We compute the matrix representation of the operator $F_{w}$ acting on $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$ with respect to the ordered Schauder basis $\{f_{n}\}_{n\geq 0}$. For this, we express a monomial $z^{n}$ as an expansion in the basis $\{f_{n}\}$; see (4.1) below. Indeed, fix $n\geq 0,$ and write $z^{n}=\sum_{j=0}^{\infty}\alpha_{j}f_{j}$ for some $\alpha_{j}\in\mathbb{C}.$ Then $z^{n}=\alpha_{0}a_{0}+\sum_{j=1}^{\infty}(\alpha_{j-1}b_{j-1}+\alpha_{j}a_{j})z^{j}.$ Comparing the coefficients of the same powers, a simple computation gives that $\alpha_{0}=\alpha_{1}=\cdot\cdot\cdot=\alpha_{n-1}=0,$ $\alpha_{n}=\frac{1}{a_{n}}$, and $\alpha_{n+k}=\frac{(-1)^{k}}{a_{n}}\frac{b_{n}b_{n+1}\cdot\cdot\cdot b_{n+k-1}}{a_{n+1}a_{n+2}\cdot\cdot\cdot a_{n+k}}$, where $k\geq 1$. From these expressions, we obtain
$z^{n}=\frac{1}{a_{n}}\sum_{j=0}^{\infty}(-1)^{j}\left(\frac{\prod_{k=0}^{j-1}b_{n+k}}{\prod_{k=0}^{j-1}a_{n+k+1}}\right)f_{n+j},\,\,\,\,(n\geq 0),$ (4.1)
where the coefficient for $j=0$ should be understood as $1$. (See, also, [2].) Note that $z^{n}\in\ell^{p}_{a,b}(\mathbb{D})$ if and only if
$\|z^{n}\|^{p}:=\frac{1}{|a_{n}|^{p}}\sum_{j=0}^{\infty}\left|\frac{\prod_{k=0}^{j-1}b_{n+k}}{\prod_{k=0}^{j-1}a_{n+k+1}}\right|^{p}<\infty.$ (4.2)
Acting $F_{w}$ on the normalized Schauder basis $f_{n}(z)=(a_{n}+b_{n}z)z^{n},n\geq 0$, of $\ell^{p}_{a,b}(\mathbb{D})$, we obtain
$F_{w}(f_{n})(z)$ $=$ $w_{n}a_{n}z^{n+1}+w_{n+1}b_{n}z^{n+2}=\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+\left(\frac{w_{n+1}b_{n}}{a_{n+2}}-\frac{w_{n}a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\right)a_{n+2}z^{n+2}$
$\hskip 48.36958pt=\frac{w_{n}a_{n}}{a_{n+1}}f_{{n+1}}+c_{n}\sum_{j=0}^{\infty}(-1)^{j}\left(\frac{\prod_{k=0}^{j-1}b_{n+2+k}}{\prod_{k=0}^{j-1}a_{n+3+k}}\right)f_{n+2+j},$
where $c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}$.
Therefore, the matrix of $F_{w}$ with respect to the basis $\{f_{n}\}_{n\geq 0}$ is
$[F_{w}]:=\begin{bmatrix}0&0&0&0&\cdots\\ w_{0}\frac{a_{0}}{a_{1}}&0&0&0&\ddots\\ c_{0}&w_{1}\frac{a_{1}}{a_{2}}&0&0&\ddots\\ -c_{0}\frac{b_{2}}{a_{3}}&c_{1}&w_{2}\frac{a_{2}}{a_{3}}&0&\ddots\\ c_{0}\frac{b_{2}b_{3}}{a_{3}a_{4}}&-c_{1}\frac{b_{3}}{a_{4}}&c_{2}&w_{3}\frac{a_{3}}{a_{4}}&\ddots\\ -c_{0}\frac{b_{2}b_{3}b_{4}}{a_{3}a_{4}a_{5}}&c_{1}\frac{b_{3}b_{4}}{a_{4}a_{5}}&-c_{2}\frac{b_{4}}{a_{5}}&c_{3}&\ddots\\ \vdots&\ddots&\ddots&\ddots&\ddots\end{bmatrix}.$
We now provide a sufficient condition for $F_{w}$ to be bounded. Only an outline of the proof will be given. For the particular cases of $p=2$ and the unweighted forward shift $F$, we refer to [2].
Theorem 4.1.
If
$\beta_{1}:=\sup_{n\geq 0}~{}\Big{\{}\left\lvert\frac{w_{n}a_{n}}{a_{n+1}}\right\rvert,~{}|c_{n}|\Big{\}}<\infty,~{}\text{and}~{}\beta_{2}:=\sum_{n=3}^{\infty}\sup_{j\geq 0}~{}\left|c_{j}\frac{\prod_{k=3}^{n}b_{j+k-1}}{\prod_{k=3}^{n}a_{j+k}}\right|<\infty,$
then $F_{w}$ is a bounded operator on $\ell^{p}_{a,b}(\mathbb{D})$ and $c_{0,a,b}(\mathbb{D})$, where $1\leq p<\infty$. Moreover, $\|F_{w}\|\leq\beta_{1}+\beta_{2}$.
Proof.
The matrix of $F_{w}$ can be formally written as $[F_{\alpha}]$+$\sum_{n=2}^{\infty}[F_{n}]$. Here, $[F_{\alpha}]$ is the matrix of the standard weighted forward shift $F_{\alpha}(e_{i})\mapsto\alpha_{i}e_{i+1}$ on $\ell^{p}$, $i\geq 1$, having weights
$\alpha_{i}=\frac{w_{i}a_{i}}{a_{i+1}},\hskip 28.45274pt(i\geq 0),$
on $\ell^{p}$. Note that the matrix $[F_{n}]$ is obtained by deleting all the entries of $[F_{w}]$, except those at the $n$-th subdiagonal, where $n\geq 2$. Observe that $[F_{n}]$ is the matrix of suitable powers of a weighted forward shift $F_{n}$ for $n\geq 2$.
It follows, by the first assumption in the theorem, that the weighted shifts $F_{\alpha}$ and $F_{2}$ are bounded on $\ell^{p}$. Since
$\|F_{n}\|=\sup_{j\geq 0}~{}\left|\frac{c_{j}b_{j+2}b_{j+3}\cdots b_{j+n-1}}{a_{j+3}a_{j+4}\cdots a_{j+n}}\right|,$
the second condition in the theorem gives that $F_{n}$ is bounded, and $\sum_{n\geq 3}\|F_{n}\|<\infty$, with respect to the operator norm. Hence, the shift $F_{w}$ is bounded. This completes the proof of the theorem. ∎
Remark 4.2.
The following conditions for the boundedness of $F_{w}$ are simpler and stronger, compared to those in the above theorem. If
$\sup_{n\geq 0}~{}\left\lvert\frac{w_{n}a_{n}}{a_{n+1}}\right\rvert<\infty\hskip 11.38092pt\text{and}\hskip 11.38092pt\lim\sup_{n}\left\lvert\frac{b_{n}}{a_{n+1}}\right\rvert<1,$ (4.3)
then the assumptions in Theorem 4.1 are easily satisfied so that $F_{w}$ is bounded.
To infer the dynamics of $F_{w}^{*}$, we need the following. Its proof is omitted since it is similar to the bilateral version proved earlier. Also, we refer to [32] wherein the authors studied the chaos for $F^{*}$ in the tridiagonal spaces.
Proposition 4.3.
Let $\mathcal{E}$ be a Banach space of analytic functions on the unit disc, having evaluation functionals continuous. Let $k_{n}$, $n\geq 0$, denote the $n$-th Taylor coefficient functional. Then the span of $\{k_{n}:n\geq 0\}$ is weak-dense in $\mathcal{E}^{*}$, and norm dense if $\mathcal{E}$ is reflexive.
In the theorem below, we state the results for hypercyclicity only. For mixing of $F_{w}^{*}$, “liminf” is replaced by “lim”, and we do not state it.
Theorem 4.4.
Let $\ell^{p}_{a,b}(\mathbb{D}),$ $1<p<\infty,$ be the analytic Banach space over the unit disc $\mathbb{D}$ corresponding to $a=\{a_{n}\}_{n=0}^{\infty}$ and $b=\{b_{n}\}_{n=0}^{\infty}$. Then, the following hold for the adjoint $F_{w}^{*}$, assuming that $F_{w}$ is bounded.
  1. (i)
    If
    $\liminf_{n\rightarrow\infty}~{}\frac{|a_{\nu+n}|+|b_{\nu+n-1}|}{\prod_{k=0}^{n-1}|w_{\nu+k}|}=0,~{}\forall~{}\nu\geq 0,$ (4.4)
    then $F_{w}^{*}$ is hypercyclic.
  2. (ii)
    Conversely, if $F_{w}^{*}$ is hypercyclic, then
    $\sup_{n\geq 0}~{}\frac{\prod_{k=0}^{n-1}|w_{k}|}{|a_{n}|}\left(\sum_{j=0}^{\infty}\left|\frac{\prod_{k=0}^{j-1}b_{n+k}}{\prod_{k=0}^{j-1}a_{n+k+1}}\right|^{p}\right)^{1/p}=\infty.$
  3. (iii)
    The adjoint operator $F_{w}^{*}$ is always supercyclic.
  4. (iv)
    If the conditions in (4.3) are satisfied, then $F_{w}^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$ if and only if
    $\liminf_{n\rightarrow\infty}\left|\frac{a_{\nu+n}}{\prod_{k=0}^{n-1}w_{\nu+k}}\right|=0,~{}~{}\forall~{}~{}\nu\in\mathbb{N}.$ (4.5)
Proof.
(i) First suppose that (4.4) holds, and we will use Gethner-shapiro criterion to show that $F_{w}^{*}$ is hypercyclic. Consider $D:=\text{span}\{k_{\nu}:\nu\geq 0\}$. Then $D$ is dense in $(\ell^{p}_{a,b}(\mathbb{D}))^{*},$ by the previous proposition. Define $S:D\rightarrow D$ by
$S(k_{\nu})=\frac{1}{w_{\nu}}k_{\nu+1},$
$\forall~{}~{}\nu\geq 0.$ Since $F_{w}^{*}S(f)=f$ and $(F_{w}^{*})^{n}f\to 0,$ as $n\to\infty,$ $\forall$ $f\in D,$ we just need to show that there exists $(n_{k})$ such that $S^{n_{k}}(k_{\nu})$ tends to 00, as $k\to\infty,$ for $\nu\geq 0$. Indeed, Proposition 2.2 along with our assumption in (i) gives that, for some $(n_{k})$,
$\lVert S^{n_{k}}(k_{\nu})\rVert=\left\lVert\frac{k_{\nu+n_{k}}}{w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}}\right\rVert\leq M_{1}\frac{|a_{\nu+n_{k}|}+|b_{\nu+n_{k}-1}|}{|w_{\nu}w_{\nu+1}\cdots w_{\nu+n_{k}-1}|}\longrightarrow 0,$
as $k\rightarrow\infty$, where $M_{1}$ is a constant. Thus, $F_{w}^{*}$ satisfies the Gethner-Shapiro criterion.
Conversely, if $F_{w}^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$, then by the previously mentioned result of Bonet, we have
$\sup_{n\geq 1}\|w_{0}w_{1}\cdots w_{n-1}z^{n}\|=\sup_{n\geq 1}\|F_{w}^{n}(1)\|=\infty.$
By the norm expression (4.2), we get the result in (ii).
(iii) The supercyclicity criterion with $D$ and $S$ as above, taking $n_{k}=k$, $k\geq 1$, is satisfied. Indeed, by Proposition 3.2, for each $\nu\geq 0$ the sequence $F_{w}^{*n}(k_{\nu})$ vanishes for large $n$.
(iv) By the assumptions in (4.3), $F_{w}$ becomes a bounded operator. Also, one can find $r<1$ and $N\in\mathbb{N}$ such that $\big{|}b_{n}/a_{n+1}\big{|}<r$ for every $n\geq N$. Hence, the conditions that appear in (i) and (ii) are easily seen to be equivalent.
Our next result on the chaos of $F_{w}^{*}$. In order to apply the chaoticity criterion, we need a representation of $f_{n}^{*}$ in terms of $k_{n}$. The proof is omitted as it runs parallel to the bilateral case, see Proposition 3.11.
Proposition 4.5.
Let $k_{n}$ denote the co-ordinate functional $f\mapsto\frac{f^{(n)}(0)}{n!},$ defined on $\ell^{p}_{a,b}(\mathbb{D})$ or $c_{0,a,b}(\mathbb{D})$, where $n\geq 0$. We have
$f_{n}^{*}=\frac{1}{a_{n}}k_{n}+\sum_{j=1}^{n}(-1)^{j}\frac{b_{n-1}b_{n-2}\cdots b_{n-j}}{a_{n}a_{n-1}\cdots a_{n-j}}k_{n-j},$
for all $n\geq 1$ and $f_{0}^{*}=\frac{1}{a_{0}}k_{0}.$
We now have necessary and sufficient conditions for $F_{w}^{*}$ to be chaotic on the dual of $\ell^{p}_{a,b}(\mathbb{D})$, stated below. We recall that, in [32] the authors provided a sufficient condition for the adjoint $F^{*}$ of the unweighted shift $F$ to be chaotic on a general analytic reproducing kernel space over $\mathbb{D}$ which includes the case of $\ell^{2}_{a,b}(\mathbb{D})$.
Theorem 4.6.
Assume that the conditions stated in (4.3) are satisfied for $\ell^{p}_{a,b}(\mathbb{D})$, where $1<p<\infty$. Consider the following statements:
  1. (i)
    $F_{w}^{*}$ is chaotic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$
  2. (ii)
    $F_{w}^{*}$ has a non-trivial periodic vector.
  3. (iii)
    $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{q}<\infty.$
Then $(iii)\Rightarrow(i)\Rightarrow(ii)$. In addition, if $\frac{1}{p}+\frac{1}{q}=1,$ and
$\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}b_{n+k-1}}{\prod_{k=1}^{j}a_{n+k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty.$
then $(ii)\Rightarrow(iii)$.
Proof.
Clearly, (i) gives (ii).
To see that (ii) implies (iii), let $\psi=\sum_{m=0}^{\infty}\lambda_{m}f_{m}^{*}$ be a non-zero periodic vector for $F_{w}^{*}$ on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$, where $f_{m}^{*}$, $m\geq 0$, forms an unconditional Schauder basis for $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$. Now, using Proposition 4.5, we get
$\psi=\sum_{m=o}^{\infty}A_{m}k_{m},~{}~{}~{}~{}\text{where}~{}~{}~{}A_{m}=\lambda_{m}\frac{1}{a_{m}}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{m+j}\frac{b_{m}b_{m+1}\cdots b_{m+j-1}}{a_{m}a_{m+1}\cdots a_{m+j}},$
for all $m\geq 0.$ Now, let $\nu\in\mathbb{N}$ be such that $(F_{w}^{*})^{\nu}\psi=\psi$. Then $(F_{w}^{*})^{n\nu}\psi=\psi$ for all $n\geq 1$. Equating the coefficient of $k_{0}$ both sides, we get
$\frac{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}{a_{n\nu}}\left(\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right)=\frac{1}{a_{0}}\left(\lambda_{0}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{j}\frac{b_{0}\cdots b_{j-1}}{a_{1}\cdots a_{j}}\right).$
Since the series in the right hand side converges, we get that
$C\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{q}=\left\lvert a_{0}\right\rvert^{q}\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right\rvert^{q},$
for some $C>0$. Note that using Hölder inequality, we can get
$\sum_{n=1}^{\infty}\left\lvert\lambda_{n\nu}+\sum_{j=1}^{\infty}(-1)^{j}\lambda_{n\nu+j}\frac{b_{n\nu}\cdots b_{n\nu+j-1}}{a_{n\nu+1}\cdots a_{n\nu+j}}\right\rvert^{q}<\infty,$
as $\{\lambda_{n}\}_{n}\in\ell^{q}(\mathbb{N})$ and $\sum_{n\geq 0}\left(\sum_{j\geq 1}\left\lvert\frac{\prod_{k=1}^{j}b_{n+k-1}}{\prod_{k=1}^{j}a_{n+k}}\right\rvert^{p}\right)^{\frac{q}{p}}<\infty.$ Therefore,
$\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu}}{w_{n\nu-1}w_{n\nu-2}\cdots w_{0}}\right\rvert^{q}<\infty.$ Similarly, equating the coefficient of $k_{j},j=1,2,\cdots,\nu-1$, and performing calculations as above, it is possible to observe that $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n\nu+j}}{w_{n\nu+j-1}w_{n\nu+j-2}\cdots w_{j}}\right\rvert^{q}<\infty.$ We conclude that $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{q}<\infty.$
Now we prove that (iii) implies (i), by applying the chaoticity criterion to show that $F_{w}^{*}$ is chaotic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Consider $D=\text{span}\{k_{n}:n\geq 0\}$ which is dense in $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Define $S:D\rightarrow D$ by $S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\geq 0.$ Clearly, $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ and $\sum_{n=0}^{\infty}(F_{w}^{*})^{n}(f)$ converges unconditionally for each $f\in D.$ We just need to show that $\sum_{n=0}^{\infty}S^{n}(f)$ converges unconditionally, for each $f\in D.$ It suffices to prove that $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is unconditionally convergent in $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Recall that $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\geq 1$ and $k_{0}=a_{0}f_{0}^{*},$ and so we have
$\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ $=$ $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}(a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*})$
$=$ $\frac{a_{1}}{w_{0}}\frac{b_{0}}{a_{1}}f_{0}^{*}+\sum_{n=1}^{\infty}\left(\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}+\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\frac{b_{n}}{a_{n+1}}\right)f_{n}^{*}.$
As $\limsup_{n}|b_{n}/a_{n+1}|<1$, one gets $N\in\mathbb{N}$ and $r<1$ such that $|b_{n}/a_{n+1}|<r$ for all $n\geq N$. Hence,
$\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}+\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\frac{b_{n}}{a_{n+1}}\right\rvert\leq\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert+r\left\lvert\frac{a_{n+1}}{w_{0}w_{1}\cdots w_{n}}\right\rvert,$
for all $n\geq N$. The right hand side of the above inequality is the $\ell^{q}$ element
$\left\{\frac{a_{n+1}}{w_{0}\cdots w_{n}}\right\}_{n=1}^{\infty},$
and hence it is absolutely $q$-summable. Consequently, the series $\sum_{n\geq N}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is convergent. The unconditional convergence occurs because $\{f_{n}^{*}\}$ is an unconditional Schauder basis. Hence, $F_{w}^{*}$ satisfies the chaoticity criterion, and (i) follows.
We illustrate the above results on the unilateral weighted shift $F_{w}$ with an example.
Example 4.7.
Choose $a_{0}=b_{0}=w_{0}=1,$ and $a_{n}=\frac{1}{n2^{n-1}},~{}b_{n}=\frac{1}{(n+1)2^{2n}},~{}w_{n}=4,~{}n\geq 1$. Then $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}),$ as it satisfies the conditions of Theorem 4.1. By Theorems 4.4 and 4.6, the adjoint $F_{w}^{*}$ is mixing and chaotic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*},$ $1<p<\infty$.
A more general result on the dynamics of $F_{w}^{*}$ is given as follows in the next theorem. We omit its proof. A bilateral analogue was stated in Theorem 3.13.
Theorem 4.8.
Let $\mathcal{E}$ be a Banach space of analytic functions on the unit disc $\mathbb{D}$, having norm-separable dual $\mathcal{E}^{*}$. Let $k_{n}\neq 0$ for all $n\geq 0$. The following hold, assuming that the operator $F_{w}$ is bounded on $\mathcal{E}$.
  • (i)
    If $\inf_{n\geq 0}\big{(}|\prod_{k=1}^{n}w_{k}|^{-1}\|k_{n+1}\|\big{)}=0$, then $F_{w}^{*}$ is hypercyclic.
  • (ii)
    If $\lim_{n\rightarrow\infty}\big{(}|\prod_{k=1}^{n}w_{k}|^{-1}\|k_{n+1}\|\big{)}=0$, then $F_{w}^{*}$ is mixing.
  • (iii)
    If $\mathcal{E}$ is reflexive, and
    $\sum_{n\geq 1}\left|\left(\prod_{k=1}^{n}w_{k}\right)^{-1}\frac{f^{(n+1)}(0)}{{(n+1)}!}\right|<\infty,$
    for all $f\in\mathcal{E}$, then $F_{w}^{*}$ is chaotic.
  • (iv)
    The adjoint $F_{w}^{*}$ is supercyclic.
We close this section with results on compact perturbations, essential spectrum, and hypercyclic subspaces for $F_{w}^{*}$ when $F_{w}$ is defined on $\ell^{p}_{a,b}(\mathbb{D})$. Recall that the essential spectrum of a unilateral shift on $\ell^{2}$ was obtained in Shields [36]. The same proof works for $\ell^{p}$, $1<p<\infty$.
Proposition 4.9.
Suppose that the weighted shift $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathbb{D})$, where $1<p<\infty$. If $c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\rightarrow 0$, as $n\rightarrow\infty$, then $F_{w}$ is similar to a perturbation $F_{\alpha}+K$ for some compact operator $K$ and a weighted forward shift $F_{\alpha}$ defined on $\ell^{p}$, having weights $\alpha_{i}=\frac{w_{i}a_{i}}{a_{i+1}}$, and hence $\sigma_{e}(F_{w})=\sigma_{e}(F_{\alpha})$.
Proof.
Recall from the matrix representation of $F_{w}$ on $\ell^{p}_{a,b}(\mathbb{D})$, it follows that $F_{w}$ is similar to an operator of the form $K+F_{\alpha}$ for some bounded operator $K$ and a weighted forward shift $F_{\alpha}$ on $\ell^{p}$. Here, $K=\sum_{n\geq 2}F_{n}$. Since $c_{n}$ is a null sequence, each $F_{n}$ becomes compact, and so $K$ is compact. As the essential spectrum is invariant under compact perturbations, we also have $\sigma_{e}(F_{w})=\sigma_{e}(F_{\alpha})$. ∎
Corollary 4.10.
Suppose that $F_{w}^{*}$ is hypercyclic on the dual $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$, where $1<p<\infty$, and $11/p+1/q=$ 1. Under the assumptions as in the previous proposition, the adjoint $F_{w}^{*}$ possesses a hypercyclic subspace if and only if $F_{\alpha}^{*}$ has a hypercyclic subspace in $\ell^{q}(\mathbb{Z})$.
5. Notable differences: Eigenvectors, periodic vectors, and a zero-one law
We establish a couple of surprising results which show that the forward weighted shifts $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $\ell^{p}_{a,b}(\mathbb{D})$ are very different from the classical bilateral/unilateral weighted shifts. The adjoint of an unweighted forward shift operator $F$ can have a non-zero periodic vector without $F$ being hypercyclic, which is in sharp contrast to the case of the classical weighted shifts. On another direction, a well known result states that if a bilateral weighted shift on $\ell^{p}(\mathbb{Z})$ has an orbit having a non-zero limit point, then the operator has a dense orbit, cf. [11]. (This property is known as a zero-one law of orbital limit points for an operator, also see [1], [8], and [10].) Such a law does not hold for $F_{w}$ defined on $\ell^{p}_{a,b}(\mathtt{A})$ and $\ell^{p}_{a,b}(\mathbb{D})$, as proved below. Further distinguishing characteristics are obtained in this section. These interesting properties of $F^{*}$ will be based on a suitable extension of the domain of the map $z\mapsto ev_{z}$, where $ev_{z}$ denotes the evaluation functional at $z$. The space we consider here, is based on Theorem $7$ in [2], where the authors studied the unweighted shift on a tridiagonal kernel space.
Proposition 5.1.
Let $|\lambda|=1$. Consider a space of the form $\ell^{p}_{a,b}(\mathbb{D})$, $1<p<\infty$, wherein
$a_{n}=1,~{}0<|b_{n}|<1,~{}(n\geq 0).$
Suppose that the unweighted forward shift $F$ is bounded, and that
$\lim_{n\rightarrow+\infty}(b_{n}-b_{n+1})=0$.
Then, $F$ is non-hypercyclic, and it is similar to $S+K$, where $S$ is the right shift on $\ell^{p}$ and $K$ is some compact operator. Also, the following hold, assuming that
$\sum_{n\geq 0}|1+\lambda b_{n}|^{q}<\infty,$
$11/p+1/q=$ 1.
  1. (i)
    The zero-one law of orbital limit points fails for $F^{*}$.
  2. (ii)
    If $\lambda$ is a root of the unity, then $F^{*}$ has a non-trivial periodic vector in $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$.
Proof.
By Theorem $4.1$ we can see that $F$ is a bounded operator, which is similar to a “compact + right shift” on $\ell^{p}(\mathbb{N})$.
By Theorem 4.4 (ii), we see that $F^{*}$ is not hypercyclic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$.
The forward shift operator $F$ in $\ell^{p}_{a,b}(\mathbb{D})$ is given by $F(f(z))=zf(z),$ for all $f\in\ell^{p}_{a,b}(\mathbb{D})$ and $z\in\mathbb{D}.$ Now, since $F$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}),$ for $f\in\ell^{p}_{a,b}(\mathbb{D})$ and $z\in\mathbb{D},$ we have
$F^{*}(ev_{z})f=(ev_{z}\circ F)f=ev_{z}(zf)=zf(z)=z(ev_{z})f,$
that is, $F^{*}(ev_{z})=z(ev_{z}).$ It shows that $z\in\mathbb{D}$ is an eigenvalue of $F^{*}$ corresponding to the eigenvector $ev_{z}$ on $(\ell^{p}_{a,b}(\mathbb{D}))^{*},$ provided $ev_{z}$ is a non-zero functional. This fact is well known, cf. Godefroy and Shapiro [19].
In our case, $a_{n}=1,$ $0<\lvert b_{n}\rvert<1,$ $n\geq 0,$ and $\lambda$ is unimodular. Now,
$\sum_{n=0}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=0}^{\infty}|1+\lambda b_{n}|^{q}<\infty,$ (5.1)
for $1<q<\infty.$ Therefore, the natural domain of $\ell^{p}_{a,b}(\mathbb{D})$ is given by $\text{Dom}_{(p)}=\mathbb{D}\cup\{\lambda\}.$ Note that if $f\in\ell^{p}_{a,b}(\mathbb{D})$, then $f(z)=\sum_{n\geq 0}\widehat{f}(n)f_{n}(z)$, and $\sum_{n\geq 0}|\widehat{f}(n)|^{p}<\infty$. In view of (5.1), we obtain that each $f(z)$ in $\ell^{p}_{a,b}(\mathbb{D})$ extends to $z=\lambda$, and thus, it is meaningful to define
$ev_{\lambda}(f)=f(\lambda):=\sum_{n\geq 0}\widehat{f}(n)f_{n}(\lambda),$
for all $f$. Since $|f(\lambda)|\leq M_{\lambda}\|f\|$ for some constant $M_{\lambda}>0$, we have that $ev_{\lambda}\in(\ell^{p}_{a,b}(\mathbb{D}))^{*}$. It follows that $F^{*}(ev_{\lambda})=\lambda(ev_{\lambda})$, i.e. $ev_{\lambda}$ is an eigenvector of $F^{*}$. Note that $ev_{\lambda}$ is non-zero as, for example, $ev_{\lambda}(g)\neq 0$, where
$g(z)=\sum_{n\geq 0}\frac{1}{\lambda^{n}2^{n}}\frac{|1+\lambda b_{n}|}{1+\lambda b_{n}}f_{n}(z)\in\ell^{p}_{a,b}(\mathbb{D}).$
If $\lambda$ is a root of unity, say $\lambda^{m}=1$, then $F^{*m}(ev_{\lambda})=ev_{\lambda}$. Hence, $F^{*}$ has a nontrivial periodic vector, namely $ev_{\lambda}$, which gives (ii).
To prove (i), first consider the orbit of $f:=ev_{z}+ev_{\lambda}$ under $F^{*}$, for any $z\in\mathbb{D}$. If $\lambda^{j}=1$, then $(F^{*})^{jn}(ev_{z}+ev_{\lambda})\rightarrow ev_{\lambda}$, as $n\rightarrow\infty$. In the other case when $\lambda$ is not a root of unity, we use a well known fact, for such $\lambda$, that the sequence $\{\lambda^{n}:n\geq 1\}$ is dense in the unit circle of $\mathbb{C}$, and obtain a sequence $(j_{n})$ such that $\lambda^{j_{n}}\rightarrow 0$, as $n\rightarrow\infty$. Consequently, we have $(F^{*})^{j_{n}}(ev_{z}+ev_{\lambda})\rightarrow ev_{\lambda}$, as $n\rightarrow\infty$. Thus, the zero-one aw of orbital limit points fails for $F^{*}$. We would like to recall that $F^{*}$ is not hypercyclic. ∎
Example 5.2.
For $|\lambda|=1$, if we set
$a_{n}=1~{}\text{and}~{}b_{n}=\frac{-1}{\lambda}\frac{n+1}{n+2},~{}n\geq 0,$
$F^{*}$ has the properties stated in (i) and (ii) of the above proposition.
Similarly, we have a bilateral analogue of the above result.
Proposition 5.3.
Let $|\lambda|=1$, and let $\mathtt{A}$ be an annulus $r<|z|<1$. Consider $\ell^{p}_{a,b}(\mathtt{A})$, $1<p<\infty$, where $10<|b_{n}|<$ 1 for all $n\in\mathbb{Z}$, and
$a_{n}=1,~{}n\geq 0,~{}~{}~{}~{}\text{and}~{}~{}~{}~{}0<|a_{n}|<1,~{}n\leq-1.$
If the right shift $F$ is bounded, and
$\sum_{n=2}^{\infty}\left\lvert\prod_{k=1}^{n-1}\frac{a_{-k}}{b_{-k-1}}\right\rvert^{p}<\infty,~{}~{}~{}~{}\sum_{n\geq 0}|1+\lambda b_{n}|^{q}<\infty~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n\geq 1}|a_{-n}+\lambda b_{-n}|^{q}<\infty,$
$11/p+1/q=$ 1, then the following hold.
  1. (i)
    The adjoint $F^{*}$ is non-hypercyclic.
  2. (ii)
    The zero-one law of orbital limit points fails for $F^{*}$.
  3. (iii)
    If $\lambda$ is a root of the unity, then $F^{*}$ has a non-trivial periodic vector in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$.
Proof.
In view of Proposition 3.15 (ii), we see that $F^{*}$ is not hypercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$. Indeed, using the basis expansion of $z^{n}$, $n\geq 1$, the set
$\{\|z^{n}\|:n\geq 1\}$
can be shown to be bounded.
Now,
$\sum_{n=-\infty}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=1}^{\infty}|a_{-n}+\lambda b_{-n}|^{q}+\sum_{n=0}^{\infty}|1+\lambda b_{n}|^{q}<\infty,$ (5.2)
for $1<q<\infty.$ If $f\in\ell^{p}_{a,b}(\mathtt{A})$, then $f(z)=\sum_{n=-\infty}^{\infty}\widehat{f}(n)f_{n}(z)$, and $\sum_{n}|\widehat{f}(n)|^{p}<\infty$. In view of (5.2), the evaluation functional $ev_{\lambda}$ is well-defined and it is given by $ev_{\lambda}(f)=f(\lambda):=\sum_{n=-\infty}^{\infty}\widehat{f}(n)f_{n}(\lambda)$, for all $f$. Also, $ev_{\lambda}$ is continuous and non-zero functional because $ev_{\lambda}(g)\neq 0$, where
$g(z)=\sum_{n\geq 1}\frac{\lambda^{n}}{2^{n}}\frac{|a_{-n}+\lambda b_{-n}|}{a_{-n}+\lambda b_{-n}}f_{-n}(z)+\sum_{n\geq 0}\frac{1}{\lambda^{n}2^{n}}\frac{|1+\lambda b_{n}|}{1+\lambda b_{n}}f_{n}(z)\in~{}~{}\ell^{p}_{a,b}(\mathtt{A}).$
Hence, $ev_{\lambda}$ is an eigenvector of $F^{*}$. The remaining proof, similar to the unilateral case, yields $(ii)$ and $(iii)$. ∎
Example 5.4.
Let $\lvert\lambda\rvert=1,$ and consider
$a_{n}=1,~{}~{}~{}b_{n}=-\frac{-1}{\lambda}\frac{n+1}{n+2},~{}\forall n\geq 0,~{}~{}~{}\text{and}~{}~{}~{}a_{n}=\frac{2^{n}}{2^{2}},~{}~{}~{}b_{n}=-\frac{1}{\lambda}2^{n},~{}~{}~{}~{}\forall n\leq-1.$
Then, we have
$c_{-1}=-\frac{1}{2\lambda}\frac{7}{8},~{}~{}~{}~{}~{}~{}~{}c_{n}=\frac{b_{n}}{a_{n+2}}-\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}=-\frac{1}{\lambda}\frac{n+1}{n+2}+\frac{1}{\lambda}\frac{n+2}{n+3}=\frac{1}{\lambda}\frac{1}{(n+2)(n+3)},~{}~{}~{}(n\geq 0)$
and
$d_{-n}=\frac{a_{-n}}{b_{-n}}-\frac{a_{-n+1}}{b_{-n+1}}=-\frac{\lambda 2^{n}}{2^{n+2}}+\frac{\lambda 2^{n-1}}{2^{n+1}}=0,~{}~{}~{}(n\geq 2).$
Therefore, we get the matrix of $F$ corresponding to $\{f_{n}\}$
$[F]:=\begin{bmatrix}\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots&\vdots\\ \ddots&0&\frac{\alpha\lambda^{3}}{2}\frac{7}{8}&-\frac{\alpha\lambda^{4}}{2.2}&-\frac{\alpha\lambda^{5}}{2.3}&-\frac{\alpha\lambda^{6}}{2.4}&\cdots\\ \ddots&0&\alpha\lambda^{2}\frac{7}{8}&-\frac{\alpha\lambda^{3}}{2}&-\frac{\alpha\lambda^{4}}{3}&-\frac{\alpha\lambda^{5}}{4}&\ddots\\ \ddots&\frac{1}{2}&2\alpha\lambda\frac{7}{8}&-\alpha\lambda^{2}&-\frac{2\alpha\lambda^{3}}{3}&-\frac{2\alpha\lambda^{4}}{4}&\ddots\\ \ddots&0&\frac{1}{8}(1+7\alpha)&\bm{-\frac{\alpha\lambda}{1.2}}&-\frac{\alpha\lambda^{2}}{1.3}&-\frac{\alpha\lambda^{3}}{1.4}&\ddots\\ \ddots&0&-\frac{(1-\alpha)}{2\lambda}\frac{7}{8}&1-\frac{\alpha}{2^{2}}&-\frac{\alpha\lambda}{2.3}&-\frac{\alpha\lambda^{2}}{2.4}&\ddots\\ \ddots&0&-\frac{(1-\alpha)}{3\lambda^{2}}\frac{7}{8}&\frac{(1-\alpha)}{\lambda}\frac{1}{2.3}&1-\frac{\alpha}{3^{2}}&-\frac{\alpha\lambda}{3.4}&\ddots\\ \ddots&0&-\frac{(1-\alpha)}{4\lambda^{3}}\frac{7}{8}&\frac{(1-\alpha)}{\lambda^{2}}\frac{1}{2.4}&\frac{(1-\alpha)}{\lambda}\frac{1}{3.4}&1-\frac{\alpha}{4^{2}}&\ddots\\ \vdots&\ddots&\ddots&\ddots&\ddots&\ddots&\ddots\end{bmatrix}.$ (5.3)
Note that the boundedness of $F$ cannot be conluded using Theorem 3.3. So, to show that $F$ is bounded, recall the matrix representation of $F$ and invoke the proof of Theorem 3.3. We see that $\sum_{i=1}^{\infty}\lVert E_{-i}\rVert+\sum_{i=1}^{\infty}\lVert E_{i}\rVert<\infty.$ Therefore, $F$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}).$ For the given values of $\{a_{n}\}_{n\in\mathbb{Z}}$ and $\{b_{n}\}_{n\in\mathbb{Z}},$ we get
$\sum_{n=-\infty}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=1}^{\infty}\left\lvert\left(\frac{1}{2^{2}2^{n}}-\frac{1}{\lambda 2^{n}}\lambda\right)\lambda^{-n}\right\rvert^{q}+\sum_{n=0}^{\infty}\left\lvert\left(1-\frac{1}{\lambda}\frac{n+1}{n+2}\lambda\right)\lambda^{n}\right\rvert^{q}<\infty,$ (5.4)
for $1<q<\infty.$ Here $r=\frac{1}{2}$ and $R=1.$ Therefore the natural domain of $\ell^{p}_{a,b}(\mathtt{A})$ is given by $\text{Dom}_{(p)}=\mathtt{A}\cup\{\lambda\},$ where $\mathtt{A}$ is a annulus $\frac{1}{2}<\lvert z\rvert<1.$ Then, $F^{*}$ has the properties stated in (i), (ii), and (iii) of the above proposition.
In view of these propositions, the following corollaries are immediate.
Corollary 5.5.
There exists a weighted backward shift $B_{w}$ with respect to the standard Schauder basis in the sequence space $\ell^{p}$ and a compact operator $K$ such that $B_{w}+K$ has non-trivial periodic vectors, but $B_{w}+K$ is not hypercyclic.
There is yet another difference.
Corollary 5.6.
There exists a weighted forward shift $F_{w}$ on a space of the form $\ell^{p}_{a,b}(\mathbb{D})$, which is not similar to a weighted shift on $\ell^{p}$. A similar conclusion holds for bilateral shifts as well.
Acknowledgments. Bibhash Kumar Das is supported financially by CSIR research fellowship (File No.: 09/1059(0037)/2020-EMR-I), and Aneesh Mundayadan is partially funded by a Start-Up Research Grant of SERB-DST (File. No.: SRG/2021/002418).